A Complete Bibliography of the Journal on Computing and Cultural Heritage (JOCCH)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 December 2024 Version 1.42

Title word cross-reference

#AtHomeAtTheMuseum [MLL^{+}22]. #MuseumAtHome [MLL^{+}22].

2.5 [Sap19]. 3 [BPV^{+}15, BCFS15, CNS15, CMGFOAFH23, CC23, CS08, DCC^{+}10, DNS23, DSH^{+}15, ES16, ESW20, EES^{+}20, GMLLHFTT21, GSB18, GBG^{+}22, GA16, HSB^{+}09, HRS^{+}16, HSMAK^{+}22, HWSK14, KRF^{+}14, KFH09, KKRS22, LBM^{+}12, LBC^{+}15, LCD^{+}15, LTP21, MSWS20, MWL^{+}16, MKM21, NTM^{+}23, PG15, PHB22, RGRG09, RRC^{+}10, RAB^{+}16, RTA22, SB22, SM18, YH13, ZA24, ZPS^{+}16, ZSV^{+}16, vBTS^{+}22]. < [MANC24]. > [MANC24]. n [KT16].

-Grams [KT16].

14th [CD18]. 15th [GLF^{+}20]. 15th-Century [GLF^{+}20]. 1650 [RPA^{+}10]. 19 [BKPW24, MLL^{+}22]. 19100 [MGB18]. 19th [LBM^{+}12, LBC^{+}15]. 19th-Century [LBC^{+}15].

2 [DS21, SPC^{+}23]. 2-NN [DS21]. 2019 [ERG^{+}19]. 2nd [DNR17b].

360-Degree [ZM20]. 3D [DPG22, LXW24, MPPB24]. 3DEmo [LA24]. 3VR [VQP^{+}24].

Abbey [PZF11]. Acceptance [CLCL20, HMS20, JML24]. Access [BZ24, BSO22, CCAS23, DOHO08, KHO^{+}23, PLSL22, PMGZ^{+}24, RNP21, SBO^{+}17, VMH^{+}18, ATP11, IA14, SLD^{+}12]. Accessibility [APH24, CSSC^{+}24, SRH^{+}20]. Accessible [WRKK^{+}22]. Accessing [LPD16]. Accurate [SKK^{+}20]. Acoustic [AY23]. Acquisition [NKRS15, RFA19, CCD^{+}11, DCC^{+}10]. across [RCB^{+}18]. active [KSZ^{+}11, VC11]. Activities [RFA19]. Acuity [PPBH21]. Adaptability [TLL^{+}24]. adaptation [AL10]. Adaptive [GSL15, RTF22, KSSW24]. Added [HMP17]. Addendum [Nic19a]. Adding [KC20]. adjacent [Erk22]. Advances [BM22]. Advancing [MNL^{+}21]. Adversarial [YZH^{+}24]. Aesthetic [SGMT21]. aesthetics [MA12]. Affect [RFA19]. Affected [KOIC20]. Affection [BCF^{+}21]. Affordances [Dry24]. Age [CFRCDSC21]. agenda [DOHO08]. Aggregate [ZDC21]. Aggregation [SZGE23]. Agnew [NVCD22]. Agonistic [DFSO21]. AI [BGP22, CBJN22, Sto22]. Aided [DKY24, CGL^{+}22]. AIR [DNS23]. Alexa [VYBG22]. Alexandria [CCD^{+}21]. Algerian [AKK^{+}24]. Algiers [ZDC21]. Algorithm [FUD^{+}19, GWZ^{+}24, HWSK14, NDN^{+}21, SF18, VQP^{+}24]. Algorithmic [Ben23]. Algorithms [YPGR17]. Alignment [PFZG23]. Altarpiece [CDV21, PVPMP24]. altarpieces [PVPMP24]. Alternative [MWL^{+}16, CKA13]. Amazonia [PMHdF24]. Ambient [HMS20]. American [KA20]. Amire [AAÇ22]. Amphitheatres [RTF22]. Analog [FCP24]. Analogue [CLN18]. analyses [YH13]. Analysing [SGMT21]. Analysis [AAÇ22, BM22, CSS^{+}17, CYH^{+}22, CMST22, CFRCDSC21, DLI^{+}17, GFE19, GBAA22, GCL^{+}19, GWZ^{+}24, JTL19, Kim22, KCJP24, NB17, Nic19a, OAB^{+}24, Pay23, PYR15, PHB22, RAD^{+}23, RAB^{+}19, RSM23, RTA22, SP16, Ulv22, WSA^{+}18, WZAM^{+}21, YPGR17, HS08, RPA^{+}10, ST12, ASC^{+}15]. analytical [GBG^{+}22]. Analytics [ACH24, BKPW24, LMWZ17]. Analyze [AFGG21]. Analyzing [ES16, FKS22, GSL15, SDF^{+}12]. anastylosis [TFK^{+}09]. Ancient [ASBD21, AFV16, COGP18, CPR^{+}24, CCZ^{+}17, FBP20, FHA^{+}21, HOGP17, LLB24, LXW24, LW16, LMWP24, LWZ^{+}24, MVCK^{+}23, PKO23, PFF^{+}15, SP16, Sap19, SBS24, SCDS15, SZF^{+}10, Tai23, WRKK^{+}22, DG11, GMMC09, HAK21, MQD^{+}23]. Andalusian [SB22]. Andean [BCH^{+}15]. Andrea [BG14]. Animation [KKRS22]. Annales [HH13]. Annotated [JVAP^{+}21, MPFC16]. Annotating [MWX^{+}12]. Annotation [FKW17, DRAT11]. annotations [YH13]. answering [SZGE23]. Antique [PMG^{+}20]. Apart [COGP18]. App [OAB^{+}24]. Appearance [EES^{+}20, NKRS15]. Application [BZ24, BSO22, CCF20, DS21, HLT23, PKO23, PLC^{+}20, TJTHW24]. Applications [CCD^{+}21, DSH^{+}15, HDZM16, HGLS23, KC20, MC22, OZY^{+}21, ORD^{+}21, PPF^{+}23, RJIH22, RKG^{+}24, SCM^{+}24, CS08, FPMT10]. Applied [AKK^{+}24, GMLLHFTT21, GAKL24, GSB18]. Applying [APHGG23, BSDG^{+}21, PP23, TLL^{+}24]. Appreciation [ZKB^{+}20, ZM20]. Approach [APHGG23, BV17, BCH^{+}15, FL20, FCL^{+}24, GFE19, GSB18, GCL^{+}19, GBG^{+}22, HVS^{+}15, HZ24, KA20, MKM21, MPPB24, OAB^{+}24, RDH15, SWZG23, SFB^{+}22, Tro22, VVSB14, Vio23, ALBA13, DI08]. Approaches [PMGZ^{+}24, VKS^{+}24, SS13]. Arabic [BEES15]. Aramaic [LLB24]. ARCA [MSTG18]. Archaeological [CMGFOAFH23, DNS23, EEG^{+}21, FUD^{+}19, MMGGA21, MGB18, MQB22, Pay23, PSBM15, RRJBMM16, SBVVSL^{+}15, VVSB14, WMJ^{+}18, YN20, ALBA13, MD12, SHFS12]. Archaeologist [PLP^{+}17]. Archaeology [BVLW22, CTRG19, DSH^{+}15, KTHP15, MSR^{+}17, MSTG18, PT19, HLB14, WD13]. Archeological [Ulv22, KLST13]. ArcheoTUI [RRC^{+}10]. Architectural [CLP24, FL20, IAY15, JK22, JK23, MM21, SU22, DG11]. Architecture [AKK^{+}24, DTM17, LW16, MKM21, ZDC21, HLZ12]. Archival [FYYG22, HMG22, PLSL22]. Archive [EKS^{+}22, NVCD22, PLP^{+}17, DMP^{+}17, SH22]. Archives [CDM^{+}24, CBJN22, CGL^{+}22, CFdS22, DAD^{+}22, FYYG22, GB22, GCM^{+}20, GGN22, HLZL18, LMWZ17, MANC24, PS22, RRDC23, YAGU09, dMKBD22, KFH09]. Archiving [FF22, Kim22, MA12, Tai23]. Archivists [HKvF22]. area [JZF23]. Argentino [SKD19]. argumentation [DKB11]. ARIADNE [MSR^{+}17]. Arnold [SE17]. Arrange [ZJCN24]. Arrigo [HSB^{+}09]. Art [AFGG21, CCAS23, DAD^{+}22, Fra20, FBB^{+}24, FMAG13, HKP^{+}22, Kim22, MGBM14, MHJ22, Por21, SSLdB24, SV24, TJTHW24, ZPS^{+}16, IA14, LLS^{+}12, ST12, CCAS23, VC21]. Art-related [CCAS23]. Artefacts [ESW20]. Artifacts [APH24, BCFS15, FF23, RRL^{+}14, SBVVSL^{+}15, SM18, CCD^{+}11, KRF^{+}14, MSN^{+}13, YH13]. Artificial [LMWP24]. Artistic [GSL15, HGCS11, SBvE^{+}23, Yan13]. Artists [VKS^{+}24]. Arts [APHGG23, BDK23, Ben23, SIS^{+}23, BB08]. Artwork [KCJP24, KSSW24]. Artworks [BPV^{+}15, FCP24, SGMT21, Sto22, VKS^{+}24, SLD^{+}12]. Asia [MKM21]. aspects [AL10]. Assembly [ASBD21]. assertions [DKB11]. Assessing [HMP17, KTHP15, PPBH21]. Assessment [DCC^{+}20, NTM^{+}23]. Assistant [AMP^{+}17]. Assisted [FKW17, WMJ^{+}18, RSM23]. Assisting [PGD^{+}22]. Assyriology [BM22]. Asymmetry [SWZG23]. ATHENA [PYR15]. Atina [FMG^{+}21]. Attended [MRFF24]. Audience [ESW20, EKS^{+}22].

Audio [CLN18, WM24]. Augmentation [NCM^{+}19, PLSL22, RFW^{+}22, YLFM22]. Augmentation-based [PLSL22]. Augmented [BPF^{+}18, CLP24, CWR^{+}23, CCFC23, CCD^{+}21, DBYA18, Dim22, DMPC14, EKS^{+}22, Hu24, JML24, KKK^{+}20, OZY^{+}21, PGMBR17, RRL^{+}14, SLRV23, SOÇT23, TSB21, XLL^{+}24]. Austria [Eic21]. authentic [Vio23]. Authentication [ZOZ21]. Authenticity [GLS^{+}24]. author [MBD^{+}13]. Authored [Sto22]. Authoring [BCF^{+}21, GHC21, PDK^{+}22, SPZ^{+}22, ZM20]. Authorities [HKvF22]. Authorship [CMST22, SCM^{+}24]. Automated [TS14, MDCS19]. Automatic [DDGP20, BGB24, DLI^{+}17, FUD^{+}19, HWSK14, JK23, MM23, MD12, MCV23, MSTG18, NWLZ23, NKRS15, PYR15, PG15, PSBM15, SBVVSL^{+}15, SBC^{+}22, Sto22, WZL^{+}21, YPGR17, ZLSD13, KML22]. automatized [HSMAK^{+}22]. Automatizing [PGD^{+}22]. Avatars [CPR^{+}24]. awaRe [AMP^{+}17]. Awareness [MC22]. Ayazini [SOÇT23].

B [PKO23]. B.C. [RPA^{+}10]. Bacchus [Erk22]. Background [FWZ23]. Baroque [GSB18]. Base [BCH^{+}15, PLC^{+}20]. Based [ASBD21, AASFB23, BCFS15, CAMR22, CSS^{+}17, CN23, DYW21, ES16, FST24, FKW17, Geo24, GLS^{+}24, HHJ^{+}24, HWSK14, JTL19, KSSW24, LW16, MRAB21, MBEB23, MGB18, MCKC24, MR15, NDN^{+}21, NWLZ23, OKY15, RBXM15, SM18, TWZF24, Tro22, TV23, WHC^{+}19, YPGR17, BV17, BBD^{+}24, DCC^{+}10, DRAT11, ESD^{+}22, ESKI21, HLB14, HLYH14, HH13, HGCS11, JVAP^{+}21, JML24, KCJP24, KKW22, KY24, LLA^{+}19, MPPB24, MSN^{+}13, OAAA21, PLSL22, RFW^{+}22, RNP21, SRWK13, SLD^{+}12, SBN22, STLL13, YLFM22]. Basic [ZJCN24, Arn08b]. Basilica [CCD^{+}21]. Bath [AY23]. Batik [WSHH24]. BBC [SH22]. Beacons [CAMR22]. Beauty [CRR^{+}14]. Beginners [ESKI21]. Behavior [MNL^{+}21]. Behaviors [CAMR22]. Behaviour [CC23, RBXM15]. Belo [CCN^{+}24]. Belonging [JDDAB^{+}24]. Benchmark [PP23, HS08]. Benefits [PGMBR17, LRD08]. BERT [BVLW22, HZ24]. BERT-enhanced [HZ24]. Best [AFGG21, CD18]. Between [RTA22, TF23, ASC12, DCC^{+}20, FST24, GLS^{+}24, HSB^{+}09, JK23, LRD08, SS13]. Beyond [EKS^{+}22, MLL^{+}22, DMK^{+}22]. Bibliometric [Mün19]. bidirectional [SRWK13]. Big [ABG^{+}17, AMP^{+}17, FF22]. Bilaterally [SWZG23]. binning [MGM13]. BiographScope [ASY^{+}24]. Blend [YL24]. Blood [RTA22]. Bluetooth [CAMR22]. Board [MM21]. Bone [YLFM22]. Book [CLP24, BS12]. Books [LMWP24, NVCD22]. Botanical [CDV21]. BP [DCC^{+}20]. Brazilian [GSB18]. BRDF [PAZ^{+}23]. breath [HLYH14]. Bring [LCD^{+}15]. Bringing [CGL^{+}22]. Brokers [HKvF22]. Bronze [NCP^{+}24]. broom [WWS^{+}22]. Browsing [AMP^{+}17]. Brushstrokes [Geo24]. Bryn [EEG^{+}21]. Buddhist [LLS^{+}12]. Building [AAÇ22, CCN^{+}24, DMS23, FL20, GAKL24, GCM^{+}20, OMRP22, SFB^{+}22, PT19]. Buildings [BA24, Che15, FBP20, HSMAK^{+}22, SBL^{+}13]. Built [AGGD^{+}23, MMGGA21].

CAD [LBC^{+}15]. CAM [BA24]. camera [LLS^{+}12, MSN^{+}13]. Cameras [CTRG19, GSB18]. Campaigns [MDCS19]. Can [BVLW22, LMWP24]. Cantonese [LBM^{+}12, LBC^{+}15]. Capacity [SFB^{+}22]. Capitalization [AKK^{+}24]. Capture [MGBM14]. Capturing [AASFB23, FL20]. CARE [CCF20]. CARE-GIS [CCF20]. Carvings [WHC^{+}19]. Case [ALBA13, BSO22, BSDG^{+}21, Cak22, CWR^{+}23, CCD^{+}21, EEG^{+}21, Eic21, EMK^{+}21, FMG^{+}21, FCL^{+}24, FHA^{+}21, GK19, GWZ^{+}24, HDZM16, LLB24, NCP^{+}24, OMRP22, PPBH21, RDH15, RSM23, RTA22, SP16, SBC^{+}22, SKD19, UY22, WRKK^{+}22, ZDC21, CCD^{+}11, PVPMP24, CMN13]. Castle [MM16]. Casual [MNL^{+}21]. CAT [FPS^{+}21]. Cathedral [BZ24, PPBH21]. Cathedrals [BZ24]. Catherine [CCD^{+}21]. CAVE [KTHP15]. Caves [LLS^{+}12]. celebrated [RPA^{+}10]. Celli [EEG^{+}21]. Celsus [Cak22]. Cemeteries [Eic21]. CemoMemo [MKZ21]. CENDARI [BBB^{+}18]. Center [FMG^{+}21]. Centered [RDH15]. Central [CLP24, MKM21]. Centred [PLP^{+}17]. Century [GLF^{+}20, LBM^{+}12, Mim12, LBC^{+}15]. Ceramic [DDGP20, ZJCN24]. cesspit [SHFS12]. CH [CS08]. Challenge [Sto22]. Challenges [DAD^{+}22, SGH^{+}21, WHMW23, WD13, KFH09]. Change [AVK^{+}22, Ulv22, HLZ12]. Chapel [MM16]. Character [CGG^{+}18]. Characteristics [Geo24]. Characterization [CAMR22, GWZ^{+}24]. Characterizing [CNS15]. CHER [WSA^{+}18]. CHER-Ob [WSA^{+}18]. Chess [LBC^{+}15, LBM^{+}12]. Children [OAAA21, TAR20]. Chinaware [NWLZ23]. Chinese [APHGG23, Che14, HZ24, LXW24, LW16, LLL^{+}24, MWX^{+}12, NJDZ11, SZF^{+}10, TWZF24, WKNC14, XLL^{+}24, ZKB^{+}20, ZM20]. Chopin [Dry24]. Choreography [Por21]. Christ [RTA22]. Christian [PFF^{+}15]. Christianity [CCF20]. Chronological [CCZ^{+}17]. Church [Erk22, SOÇT23]. CIDOC [CFdS22, FKD23, RTF22, TF23]. CIDOC-CRM [FKD23]. Cinema [ASY^{+}24]. Cinematic [TLL^{+}24]. Cinematographic [BCF^{+}21, SPC^{+}23]. circa [RPA^{+}10]. Cisneros [HDZM16]. Cited [BPLP23]. Cities [FST24]. Citizen [ADD^{+}24, CDM^{+}24, DAD^{+}22, JDDAB^{+}24, TJTHW24]. City [SBN22, PPGLPMR24]. Class [WM24]. Classic [LLL^{+}24]. Classical [EMK^{+}21, HZ24, ZM20, MCG11]. classics [Mim12]. Classification [AFG15, BA24, COGP16, CCZ^{+}17, FUD^{+}19, Fra22, Geo24, MF21, NDN^{+}21, RFW^{+}22, RRJBMM16, TJTHW24, WSHH24, DT08, MD12]. Classifying [VKS^{+}24]. Cleaning [MDCS19]. CLEF [DWD^{+}23]. Close [SDH^{+}16]. Close-Range [SDH^{+}16]. Cloud [CSS^{+}17, LXW24, Nic19a, RAB^{+}19, UY22, ZDC21]. Cloud-Based [CSS^{+}17]. Clouds [HWSK14, DKC23]. Clustering [KCJP24]. Clusters [JDDAB^{+}24]. Cluttered [BPV^{+}15]. CNN [KY24]. CNNs [COGP18]. Co [HHJ^{+}24]. Co-Creation [HHJ^{+}24]. Code [FKS22]. Code-mixing [FKS22]. Codes [BDK23]. Codex [SBC^{+}22]. coding [HLZ12]. COFLA [KDBMG16]. Cognitive [RFA19]. Cohesion [LSA^{+}24]. Coins [FHA^{+}21, KZK14]. Collaboration [BSW^{+}22, DH21, FL20]. Collaborative [ACF^{+}18, FPS^{+}21, NPR^{+}20]. Collection [GBG^{+}22, ORD^{+}21, TV23, UY22, VC21, CDV21]. Collections [ASC20, APH24, CSSC^{+}24, ES16, ESW20, FKW17, GFE19, HVS^{+}15, MNL^{+}21, OMRP22, PAZ^{+}23, RRDC23, RAB^{+}16, SU22, UA20]. Collective [AGGD^{+}23, CC23]. College [OZY^{+}21]. Cologne [SHFS12]. Colombia [HDZM16]. Color [Ben23, BPV^{+}15, EES^{+}20, GCL^{+}19, HWSK14, MSTG18, DCC^{+}10]. Color-Based [HWSK14]. Colorization [LKK^{+}20]. Column [SP16]. Columns [LLD15]. Combination [CCZ^{+}17, LBC^{+}15]. Combining [LBM^{+}12]. Comedy [AD21]. Comes [AVK^{+}22]. Commerce [CFRCDSC21]. Commodity [JZF23]. Communicate [AFGG21]. Communicating [CCFC23, GFE19]. Communication [CFRCDSC21, FWZ23, FMMM22, GFE19, TF23]. communications [KSZ^{+}11]. Communities [ESD^{+}22]. Comparative [GLS^{+}24, JML24, RTA22, YH13].

Comparing [KF23, PFZG23, ZJCN24]. Comparison [DCC^{+}20, UY22, VC21]. Complete [PSA^{+}17]. Completion [PSA^{+}17]. complex [RGRG09]. Compliant [CFdS22]. Components [SOÇT23]. Comprehensive [DRH19, OAB^{+}24, SBvE^{+}23]. Compression [vBTS^{+}22]. Computation [Sto22]. Computational [AVV21, CMST22, DS21, GB22, GBAA22, HMG22, KDBMG16, Mim12, PS22, PMHdF24, dMKBD22, CL12]. Computed [LBC^{+}15]. Computer [BEES15, CGP15, CGL^{+}22, DKY24, FUD^{+}19, RSM23, ST12, VC21, WMJ^{+}18, Arn08b]. Computer-Aided [DKY24, CGL^{+}22]. Computer-Assisted [WMJ^{+}18, RSM23]. Computerized [MM21]. Computing [ASC12]. concepts [FMAG13]. Conceptual [IA18]. Concerts [EMK^{+}21]. Concurrent [BCFS15]. Conditions [FHA^{+}21]. Conducting [MR15]. Conforming [CFC23]. Connected [SH22]. Connecting [RRDC23]. Conservation [BA24, HSMAK^{+}22, MCKC24, NMM^{+}17, RAD^{+}23, HS08]. Conservation-Restoration [NMM^{+}17]. Considering [PT22]. Construct [AKK^{+}24, BSDG^{+}21]. Construction [FL20, LDS16, ZLSD13]. Constructivist [GFE19]. consultation [SBL^{+}13]. Contact [APHGG23]. Contactless [JZF23]. Contemporary [BTIR^{+}23, IRHH^{+}22, Por21]. Content [AASFB23, BGB24, JK22, KSSW24, LCD^{+}15, PMGZ^{+}24, SLRV23, HGCS11, VC11]. Context [AMP^{+}17, MC22, NCP^{+}24, NKRS15, PDK^{+}22, Por21, SLRV23]. Context-awaRe [AMP^{+}17]. Contexts [CFC23, VRC24]. Contextualizing [MHJ22]. Continuity [WLZ19]. Contour [NWLZ23]. Control [HRS^{+}16]. Convergence [BSW^{+}22]. Conversational [CPR^{+}24]. Convert [CFC23]. Converter [CFC23]. Convolutional [AFV16, MF21]. core [DKB11]. Coronavirus [PT22]. Corpora [FKS22]. Corpus [KDBMG16, CL12, HLZ12, ZLSD13, KDBMG16]. Correction [SZGE22, WWS^{+}22]. Corrupted [MYS22]. Cost [ZSV^{+}16]. Counting [JZF23]. Coupling [BGP22]. Covid [GB22, BKPW24, MLL^{+}22]. COVID-19 [BKPW24, MLL^{+}22]. CPF [CFC23]. CR [HBC24]. CR-TransR [HBC24]. Craft [ZMD^{+}22]. Crafts [NVN^{+}22, SPZ^{+}22]. Create [BPLP23]. Creating [BCF^{+}21, BCH^{+}15, DMK^{+}22, HLYH14, KA20, PLC^{+}20, Tro22]. Creation [EEG^{+}21, HHJ^{+}24, HH13]. Creative [ACH24, ESD^{+}22]. Creativity [BSW^{+}22]. Criteria [NTM^{+}23]. Critical [GAKL24, LRS^{+}21]. CRM [CFdS22, FKD23, TF23]. CRM-Compliant [CFdS22]. CRMba [RTF22]. Cross [CDV21, LCD^{+}15, LDS16, MANC24, ZKB^{+}20, ZLSD13]. Cross-collection [CDV21]. Cross-cultural [ZKB^{+}20]. Cross-lingual [MANC24]. Cross-Period [LDS16]. Crowd [MKZ21, MBEB23]. Crowd-sensing [MBEB23]. Crowdsourcing [AGGD^{+}23, CC23, DWD^{+}23]. CT [LBM^{+}12]. Cube [ESKI21]. Cultural [APHGG23, ACH24, AMP^{+}17, ACF^{+}18, BA24, BV17, BZ24, BED^{+}23, BPF^{+}18, BFJ^{+}23, BGP22, CDM^{+}24, CD18, CKNK18, CC23, CWR^{+}23, CLCL20, CCFC23, CCD^{+}21, CCAS23, DAD^{+}22, DDDK^{+}24, DMP^{+}17, DH21, DNR17a, DNR17b, DBYA18, DMS23, DTM17, DSP24, ESD^{+}22, ERG^{+}19, EEG^{+}21, FCWY21, FMG^{+}21, FMMM22, GK19, GSB18, GBG^{+}22, HLZL18, HMP17, HSMAK^{+}22, HDZM16, HKP^{+}22, HBC24, Hu24, HGLS23, IAY15, IA18, JDDAB^{+}24, JTL19, KT16, KC20, KKK^{+}20, KKW22, KOIC20, KF23, LhL19, LPD16, LSA^{+}24, MM23, MRAB21, MRB^{+}14, MRFF24, MC22, MM21, MCKC24, NCP^{+}24, NKRS15, ORD^{+}21, OMRP22, OAAA21, OAB^{+}24, PMGZ^{+}24, PPF^{+}23, PGMBR17, PWM^{+}17, PT22, PMG^{+}20, PT19, RFB^{+}19, RFA19, RDH15, RRL^{+}14, RKG^{+}24, SSLdB24, SBO^{+}17, SRH^{+}20, SGH^{+}21, SLRV23, SM18, SCM^{+}24, SFB^{+}22, SKD19, SOÇT23, Tai23, TWZF24]. Cultural [TAR20, UY22, VRC24, VMH^{+}18, WZL^{+}21, WHMW23, WWS^{+}22, ZOZ21, dMKBD22, ASC12, ABV11, ALBA13, BBD^{+}12, Che14, CMN13, FPMT10, HH13, HGP12, HES09, KRF^{+}14, KFH09, LRD08, LDD08, MB13, PZF11, RGRG09, STLL13, SBL^{+}13, SS13, ZKB^{+}20, Arn08b]. Culture [FWZ23, GB22, KC20, LLL21, MPRS17, WKNC14]. Cultures [NDN^{+}21]. Cuneiform [BM22, RFW^{+}22]. Curated [SSLdB24]. Curating [KSSW24]. Curation [ADD^{+}24, DMK^{+}22, FPS^{+}21, JK22]. Curators [BGP22]. Current [CBJN22]. Cynic [JTL19]. Cypro [PHB22]. Cyprus [APH24, PPBH21]. CyprusArk [APH24]. Czechia [Eic21].

D

[DNS23, DSH^{+}15, BPV^{+}15, BCFS15, CNS15, CMGFOAFH23, CC23, CS08, DCC^{+}10, ES16, ESW20, EES^{+}20, GMLLHFTT21, GSB18, GBG^{+}22, GA16, HSB^{+}09, HRS^{+}16, HSMAK^{+}22, HWSK14, KRF^{+}14, KFH09, KKRS22, LBM^{+}12, LBC^{+}15, LCD^{+}15, LTP21, MSWS20, MWL^{+}16, MKM21, NTM^{+}23, PG15, PHB22, RGRG09, RRC^{+}10, RAB^{+}16, RTA22, Sap19, SB22, SM18, YH13, ZA24, ZPS^{+}16, ZSV^{+}16, vBTS^{+}22]. D-Reconstructed [MWL^{+}16]. Dacian [NCP^{+}24]. Damage [DCC^{+}20, MBEB23]. Damaged [CLN18, DS21, SP16]. Dance [ASC^{+}15, ASC20, PP23, Por21, RKWC24, MCG11]. Dare [IEL21]. Data [ABG^{+}17, AMP^{+}17, CDM^{+}24, CCF20, DMP^{+}17, DWD^{+}23, DH21, DS21, FPS^{+}21, FL20, FF22, GFZ24, GNL20, GGN22, HSMAK^{+}22, HLT23, JK23, KCRE23, MGB18, MQB22, MWL^{+}16, Mim12, MPFC16, Nic19a, ORD^{+}21, Pay23, RAB^{+}19, SPC^{+}23, UY22, ATP11, RGRG09]. Data-Driven [GNL20, DS21, FL20]. Database [ORD^{+}21, SPC^{+}23]. Databases [CFdS22]. Dataset [HZ24, LA24, MF21, PP23, SV24, WHC^{+}19, YLFM22, WZAM^{+}21]. Datasets [CFC23, CMST22]. David [SE17]. Daxi [CWR^{+}23]. Daylight [GNL20]. Dazu [WHC^{+}19]. Ddu [EEG^{+}21]. Debates [CBJN22]. DECHO [ABV11]. Decision [MCKC24]. Decision-Making [MCKC24]. Decisions [RFA19]. Deconstructed [CCN^{+}24]. Decorative [ES16]. Deep [AVV21, BA24, DSH^{+}15, GAKL24, KOIC20, MBEB23, MYS22, PLSL22, PP23, RFW^{+}22, SZGE23, TJTHW24, WM24, YLFM22, ZLS^{+}22]. Deep-Sea [DSH^{+}15]. Degraded [HOGP17]. Degree [ZM20]. Delaminated [CLN18]. Deliberative [BSO22]. Dellepiane [Nic19b]. Depth [FHA^{+}21, ZPS^{+}16, vBTS^{+}22]. Depth-of-Field [vBTS^{+}22]. described [DG11]. Describing [MQD^{+}23]. Description [PLSL22]. Design [AMP^{+}17, ACF^{+}18, BSW^{+}22, BSDG^{+}21, BFJ^{+}23, BBC21, CLP24, CRR^{+}14, Dim22, DKY24, HDZM16, NB17, NCP^{+}24, NVN^{+}22, PLP^{+}17, STLL13, XLL^{+}24, YL24, YK21]. Designed [EMK^{+}21]. Designer [RFA19]. Designing [EKS^{+}22, GGN22, IA18, KRJ23, LRS^{+}21, Mas16, REBK^{+}16, YL24]. designs [CKA13]. Destruction [Pay23]. Detail [RRL^{+}14]. Detailed [RAB^{+}16, RGRG09]. Detect [MRFF24]. Detection [APHGG23, KIM24, KML22, KOIC20, MANC24, SWZG23, TWZF24, TG15, Ulv22, HLYH14, SZMB11]. deterioration [HLYH14]. determination [RPA^{+}10]. Develop [ACF^{+}18]. Developing [HRS^{+}16, KML22]. Development [HDZM16, KKK^{+}20, PLC^{+}20, ALBA13, HLB14]. Device [DMPC14]. Devices [REBK^{+}16]. Diachronic [CSS^{+}17].

Diagnostic [SCCF22]. Dialogical [MLL^{+}22]. dialogue [MBD^{+}13]. Diars [AKK^{+}24]. Did [PT19]. Differences [CC23, GLS^{+}24]. Different [ADN21, FHA^{+}21, NDN^{+}21, PT19]. Dig [BVLW22]. Digging [PWM^{+}17]. Digital [ABG^{+}17, ACH24, AMP^{+}17, AVK^{+}22, ASC20, BSW^{+}22, BM22, BTIR^{+}23, CNS15, CCF20, CCAS23, CFdS22, DG11, DNR17a, DNR17b, DTM17, ERG^{+}19, Eic21, FPS^{+}21, FF23, FTM19, GK19, GLF^{+}20, HHJ^{+}24, KPV^{+}20, KHO^{+}23, LKK^{+}20, LCD^{+}15, LXW24, LLL^{+}24, LWZ^{+}24, MM23, MPRS17, MNL^{+}21, MLL^{+}22, Mün19, NTM^{+}23, NCM^{+}19, OTBC11, PVPMP24, PMGZ^{+}24, PPF^{+}23, PGMBR17, PWM^{+}17, RJIH22, RDH15, RAB^{+}16, SV24, SBO^{+}17, SLRV23, SH22, SFB^{+}22, SIS^{+}23, SGMT21, SKD19, Tai23, TFK^{+}09, TV23, WHMW23, ZKB^{+}20, ZLS^{+}22, ASC12, ABV11, DS09, KFH09, LBM^{+}12, LLS^{+}12, MSN^{+}13, SLD^{+}12, SS13, YAGU09, YH13, CFRCDSC21, MFRG10]. Digitalization [GMLLHFTT21]. Digitally [ATP11, WRKK^{+}22]. Digitisation [ZMD^{+}22]. Digitising [TJTHW24]. Digitization [BJC^{+}22, BPV^{+}15, CGG^{+}18, GA16, HVS^{+}15, Che14]. Digitized [APH24]. Digitizing [HKP^{+}22]. Dimension [HSMAK^{+}22]. Dimensional [SGH^{+}21, GBG^{+}22]. Diokaisareia [UY22]. directing [Yan13]. DirectorNotation [Yan13]. Disappeared [SOÇT23]. Disaster [KOIC20]. Disaster-Affected [KOIC20]. Disc [CLN18]. Discovering [PFF^{+}15]. Discovery [KCRE23]. Discrete [VVSB14, DT08]. Displacement [GA16]. Display [ZJCN24]. Displays [PGMBR17]. Dissemination [FWZ23, NCP^{+}24, WSA^{+}18]. Dissipation [AS23]. District [CLP24]. Diversity [CDV21, HZ24]. Divine [AD21]. Do [LLD15, RFA19]. Document [PLSL22, GBBS11]. Documentation [FBB^{+}24, GMLLHFTT21, HAK21, MM23].

Documenting [RTF22, YH13]. Documents [HOGP17]. does [EKS^{+}22]. Domain [BVLW22, DMP^{+}17, HBC24, NMM^{+}17, PFZG23, SH22]. Domes [MM23, SB22]. Domestic [AKK^{+}24]. Drama [LPD16]. Drawing [AFGG21, WMJ^{+}18, RPA^{+}10]. Drawings [SWZG23, HGCS11, KLST13]. Dream [LCD^{+}15, DI08]. Driven [GNL20, APHGG23, AKK^{+}24, DS21, FL20, GCM^{+}20]. Driving [RRC^{+}10]. Dropped [WWS^{+}22]. Drum [SP16]. Dual [KRJ23]. Dual-level [KRJ23]. Dublin [OZY^{+}21]. Dunhuang [ACH24, LLS^{+}12, WLZ19, ZLS^{+}22]. Dust [TG15]. Dutch [KHO^{+}23]. Dynamic [AAÇ22, PCC^{+}10, YLFM22, GMMC09]. Dynasties [DYW21]. Dynasty [FST24, LW16, YZH^{+}24].

E-Commerce [CFRCDSC21]. EAC [CFC23]. EAC-CPF [CFC23]. EAD [CFC23]. Early [ASY^{+}24, AS23, CGG^{+}18, TF23]. Early-stopping [AS23]. Earthen [DCC^{+}20, GWZ^{+}24]. Earthquake [FBP20]. easy [MBD^{+}13]. Eco [AKK^{+}24]. Eco-Diars [AKK^{+}24]. Ecologies [SU22]. Economy [PPGLPMR24]. Edge [YZH^{+}24]. Edition [LCD^{+}15]. Editions [FTM19]. Editorial [ADP15, Arn08a, Arn08b, BFJ^{+}23, CD18, DNR17a, ERG^{+}19, Nic18, Nic21, Sco12, Sco13]. Editors [DNR17b]. Education [EEG^{+}21, GB22, MM21, PPF^{+}23]. Educational [CLP24, HGLS23, TJTHW24]. Effect [AY23, FWZ23]. Effective [BPV^{+}15, PAZ^{+}23]. Effectiveness [FMMM22]. Effects [CLCL20, MPPB24]. Efficient [BSO22, NDN^{+}21, RKG^{+}24]. Egocentric [RFB^{+}19]. Egypt [CPR^{+}24, HES09, PMG^{+}20]. Egyptian [HAK21]. eHeritage [Fri11]. electronics [BB08]. Elementary [TAR20]. Elicitation [VMH^{+}18]. Embedding [HBC24, TMB^{+}20]. Embodied [DSP24, HKP^{+}22, KCS14, VC11]. Emergency [FBB^{+}24]. Emergent [HKvF22]. Emotion [KY24, LA24]. Emotional [VMH^{+}18]. Empathy [LSA^{+}24]. Empire [KA20]. Empirical [APH24]. emulation [BB08]. Enabling [VRC24, CS08]. Encoding [MQD^{+}23]. Encyclopedia [PWM^{+}17]. End [IAY15]. Energy [CAMR22]. Engagement [DDDK^{+}24, ESW20, MLL^{+}22]. Engaging [CN23, ESD^{+}22]. Engineering [DMK^{+}22]. Engraved [GLF^{+}20]. Enhance [ESW20, FYYG22]. Enhanced [PMG^{+}20, ZPS^{+}16, HZ24]. Enhancement [BJC^{+}22, PCC^{+}10]. Enhancing [APH24, CSSC^{+}24, DMP^{+}17, EMK^{+}23, HGLS23, LSA^{+}24, MVCK^{+}23, MHJ22]. Enrich [EMK^{+}21]. enriched [SBL^{+}13]. Enriching [KKW22]. Ensemble [HZ24]. Enterprises [LC21]. Enthusiast [JTL19]. Entity [BVLW22, PFZG23, TWZF24]. Entity-Relation [TWZF24]. Entry [FPS^{+}21]. Environment [ASY^{+}24, BCF^{+}21, DYW21, Hu24, HGLS23, IA18, KTHP15]. Environments [CWR^{+}23, CLCL20, IAY15, AMP^{+}17]. Ephemera [SU22]. Ephesos [TFK^{+}09]. Ephesus [Cak22]. Epigraphic [SDH^{+}16]. Epistolary [HLT23]. era [VKS^{+}24]. Error [SZGE22]. Established [AFGG21]. estimate [GBBS11]. Estimation [MVCK^{+}23, SV24]. Ethnography [ASC20, MA12]. Etruscanning [PA14]. Eurographics [CD18]. Europe [MANC24]. European [BBF^{+}17, TF23]. Evaluate [ACF^{+}18]. Evaluating [COGP16, GK19, KY24, NCP^{+}24, PPF^{+}23, PT19, RBXM15]. Evaluation [ACH24, ASC^{+}15, APH24, BCFS15, BEES15, Cak22, CFdS22, DRH19, FWZ23, FF22, KRJ23, RFA19, TSB21, XLL^{+}24, YK21, HGP12, ERG^{+}19]. Event [FCWY21, RNP21]. Event-based [RNP21].

Events [FL20]. Evidences [AASFB23]. Evolution [DYW21, FWZ23]. evolving [HLZ12]. examining [KRF^{+}14]. Example [FWZ23, PHB22]. Excavation [HAK21, SHFS12, Ulv22, ZSV^{+}16]. Excavation-parallel [SHFS12]. exercise [CS08]. exhibit [GBBS11]. Exhibition [BSW^{+}22, LLA^{+}19, SSLdB24, MWX^{+}12]. Exhibitions [SSLdB24]. Expectations [CLCL20]. Experience [BBD^{+}24, CLCL20, CGP15, CRR^{+}14, EEG^{+}21, ESKI21, EKS^{+}22, EMK^{+}23, EMK^{+}21, GLS^{+}24, HGLS23, JK22, JTL19, KMMK22, Mas16, NCP^{+}24, NVN^{+}22, OAAA21, OAB^{+}24, Por21, Tai23, TAR20, VQP^{+}24, YL24, YK21, HLYH14]. Experiences [ACH24, AASFB23, CPR^{+}24, CDM^{+}24, DAD^{+}22, DBYA18, Dim22, Dry24, ESD^{+}22, KPV^{+}20, KKW22, LC21, LRS^{+}21, PLP^{+}17, PT19, REBK^{+}16]. Experiences-Enriching [KKW22]. Experiential [AASFB23]. Experimental [BSW^{+}22, GLS^{+}24, WD13]. Expert [MGBM14]. Explainable [SCM^{+}24]. Exploiting [PAZ^{+}23]. Exploration [GFZ24, JVAP^{+}21, MRB^{+}14, Mas16, OKY15, RRL^{+}14, SLRV23, UA20, XLL^{+}24, ABV11]. Exploratory [EMK^{+}21, Che14]. Explore [BPLP23, CDV21, SBN22, CKA13]. Explorer [YAGU09]. Exploring [ASY^{+}24, AS23, BV17, BBD^{+}24, GSB18, HZ24, KPV^{+}20, LDD08, LMWZ17, Por21, PTP23, RAB^{+}16, SSLdB24]. Exposing [DMP^{+}17, SU22]. Expression [KRJ23, VMH^{+}18]. Expressivity [DMP^{+}17]. Extended [DDDK^{+}24, NCP^{+}24, VRC24]. Extending [GFE19]. Extension [FKD23]. Extracting [BS12, HOGP17]. Extraction [JK23, MCV23, PYR15, PSBM15]. Extrinsic [MRAB21]. Eyck [CDV21]. Eye [PGMBR17, ZJCN24].

Facade [AVV21]. Facets [APHGG23].

Facial [KRJ23]. Facilitate [LMWP24]. Facilitating [IAY15]. Factors [IAY15]. Factual [DKB11]. FakeMuse [ZOZ21]. Families [REBK^{+}16]. family [DS09]. Fast [AFG15, MGM13, PG15, FPS^{+}21]. FE [DS21]. FEA [ZDC21]. Feature [DDGP20, CCZ^{+}17, MPPB24, NWLZ23]. Feature-based [MPPB24]. Features [KRJ23, KCJP24, PSBM15, WSHH24]. Feedback [CGL^{+}21, KKW22]. FESRCT [FCWY21]. Fidelity [CPR^{+}24]. Field [Mün19, RSM23, vBTS^{+}22, LLS^{+}12]. Field-Topics [Mün19]. Files [CFC23]. Film [SPC^{+}23, TG15, SZMB11, Yan13]. Finding [HLZL18]. Finds [DDGP20]. Fine [AFGG21, IA14]. Fingerprint [SGH^{+}21]. FiRe [SPC^{+}23]. First [BCF^{+}21, CGP15]. First-person [BCF^{+}21]. Fisher [AFV16]. fitting [USF08]. Flaking [GWZ^{+}24]. Flamenco [KDBMG16]. flash [DCC^{+}10, MSN^{+}13]. flash-based [DCC^{+}10]. Flashlight [RRL^{+}14]. Fledged [PMGZ^{+}24]. Floor [KML22]. flow [HES09]. Fluorescence [DLI^{+}17]. Flute [PMG^{+}20]. FOIA [BSO22]. Folk [ASC^{+}15]. folksonomy [SLD^{+}12]. folksonomy-based [SLD^{+}12]. Food [MPRS17]. Formalism [NB17]. format [LLS^{+}12]. Former [CCN^{+}24]. Forms [ZJCN24]. Four [SS13]. Fracture [NWLZ23, SDF^{+}12]. Fractured [WZL^{+}21]. Fracturing [ESW20]. Fragile [SBVVSL^{+}15]. Fragments [ASBD21, NWLZ23, VVSB14]. Frames [WWS^{+}22]. Framework [ACF^{+}18, BBC21, CKNK18, DRH19, Dim22, FCWY21, HDZM16, HLT23, IA18, MCKC24, NWLZ23, PG15, dMKBD22, ABV11, HH13]. Frederic [Dry24]. free [AS23, SWZG23]. frequency [HS08]. Friendly [RKG^{+}24]. Full [PMGZ^{+}24]. Full-Fledged [PMGZ^{+}24]. Fully [NKRS15]. functions [SRWK13]. Funeral [RSM23]. Future [CBJN22]. Futures [KCS14]. futurism [DS09].

Galatina [CCD^{+}21]. Galgano [PZF11]. Game [BBD^{+}24, BSDG^{+}21, BCF^{+}21, CN23, EMK^{+}23, EMK^{+}21, FMG^{+}21, MM21, OAAA21, RFA19, RKWC24, RBXM15, SGMT21, XLL^{+}24, ZOZ21, ALBA13, BBD^{+}12, CMN13, HH13]. Game-Based [CN23, OAAA21]. GameBot [PMHdF24]. Games [ACF^{+}18, DFSO21, JTL19, LLL21, LC21, LRS^{+}21, RFA19, MBD^{+}13, MB13, WD13]. Gamification [BSDG^{+}21, CGL^{+}21, NPR^{+}20]. Gaming [SS13, VRC24]. Garowe [FBB^{+}24]. Gastronomic [MPRS17]. Gateway [PWM^{+}17, GMMC09]. Gaze [MRFF24]. GDOM [KMMK22]. Genealogical [SZGE23]. General [KHO^{+}23]. Generalization [TWZF24]. Generating [KB18]. Generation [CGL^{+}22, LW16, NB17, PLSL22, PP23, ZLSD13]. Generative [LMWP24, LWZ^{+}24, PKO23]. Genetic [SF18]. Geographical [BPLP23]. Geometric [Haj21, KRJ23, NB17]. Geometrical [Nic19a, RAB^{+}19, DT08]. Geometry [NKRS15, PG15, SWZG23]. Georgian [RSM23]. German [HHJ^{+}24]. Germany [SHFS12]. Gestural [VMH^{+}18]. Gesture [ZKB^{+}20]. Gestures [MGBM14, PP23]. Ghent [CDV21]. GIS [CCF20, HSMAK^{+}22, SBN22]. GIS-based [SBN22]. Giza [HES09]. Glass [Dim22, GNL20]. Glasses [Mas16]. Glassware [Mas16]. Global [Por21, DI08]. GLOBE [SZGE23]. Gloss [EES^{+}20]. Glyph [COGP16]. Glyphs [COGP18, HOGP17, MYS22]. Goose [LXW24]. Gothic [BZ24]. Government [BSO22]. Grad [BA24]. Grad-CAM [BA24]. Gradient [ZPS^{+}16]. Gradual [SZMB11]. Graffiti [TF23]. Grammar [NB17]. Grams [KT16]. Grand [Sto22]. Graph [BGP22, CFdS22, DMK^{+}22, HBC24, ORD^{+}21]. Graphics [CD18, KRF^{+}14, AFGG21]. Graphs [DMS23, DKC23, SZGE23]. Gravestones [MKZ21]. Great [DCC^{+}20, LXW24]. Greater [PLSL22]. Ground [PLSL22]. grounded [MVCK^{+}23]. Group [KPV^{+}20]. groups [DT08]. Guest [DNR17b]. Guide [ADN21, KKK^{+}20, KSZ^{+}11]. Guided [KPV^{+}20, ZLS^{+}22]. Guides [AVK^{+}22]. guqin [NJDZ11, SZF^{+}10].

Hagia [Erk22]. Hand [PP23]. Handheld [MM16]. Handling [MPRS17]. Handmade [MPRS17]. Handwriting [NJDZ11]. Handwritten [PYR15]. Haptic [CGL^{+}21, KKW22]. Haptics [CPR^{+}24]. Harmonization [KCRE23]. Harnessing [CC23]. Hawamahal [KKRS22]. hBIM [Nic19a, SCCF22, ZDC21, RAB^{+}19]. hBIM-Oriented [Nic19a, RAB^{+}19]. Hearing [EKS^{+}22]. Hebrew [LLB24, SZGE22]. HeGO [FMG^{+}21]. Height [PYR15]. Help [MKZ21]. Hemispherical [MM23]. Herbaria [MCV23]. Heritage [APHGG23, ACH24, AMP^{+}17, ACF^{+}18, AGGD^{+}23, BA24, BV17, BZ24, BJC^{+}22, BED^{+}23, BPF^{+}18, BFJ^{+}23, BGB24, BBC21, CDM^{+}24, CSSC^{+}24, CKNK18, CC23, CLP24, CWR^{+}23, CLCL20, CCFC23, CCD^{+}21, CCAS23, CCN^{+}24, DAD^{+}22, DMP^{+}17, DH21, DNR17a, DNR17b, Dim22, DMS23, DTM17, DSP24, ES16, ESW20, ESD^{+}22, EEG^{+}21, FKW17, FMG^{+}21, FCL^{+}24, FMMM22, GK19, GMLLHFTT21, GAKL24, GSB18, GBG^{+}22, GLS^{+}24, HLZL18, HMP17, HSMAK^{+}22, HDZM16, HKP^{+}22, Hu24, HGLS23, IAY15, IA18, JK22, JK23, JDDAB^{+}24, JTL19, KMMK22, KC20, KKK^{+}20, KKW22, KOIC20, KKRS22, KF23, LhL19, LPD16, LC21, LSA^{+}24, MM23, MRAB21, MRB^{+}14, MBEB23, MMGGA21, MC22, MWL^{+}16, MM21, MCKC24, Mün19, NCP^{+}24, NVN^{+}22, NPR^{+}20, NKRS15, ORD^{+}21, OMRP22, OAAA21, OAB^{+}24, PMGZ^{+}24, PPF^{+}23, PGMBR17, PWM^{+}17, PT22, PPBH21]. Heritage [PMG^{+}20, RFA19, RDH15, RRL^{+}14, RKG^{+}24, SBO^{+}17, SRH^{+}20, SGH^{+}21, SLRV23, SM18, SCM^{+}24, SKK^{+}20, SBN22, SFB^{+}22, SKD19, SOÇT23, Tai23, TWZF24, TAR20, UY22, Vio23, VMH^{+}18, WHMW23, WWS^{+}22, XLL^{+}24, ZA24, ZOZ21, dMKBD22, ASC12, ABV11, ALBA13, Arn08b, BBD^{+}12, Che14, CMN13, CL12, DS09, FPMT10, GSGC08, HH13, HGP12, HES09, KFH09, LDD08, MCG11, MB13, PZF11, RGRG09, STLL13, SBL^{+}13, SS13, WD13, CD18]. HeritageSite [XLL^{+}24]. Heterogeneous [BCFS15, SZGE23]. Heuristic [PFZG23]. Heuristics [AS23]. Hidden [LTP21]. Hierarchical [LXW24]. Hieroglyph [PBCOMTRR19]. High [CPR^{+}24, GMMC09, WWS^{+}22]. High-Fidelity [CPR^{+}24]. High-resolution [WWS^{+}22]. Hill [ATP11]. Hispanic [BZ24]. Histograms [RRJBMM16]. Historic [AAÇ22, TS14, YL24, SZMB11]. Historical [CGL^{+}22, DYW21, DLI^{+}17, FST24, FL20, FMG^{+}21, GCM^{+}20, GGN22, Haj21, KIM24, LLB24, MSWS20, MPFC16, NTM^{+}23, NCM^{+}19, PLSL22, RAD^{+}23, RNP21, SBN22, SZGE22, VYBG22, ZA24, Che14, PDK^{+}22, HLT23]. Histories [SH22]. historiography [Mim12]. History [BG14, DMK^{+}22, FKD23, FL20, GLF^{+}20, HVS^{+}15, Kim22, LTP21, LC21, PS22, PLC^{+}20, RCB^{+}18, RRDC23, SV24, SH22, FMAG13, LRD08]. HMD [JML24]. HMD-based [JML24]. Holistic [RDH15]. Hollow [NCP^{+}24]. Holocaust [BBF^{+}17]. Holographic [PGMBR17]. Holy [MRAB21, RTA22]. Holy-Land [MRAB21]. Homogenization [DS21]. Hong [CLP24]. Horizonte [CCN^{+}24]. House [AVV21, ATP11, YL24]. Houses [KA20]. HTR [SBC^{+}22]. Hub [CDM^{+}24]. Human [AVK^{+}22, MVCK^{+}23, RFA19, SV24]. Humanities [FPS^{+}21, FTM19, KML22, LLL^{+}24, LWZ^{+}24, RJIH22, SLRV23]. humanity [CL12]. Hybrid [LRS^{+}21, NVN^{+}22, OAB^{+}24, TSB21, WSHH24]. Hyper [GBG^{+}22, KRF^{+}14]. Hyper-dimensional [GBG^{+}22]. Hyperspectral [WWS^{+}22]. Hypotheses [KTHP15].

ICON [SBvE^{+}23]. Iconclass [BDK23].

Iconographic [BGB24, SBL^{+}13]. Iconography [MF21]. Identification [MBEB23, SCM^{+}24]. Identify [KT16]. Identifying [BA24, CYH^{+}22, VKS^{+}24]. IFLA [GFZ24]. Illumination [RFW^{+}22]. Illumination-based [RFW^{+}22]. Illuminations [CRR^{+}14]. Image [AVV21, AFG15, AGGD^{+}23, BA24, FF23, FHA^{+}21, GCL^{+}19, HOGP17, JVAP^{+}21, OKY15, PBCOMTRR19, PAZ^{+}23, RAD^{+}23, RPA^{+}10, TSB21, YZH^{+}24, vBTS^{+}22, DRAT11]. Image-Based [OKY15]. Imagery [CDV21, DLI^{+}17, RTA22]. Images [AVV21, KIM24, KOIC20, MM23, RKG^{+}24, SKK^{+}20, Sto22, WWS^{+}22, IA14]. Imaging [HRS^{+}16, KF23, LLS^{+}12, PCC^{+}10]. ImaNote [DRAT11]. Immersive [ASY^{+}24, ACF^{+}18, BSW^{+}22, FCL^{+}24, JML24, KMMK22, KKRS22, RDH15, ZM20, MWX^{+}12]. Immersivity [LSA^{+}24]. Impact [BBD^{+}24]. Implementation [AMP^{+}17, PMGZ^{+}24, PPGLPMR24]. Implementations [GAKL24]. Impressionist [Geo24]. Improve [BSO22, FMMM22, TAR20]. Improved [DCC^{+}10, PBCOMTRR19]. Improving [BGP22, OAAA21, PLP^{+}17, SRH^{+}20]. in-place [MGM13]. inaugural [Arn08b]. Incomplete [LBC^{+}15]. Index [BPLP23]. indexing [DT08]. Indian [DG11, MCG11, PP23]. Indigenous [TLL^{+}24, HH13]. Individually [Hu24]. Induced [DLI^{+}17, Erk22]. Industrial [HVS^{+}15, LC21]. INF [CAMR22].

INF-Based [CAMR22]. Infilling [PKO23]. Influence [KT16, PTP23]. Information [BG14, BGP22, BBC21, BEES15, BVLW22, CMGFOAFH23, DPG22, FKD23, FWZ23, GFE19, GAKL24, HMS20, HKvF22, LWZ^{+}24, MQB22, MNL^{+}21, MQD^{+}23, MANC24, YN20, HGP12, LRD08, STLL13]. INFRA [CCAS23]. INFRA-ART [CCAS23]. Infrared [GCL^{+}19]. Infrastructure [ABG^{+}17, BBB^{+}18, FKW17, LMWZ17, MSR^{+}17, BBF^{+}17]. Infrastructures [DNR17a, DNR17b]. inhabiting [WD13]. Innovation [HMP17]. Innovative [ACH24, CSS^{+}17, PPGLPMR24]. Inpainting [SM18, WLZ19, WHC^{+}19, ZLS^{+}22]. input [NJDZ11]. Inquiry [TV23]. Inquiry-Based [TV23]. Inscribed [PHB22]. Inscriptions [AFV16, MQD^{+}23, PFF^{+}15, PHB22, Sap19, YLFM22]. Inspection [KKW22]. inspiration [Arn08b]. Instagram [CYH^{+}22]. Installation [IEL21, NVN^{+}22]. Installations [RDH15]. Instruments [RAD^{+}23]. Intangible [APHGG23, CKNK18, Che14, HKP^{+}22, KMMK22, LPD16, SRH^{+}20, SKD19, TWZF24, VMH^{+}18, MCG11]. Integrated [CMGFOAFH23, Hu24, REBK^{+}16]. Integrating [CDM^{+}24, DAD^{+}22, KCJP24, RBXM15]. Integration [FKD23, HMP17, KCRE23]. Intelligence [DDDK^{+}24, LMWP24]. Intelligent [Haj21, LhL19, LMWP24, LWZ^{+}24]. Intentions [PTP23]. Interacting [BG14, CPR^{+}24, Hac14, HLYH14, PA14, WKNC14]. Interaction [GGN22, JDDAB^{+}24, KPV^{+}20, NCM^{+}19, PMG^{+}20, SU22, RRC^{+}10]. Interactions [DSP24]. Interactive [CMGFOAFH23, CSSC^{+}24, CRR^{+}14, FCP24, GHC21, Hu24, MRB^{+}14, MA12, MHJ22, RRL^{+}14, TAR20, VRC24, ZPS^{+}16]. Interest [PSBM15, SM18]. Interest-Based [SM18]. Interested [JTL19]. Interface [GCM^{+}20, UA20, VMH^{+}18, MA12]. Interfaces [BG14, HGLS23]. Interior [AY23]. internal [LBM^{+}12]. Internet [FWZ23, LhL19]. Interoperability [FPS^{+}21, NMM^{+}17]. Interoperable [MGB18]. Interpretable [JDDAB^{+}24]. Interpretations [SBvE^{+}23]. Interval [SKK^{+}20]. Interventions [MCKC24]. interviews [DOHO08]. intonarumori [DS09]. intricate [LBM^{+}12]. Introducing [ESKI21]. Introduction [CL12, DDDK^{+}24, Fri11, Hac14, HMG22, LLL21, MB13, SIS^{+}23, WHMW23]. Invasive [CTRG19]. investigate [AL10]. Investigation [FHA^{+}21, SKK^{+}20]. Involving [BSO22]. IoT [BBD^{+}24]. IoT-based [BBD^{+}24]. IPA [OAB^{+}24]. Iran [BA24, PT22]. Ireland [DMK^{+}22]. Islamic [DT08, KB18]. ISO [MGB18]. IsoCam [MRB^{+}14]. Issue [BFJ^{+}23, DDDK^{+}24, DNR17a, DNR17b, ERG^{+}19, Hac14, HMG22, LLL21, WHMW23, Arn08b, CL12, Fri11, MB13]. Issues [PHB22, BB08]. IT-requirements [LRD08]. Italian [CGP15, RNP21]. Italy [MQD^{+}23]. ItAnt [MQD^{+}23]. items [ASC12]. Iterative [HHJ^{+}24].

January [ERG^{+}19]. Japanese [VKS^{+}24]. Jewels [MM21]. Jiangnan [FST24]. Jiaohe [GWZ^{+}24]. JOCCH [Arn08b, DNR17b]. John [PPBH21]. Joint [TWZF24]. journals [Mim12]. Journey [CRR^{+}14, PMHdF24, HLYH14]. Judgements [SGMT21]. Just [PT19, CS08].

Kasbah [ZDC21]. Kayseri [GBAA22]. kinect [HLYH14]. kinect-based [HLYH14]. Knowledge [APHGG23, AKK^{+}24, BGP22, BCH^{+}15, DMK^{+}22, DMS23, DTM17, FL20, HBC24, LC21, Nic21, PLC^{+}20, RFA19, SZGE23, DI08]. Knowledge-driven [APHGG23, AKK^{+}24]. Kong [CLP24]. Kratylos [FKS22]. Kunqu [Che14].

Lab [ABG^{+}17]. Laban [ASC^{+}15]. Labanotation [ESKI21]. Label [AS23]. Label-free [AS23]. Labelling [KIM24]. Laboratories [BBC21]. Land [KCS14, MRAB21]. Landmark [SWZG23]. Landmark-free [SWZG23]. Landsat [Pay23]. Language [MPFC16, PFZG23, WKNC14, HLZ12, OTBC11]. Lapse [SKK^{+}20]. Large [ASC20, HMP17, MRB^{+}14, MKM21, RAB^{+}16, ZSV^{+}16, CCD^{+}11, LLS^{+}12]. large-format [LLS^{+}12]. Large-Scale [HMP17, ZSV^{+}16]. Laser [MDCS19, RTA22, UY22, MGM13, SHFS12]. Last [ATP11]. Latin [BS12, CMST22]. Layered [JVAP^{+}21]. layers [STLL13]. Learn [CDV21]. Learned [CN23]. Learning [ASBD21, AVV21, BA24, BDK23, BSO22, BGP22, CLP24, CWR^{+}23, CLCL20, FYYG22, GAKL24, HHJ^{+}24, HZ24, IAY15, IA18, JML24, KIM24, KOIC20, MVCK^{+}23, MDCS19, MRFF24, MBEB23, MMGGA21, MPFC16, NPR^{+}20, OAAA21, PLSL22, PPF^{+}23, PP23, Pay23, RBXM15, TV23, TJTHW24, VKS^{+}24, WSHH24, YZH^{+}24, YLFM22, AL10, HLB14]. Learning-based [YLFM22]. Legacy [HMP17, SS13]. legends [KZK14]. Legitimacy [HKvF22]. Leisure [MNL^{+}21]. Leonardo [AFGG21]. Lessons [CN23]. Let [KPV^{+}20]. Letters [HLT23]. LetterSampo [HLT23]. LetterSampo-Historical [HLT23]. level [KRJ23]. Levels [PLSL22]. Libraries [RRDC23]. Library [CCAS23, PMGZ^{+}24, ZKB^{+}20, ASC12, BS12, Cak22, OZY^{+}21]. librettos [Che14]. Lifecycle [IEL21]. Light [GMLLHFTT21, PAZ^{+}23, RTA22, GSGC08]. Lighting [FHA^{+}21, Tai23, GMMC09]. Limit [AS23]. Line [HVS^{+}15, HGCS11, KLST13]. line-drawings [HGCS11]. Linear [NWLZ23, PKO23]. Lines [PYR15]. lingual [MANC24]. Linguistic [FKS22, CL12]. linguistics [CL12]. link [HSB^{+}09]. Linkage [JK23]. Linked [CDM^{+}24, DMP^{+}17, DWD^{+}23, DH21, GGN22, KCRE23]. Linking [BGB24, CDV21]. Listen [HHJ^{+}24]. Listening [EMK^{+}21, VC11]. Literature [GFZ24, PFZG23, Che14]. Little [Erk22]. live [BB08]. Living [DYW21, LTP21, OAAA21]. Local [PAZ^{+}23]. Localization [RFB^{+}19]. Location [JTL19, MR15, RBXM15]. Location-Based [JTL19, RBXM15]. Long [SKK^{+}20, OZY^{+}21]. Look [PS22]. lossless [vBTS^{+}22]. Lost [DMK^{+}22, PMHdF24, PVPMP24]. Louvre [CKA13]. Loves [AGGD^{+}23]. Low [CAMR22, ZSV^{+}16]. Low-Cost [ZSV^{+}16]. LRM [GFZ24]. Lublin [MM21].

Machine [ASBD21, BSO22, BGP22, DCC^{+}20, FYYG22, GAKL24, HLB14, LLB24, MDCS19, MMGGA21, Pay23, TJTHW24, VKS^{+}24, SZF^{+}10]. MaDiH [SFB^{+}22]. Main [RTA22]. Making [MKZ21, MCKC24, WRKK^{+}22, DKB11]. Manage [SBN22]. Management [BBC21, DTM17, Fra22, HSMAK^{+}22, SLRV23, MBD^{+}13]. Manora [RKWC24]. Manresa [PVPMP24]. Manuscript [KCRE23, YPGR17]. Manuscripts [CRR^{+}14, PYR15, WRKK^{+}22]. Map [VQP^{+}24, DRAT11]. Mappers [CFdS22]. Mapping [CCF20, Eic21, GA16, MPPB24, RCB^{+}18, DCC^{+}10]. Maps [CSSC^{+}24, ZPS^{+}16]. Marble [NVN^{+}22]. Maria [RDH15]. Marine [KTHP15]. Maritime [FKD23, HHJ^{+}24]. Marketplace [HDZM16]. Martial [APHGG23]. Màrtir [PVPMP24]. Mary [SOÇT23]. Masked [PFZG23]. Masonry [AAÇ22, Cak22, DS21, FBP20, PGD^{+}22]. Mass [CC23, CGG^{+}18]. Massive [MRB^{+}14]. Masterpiece [CDV21].

Matching [SP16, VVSB14]. Material [PAZ^{+}23, SZMB11]. Materials [CCAS23]. Matrix [NCP^{+}24]. Matteo [Nic19b]. Maya [COGP16, COGP18, HOGP17]. Meaning [CRR^{+}14, Sto22]. Measuring [HLZ12, WM24]. Medellin [HDZM16]. Media [ACH24, BSW^{+}22, Ben23, FYYG22, FWZ23, KOIC20, OMRP22, ATP11]. Mediated [Dry24, PT19]. Medicine [CCN^{+}24]. Medieval [BPLP23, CNS15, CCF20, CMST22, CRR^{+}14, Eic21, YPGR17, SHFS12, TF23]. Medium [JML24, CS08]. MedLatinEpi [CMST22]. MedLatinLit [CMST22]. Meets [AMP^{+}17, Kim22, PGMBR17]. Meiji [VKS^{+}24]. Meiji-era [VKS^{+}24]. Memoirs [RNP21]. Memorial [RJIH22]. Memoriam [Nic19b, SE17]. Merging [HWSK14]. Merida [GMLLHFTT21]. Mesh [GA16]. Metadata [BGP22, CGL^{+}22, GCM^{+}20, KCRE23, HS08, MA12]. Method [FMMM22, KB18, Kim22, NWLZ23, SKK^{+}20, SPZ^{+}22, TWZF24, WZL^{+}21, RPA^{+}10]. Methodical [MKM21]. Methodological [HDZM16]. Methodologies [KC20, ZA24]. Methodology [SBN22]. Methods [Ben23, KF23, NTM^{+}23, SBC^{+}22, SKD19, UY22, YK21]. Metric [GA16]. Metropolitan [VC21]. Micro [LBC^{+}15, LBM^{+}12]. micro-CT [LBM^{+}12]. Microscopy [CNS15]. Midrash [WRKK^{+}22]. Migrant [Vio23]. migration [BB08]. Milan [GBG^{+}22]. million [BS12]. Minecraft [EEG^{+}21]. Ming [DYW21, DCC^{+}20]. Mingei [PDK^{+}22]. Minimally [CTRG19]. mining [Mim12]. Minoan [PHB22]. Mixed [BPF^{+}18, HMS20, NVN^{+}22, OKY15, YK21]. Mixed-Reality [OKY15]. mixing [FKS22]. Mobile [ADN21, BZ24, EMK^{+}23, MBEB23, RKWC24, REBK^{+}16, RBXM15, SLRV23, UY22, XLL^{+}24]. Moche [Pay23]. modal [BDK23]. modality [MRAB21]. Model [BCF^{+}21, DRH19, FBB^{+}24, HBC24, KCRE23, LWZ^{+}24, MGB18, MF21, Nic19a, PKO23, RAB^{+}19, RTF22, SCCF22, BBD^{+}12, DKB11]. Modeling [AL10, FKD23, GBAA22, GSGC08, HAK21, MGBM14, WLZ19, CKA13, HGP12, HES09, RGRG09]. Modelling [Che15, GAKL24, GNL20, HSMAK^{+}22, KHO^{+}23, ZA24]. Models [Che15, GBG^{+}22, GA16, JVAP^{+}21, MRB^{+}14, PFZG23, RAB^{+}16, CS08, DCC^{+}10, KFH09]. Modern [CGG^{+}18, MM21, TF23]. Modes [PT19]. Moments [Haj21]. Monitoring [BED^{+}23, Ulv22]. Mont’e [BPV^{+}15, RAB^{+}16]. Monument [AFG15]. Monuments [AGGD^{+}23, MSWS20, MKM21, MPPB24, ZA24]. Moorish [NB17]. Moroccan [SB22]. Morocco [SBN22]. Morpho [MPFC16]. Morpho-Syntactic [MPFC16]. Morphological [DKC23]. Motif [TJTHW24]. Motion [ESKI21]. Motion-based [ESKI21]. Motivation [DH21]. Mountains [FST24]. Movement [ASC^{+}15]. Moving [ESKI21]. MR [ZA24]. Multi [BDK23, CCZ^{+}17, GBG^{+}22, JVAP^{+}21, JZF23, KY24, MRAB21, MCV23, Pay23, PAZ^{+}23, RAD^{+}23, DRAT11, RGRG09]. Multi-analytical [GBG^{+}22]. Multi-area [JZF23]. Multi-CNN-based [KY24]. Multi-Layered [JVAP^{+}21]. Multi-Light [PAZ^{+}23]. Multi-modal [BDK23]. Multi-modality [MRAB21]. multi-resolution [RGRG09]. Multi-specimen [MCV23]. Multi-Temporal [Pay23, RAD^{+}23]. multi-user [DRAT11]. Multi-View [CCZ^{+}17]. multilayer [HLZ12]. Multimedia [FCP24, PPGLPMR24, MA12]. Multimodal [PDK^{+}22, WZAM^{+}21]. Multiple [ORD^{+}21]. Multiple-source [ORD^{+}21]. Multiscale [CCD^{+}11]. Multispectral [GCL^{+}19]. Multiview [HWSK14]. Multivocal [SSLdB24].

Municipal [HSMAK^{+}22]. Munsell [MSTG18]. Muqarnas [AY23, DKY24, GBAA22, SB22]. Murals [WLZ19, YZH^{+}24, ZLS^{+}22]. Museography [KCS14]. Museological [OZY^{+}21]. MUSETECH [DRH19]. Museum [ADN21, AVK^{+}22, APH24, BBD^{+}24, BKPW24, BSW^{+}22, BGP22, CAMR22, CGL^{+}21, CYH^{+}22, DRH19, EMK^{+}23, GFE19, IEL21, JZF23, Kim22, KSSW24, KY24, LLA^{+}19, LTP21, LRS^{+}21, Mas16, MNL^{+}21, NPR^{+}20, OAAA21, PT22, Por21, PTP23, RCB^{+}18, RBXM15, TMB^{+}20, UA20, VC21, YL24, YK21, AL10, ALBA13, GBBS11, KSZ^{+}11, YH13, Dry24, HHJ^{+}24]. Museums [CN23, CFRCDSC21, DSP24, HMS20, JML24, MHJ22, MLL^{+}22, PA14, REBK^{+}16, RRDC23, PLC^{+}20]. MuseUs [CMN13]. Music [ABG^{+}17, EKS^{+}22, EMK^{+}21, GLF^{+}20, KDBMG16, WZAM^{+}21, WM24, SZF^{+}10, VC11]. Musical [RAD^{+}23, VMH^{+}18, BB08, DS09]. Musicology [ABG^{+}17]. Myaamia [OTBC11]. Mycenaean [PKO23].

Named [BVLW22]. Names [LLL^{+}24]. Narrations [MPRS17]. Narrative [GHC21, TAR20]. Narratives [ESD^{+}22, OZY^{+}21, PDK^{+}22, VRC24, Vio23]. National [RJIH22]. Native [DWD^{+}23]. Natural [BBD^{+}24, BED^{+}23, HVS^{+}15, RRDC23, VMH^{+}18, WKNC14, WHMW23, LRD08]. Navigation [BV17, CSSC^{+}24]. Navigation-based [BV17]. Near [vBTS^{+}22]. Near-lossless [vBTS^{+}22]. Need [BA24, EKS^{+}22]. Nets [KRJ23]. Network [AFV16, ADD^{+}24, CYH^{+}22, DCC^{+}20, SZGE22, ZLS^{+}22, AS23, DI08]. Networking [NVCD22]. Networks [DS21, Haj21, SZGE23, Vio23]. Neural [AFV16, AS23, DS21, DCC^{+}20, Haj21, RFW^{+}22, RKG^{+}24, SZGE22, SZGE23]. Neural-network [AS23]. Nicosia [PPBH21]. NN [DS21]. Non [CDM^{+}24]. Non-Open [CDM^{+}24]. Notation [TS14, NJDZ11]. Note [ERG^{+}19]. Novel [GBG^{+}22, HZ24, PP23]. Novelty [PFF^{+}15]. Nrityakosha [MCG11]. Nugaal [FBB^{+}24]. Numerical [SZGE23]. Numismatics [CSS^{+}17].

O [CFC23, PPGLPMR24]. O-City [PPGLPMR24]. Ob [WSA^{+}18]. Object [CFdS22, KIM24, KKW22, MRFF24, PSA^{+}17]. Object-Graph [CFdS22]. Objects [DSP24, LKK^{+}20, MRFF24, NCM^{+}19, SP16, ABV11, KLST13, SRWK13]. Observations [AASFB23]. OCR [SZGE22]. Octagon [TFK^{+}09]. Octree [DKC23]. Off [CWR^{+}23]. Off-Site [CWR^{+}23]. Old [CWR^{+}23, PYR15]. Omnidirectional [NKRS15]. On-Site [CWR^{+}23, DMPC14]. One [LLD15]. Online [APH24, BKPW24, CFRCDSC21, DMS23, GFE19, PDK^{+}22, PWM^{+}17, PT22, PLP^{+}17, SPC^{+}23, TV23, ZJCN24]. Ontological [BCH^{+}15, CKNK18, HAK21, IA14, GBBS11]. Ontology [AASFB23, ADD^{+}24, BFJ^{+}23, CFC23, FKD23, GCM^{+}20, MRAB21, MCKC24, SBvE^{+}23, TF23, Tro22]. Ontology-Based [MCKC24, Tro22]. Ontology-driven [GCM^{+}20]. Open [CDM^{+}24, CCAS23, DMP^{+}17, DWD^{+}23, MSWS20, SCDS15]. Opening [SH22]. opera [Che14]. Operations [DKC23]. Opportunities [WHMW23]. Optical [CLN18, CGG^{+}18, TG15]. Optimized [SZGE22]. Optimizing [GA16]. Oracle [YLFM22]. Oral [PS22, SH22]. Orange [PPGLPMR24]. Organizing [ASC20, dMKBD22]. Orientation [WMJ^{+}18]. Orientations [RRJBMM16]. Oriented [Nic19a, RAB^{+}19]. Orienting [RCB^{+}18]. originals [MFRG10]. Originated [GA16]. Other [UY22]. Ottoman [YAGU09]. Outcomes [PPF^{+}23].

Outdoor [TSB21]. Overlapping [SKK^{+}20]. Overview [CBJN22, BB08].

Packaging [SBVVSL^{+}15]. Page [YPGR17]. Page-Based [YPGR17]. Pagoda [LXW24]. Painting [EES^{+}20, LKK^{+}20, SF18]. Paintings [CNS15, CCZ^{+}17, DYW21, GSL15, Geo24, MVCK^{+}23, MF21, VKS^{+}24, MWX^{+}12, RPA^{+}10, SDF^{+}12]. Pairwise [VVSB14]. Palestinian [MRAB21]. Palimpsest [FBB^{+}24]. Palladian [BG14]. Palladio [BG14]. Pan [PMG^{+}20]. Pandemic [BKPW24, MLL^{+}22, PT22]. Panel [CNS15]. Papers [CD18]. Papyrus [ASBD21]. Paradigms [PMGZ^{+}24]. Parallel [HLZL18, SHFS12]. Parametric [Che15, Haj21, NB17]. PARCOURS [NMM^{+}17]. part [DNR17b]. partially [PVPMP24]. Participation [DH21, DBYA18]. Participatory [HHJ^{+}24]. Passages [HLZL18]. Passive [CPR^{+}24]. Past [Hac14, PT19, TJTHW24, HLYH14]. Pasteur [Arn08b]. Paths [CAMR22]. Pattern [AASFB23, SP16, WSHH24, RPA^{+}10]. Pattern-Based [AASFB23]. Patterns [Haj21, KB18, NDN^{+}21, PFF^{+}15, SBS24, DT08, SDF^{+}12]. pavilions [DG11]. People [SV24]. Perceived [JTL19]. Perception [Hu24, PPBH21, SGMT21, HLYH14]. Perceptions [DFSO21]. Perceptually [MVCK^{+}23]. Perceptually-grounded [MVCK^{+}23]. Pere [PVPMP24]. Performance [Cak22, FCP24, TS14]. Performers [KRJ23]. Performing [SIS^{+}23, BB08]. Period [LDS16, SZGE22]. Period-specific [SZGE22]. Periodic [KB18]. Periods [KT16]. Person [LLL^{+}24, BCF^{+}21]. Personality [PTP23]. personalized [SLD^{+}12]. Personas [ADN21]. Perspective [Dry24, IAY15, LLL^{+}24, LWZ^{+}24, RFA19, SBC^{+}22]. Perspectives [CBJN22, Mün19, SSLdB24]. Persuasive [EMK^{+}21]. Peru [Pay23].

Peruvian [SBS24]. pervasive [CMN13]. Petri [KRJ23]. Petrographic [PSBM15]. PEVIAR [MFRG10]. philology [CL12]. Philosophy [GFE19]. PHiM [HSMAK^{+}22]. Photo [DMP^{+}17]. Photogrammetric [SDH^{+}16, Ulv22]. Photogrammetry [CC23, DSH^{+}15, RTA22, PZF11]. Photographic [SPC^{+}23, TG15]. Photographs [PG15]. Photography [KF23, DCC^{+}10]. Photometric [DMPC14]. Physical [SPZ^{+}22]. Piano [TS14]. pictorial [HGCS11]. Pictures [ZJCN24]. piece [LBM^{+}12]. Pieces [GMLLHFTT21, LBC^{+}15]. Pipeline [PSA^{+}17]. Pitch [WM24]. Pitch-Class [WM24]. Place [ESD^{+}22, MGM13]. Place-based [ESD^{+}22]. Placement [KSSW24]. Places [BCF^{+}21]. Plan [KML22]. Plant [CDV21]. Plateau [HES09]. Platform [AKK^{+}24, BED^{+}23, PDK^{+}22, PPGLPMR24, SPZ^{+}22, VRC24, SBL^{+}13]. Play [DFSO21, LRS^{+}21, FMAG13]. Playback [CLN18]. Playful [ESKI21]. Playing [BSDG^{+}21, LRS^{+}21, RKWC24]. Poetry [HZ24, ZM20]. Point [HWSK14, LXW24, Mün19, Nic19a, DKC23, RAB^{+}19, UY22, ZDC21, MGM13]. Pointillist [Geo24]. Points [FUD^{+}19]. Polychromatic [MSWS20]. Polychromy [SCDS15]. Polyvocal [FCL^{+}24]. Pomological [GBG^{+}22]. Pooling [OMRP22]. Portal [BBF^{+}17, MANC24]. portalada [CCD^{+}11]. Portrait [LA24]. Portuguese [CRR^{+}14]. Pose [MVCK^{+}23, SV24]. Poses [SV24]. Post [GB22, Vio23]. Post-authentic [Vio23]. Post-Covid [GB22]. Postphenomenological [Dry24]. Potential [AASFB23, BCF^{+}21]. Potsherds [RRJBMM16]. Pottery [MGBM14, PSBM15, SBS24, WMJ^{+}18, MD12]. Power [HKvF22]. Practical [KF23]. Practice [BTIR^{+}23, GB22, TS14]. Practices [Por21].

Prama [RAB^{+}16]. Pre [LLL^{+}24, LWZ^{+}24]. Pre-Qin [LLL^{+}24]. Pre-trained [LWZ^{+}24]. Prediction [FHA^{+}21]. Predictive [FMMM22, GMMC09]. Preexisting [KHO^{+}23]. Preferences [ADN21, KCJP24]. Prehistoric [PTP23]. Premodern [KCRE23]. Presence [GLS^{+}24, MQB22, PT19]. Presentation [PDK^{+}22, CCD^{+}11, SRWK13]. Presentations [KSZ^{+}11]. Preservation [BJC^{+}22, NKRS15, SRH^{+}20, ATP11, BB08, Che14, LBM^{+}12, MSN^{+}13]. Preserve [AFGG21]. Preserving [FCP24, RKWC24, ZPS^{+}16, MCG11]. Preventive [RAD^{+}23]. Principles [GHC21]. Printable [ESW20]. Printing [EES^{+}20, LBC^{+}15, NTM^{+}23, LBM^{+}12]. Privilege [BSO22]. Probabilistic [CKNK18, WZL^{+}21]. Problem [EEG^{+}21, SIS^{+}23]. Problem-Solving [EEG^{+}21]. Procedural [KA20, CKA13]. Procedure [HSMAK^{+}22]. Process [AKK^{+}24, BSO22, DNS23, FF22, MQD^{+}23]. Processes [ZMD^{+}22]. Processing [FYYG22, LWZ^{+}24, ORD^{+}21, RFA19]. Production [HVS^{+}15, SBVVSL^{+}15]. Products [ACH24]. professional [Yan13]. Profile [DMS23, SWZG23]. profiles [AL10]. Programmes [HRS^{+}16]. Progressive [OAB^{+}24]. Project [LCD^{+}15, MQD^{+}23, NMM^{+}17, DMK^{+}22, DSH^{+}15, PA14]. Projectile [FUD^{+}19]. Projection [LLA^{+}19, MRB^{+}14]. Projection-based [LLA^{+}19]. projector [MSN^{+}13]. projector-based [MSN^{+}13]. Projects [HHJ^{+}24]. Promotes [NVN^{+}22]. Promoting [PPGLPMR24]. Promotion [CCD^{+}21]. Properly [HHJ^{+}24]. Properties [BCFS15]. Proposal [AKK^{+}24, PPF^{+}23]. Prospect [MNL^{+}21, SCCF22]. Protection [MBEB23]. Prototype [BZ24, Mas16, SU22]. Provenance [MQB22]. Providing [BSO22]. Province [DCC^{+}20]. Provision [HKvF22].

Proxy [LKK^{+}20]. Pseudo

[KIM24, PLSL22]. Pseudo-Ground [PLSL22]. Public [BJC^{+}22, FWZ23]. Publishing [DNS23, HLT23]. Punchworks [CNS15]. Puppetry [ZKB^{+}20]. Pure [KCS14]. Pursuit [Nic21]. Push [WWS^{+}22]. Push-broom [WWS^{+}22]. Puzzles [ESW20].

Qin [LLL^{+}24]. Qing [DYW21, FST24]. Qinghai [DCC^{+}20]. Quadrant [Arn08b]. Qualitative [HGP12, JTL19]. Quality [AY23, BGP22, FF22, HRS^{+}16, NTM^{+}23, PHB22, PTP23]. Quantitative [Kim22]. Quasi [KB18]. Quasi-Periodic [KB18]. Querying [dMKBD22]. Question [SZGE23]. Question-answering [SZGE23].

Rabat [SBN22]. Race [OMRP22]. Rammed [GWZ^{+}24]. Range [SDH^{+}16, GMMC09, MGM13]. range-scanned [MGM13]. Rapid [PLC^{+}20]. Rare [SBC^{+}22]. Ray [DMPC14, LBM^{+}12]. Ray-on [DMPC14]. RDF [CFC23]. Reactivating [FCP24]. Reactivation [HDZM16]. Reading [KZK14]. Real [KY24, LKK^{+}20, PP23, ZSV^{+}16, AVK^{+}22]. Real-Time [PP23, ZSV^{+}16, KY24]. Real-world [LKK^{+}20]. Realities [JML24]. Reality [BJC^{+}22, BPF^{+}18, CPR^{+}24, CLP24, CWR^{+}23, CCFC23, CCD^{+}21, Dim22, DMPC14, HMS20, Hu24, HGLS23, IEL21, IRHH^{+}22, KKK^{+}20, NCP^{+}24, NVN^{+}22, OZY^{+}21, OKY15, PPF^{+}23, PGMBR17, PTP23, RRL^{+}14, SRH^{+}20, SLRV23, SOÇT23, TMB^{+}20, TLL^{+}24, TSB21, VQP^{+}24, VRC24, XLL^{+}24, YK21]. Reasoning [BGP22]. reassemblies [RRC^{+}10]. Reassembly [NWLZ23, PSA^{+}17, WZL^{+}21]. Recognition [AFG15, AFV16, DDGP20, BVLW22, CGG^{+}18, KY24, LA24, MM23, MGBM14, MSTG18, PP23, SBS24, SGH^{+}21, TSB21, YLFM22]. Recognizing [COGP18]. Recommendation [MRAB21]. Recommendations [KCJP24]. recommender [SLD^{+}12]. Reconstruct [LBC^{+}15]. Reconstructed [Cak22, MWL^{+}16]. Reconstruction [CMGFOAFH23, CCN^{+}24, EEG^{+}21, FF23, GSB18, HAK21, KLST13, LXW24, MSWS20, PAZ^{+}23, Sap19, SB22, SCDS15, SF18, SOÇT23, ZSV^{+}16, vBTS^{+}22, DG11, DS09, MSN^{+}13, PVPMP24, USF08]. Reconstructions [PA14]. Record [VYBG22]. recorded [DOHO08]. RecorDIM [PT22]. Recordings [RSM23, TS14, WM24]. Records [BSO22, CLN18, CFC23, Fra22, HMP17]. recovery [HLYH14, SZF^{+}10]. Recreation [ZA24]. Recreational [BV17]. Red [FWZ23]. Rediscovering [KKRS22]. Reference [OMRP22]. Referenced [DPG22]. reflectance [PCC^{+}10]. reflection [KSZ^{+}11]. Reflexivity [GK19]. regarding [ACH24]. Region [CDV21, FBB^{+}24, SM18, FST24]. Regional [BSW^{+}22]. Registration [LXW24, PG15]. Regularization [ZPS^{+}16]. Reinterpreting [LTP21]. Related [RFA19, CCAS23]. relatedness [GBBS11]. Relation [TWZF24, SS13]. Relationship [FST24]. Relationships [VKS^{+}24]. Reliability [BEES15]. Relics [WZL^{+}21]. relief [KLST13]. Relightable [JVAP^{+}21, RKG^{+}24]. Religious [BZ24, RTA22, ZA24]. reloaded [HSB^{+}09]. Remains [CPR^{+}24]. remediating [ATP11]. Rendering [OKY15]. Repetitive [SBS24]. Rephotography [AVV21]. Replica [NTM^{+}23]. Replicas [DSP24]. report [LLS^{+}12]. Reports [DNS23]. Repositories [FF23]. Representation [GFZ24, GSL15, MM16, PDK^{+}22, PBCOMTRR19]. Representations [COGP16, MWL^{+}16, WM24]. Representing [JDDAB^{+}24]. Reproducing [EES^{+}20]. Republican [KZK14]. Requirements [DAD^{+}22, FTM19, LRD08]. Research [BBF^{+}17, FWZ23, FL20, GAKL24, KML22, KFH09, KC20, KDBMG16, LRD08, LLB24, LMWP24, MSR^{+}17, PLC^{+}20, RJIH22, SKD19, YK21, Arn08b, DOHO08]. Researchers [Mün19]. Residential [KML22]. Resolution [PHB22, RGRG09, WWS^{+}22]. Resolutions [KHO^{+}23]. Resources [ERG^{+}19, HMP17, MRAB21]. Response [AAÇ22, Erk22]. Responses [BKPW24]. Restoration [Haj21, MCKC24, NMM^{+}17, YZH^{+}24]. Retrieval [BDK23, BGP22, BVLW22, FCWY21, MANC24, PBCOMTRR19, RRJBMM16, HGCS11, YAGU09]. Reuse [RTF22]. Reusing [CCF20]. Revealing [RRL^{+}14]. Review [BSO22, DH21, GAKL24]. revitalization [OTBC11]. RGB [GSB18]. RGB-D [GSB18]. RiC [CFC23]. RiC-O [CFC23]. Rich [CGL^{+}22, MNL^{+}21]. Rich-Prospect [MNL^{+}21]. Richards [DNR17b]. Richness [CDV21]. RIS3D [DPG22]. Risk [BSDG^{+}21]. Robotic [BJC^{+}22]. Robust [BED^{+}23, PG15]. Rock [FBB^{+}24, TJTHW24, WHC^{+}19, ZPS^{+}16]. Rock-Art [ZPS^{+}16]. Rocks [BJC^{+}22]. Role [CGL^{+}21, RKWC24]. Role-Playing [RKWC24]. Roman [FHA^{+}21, GNL20, KZK14, PZF11, RTF22]. Rome [PFF^{+}15]. Room [OZY^{+}21]. Rosary [PVPMP24]. Rothéneuf [BJC^{+}22]. Routing [BV17]. ROV [DSH^{+}15]. RTR [AFGG21]. Ruin [MM16]. Ruins [GWZ^{+}24]. Rule [LW16]. Rule-Based [LW16]. Rules [PFZG23]. Runicus [SBC^{+}22]. Rural [BED^{+}23].

Sabratha [PZF11]. Safeguarding [CKNK18, LPD16]. Saint [CCD^{+}21]. Saliency [ES16]. Sample [WHC^{+}19].

Sampled [CS08]. San [PZF11]. Sant [PVPMP24]. Santo [DNR17b]. Saphir [CLN18]. Satellite [Pay23]. Satisfaction [KY24]. Say [LLD15]. Scalable [LMWZ17]. Scale [Che15, HMP17, ZSV^{+}16]. Scaling [MPPB24]. Scan [BPV^{+}15]. scanned [MGM13]. Scanner [RTA22, UY22]. Scanning [EES^{+}20, GMLLHFTT21, LCD^{+}15, MDCS19, MKM21, RTA22, SHFS12]. Scattering [GNL20, GSGC08]. Scenario [FMMM22, WZAM^{+}21]. Sceptic [JTL19]. Scholarly [Mün19]. School [CCN^{+}24, TAR20, HH13]. School-based [HH13]. Schubert [WZAM^{+}21]. Science [BBD^{+}24, BBC21, CCAS23, HMG22, RRDC23, TJTHW24, Arn08b]. Sciences [BEES15]. Scientific [HVS^{+}15]. Score [GLF^{+}20]. Scores [WM24]. SCRABS [AMP^{+}17]. Scratches [TG15]. Script [PKO23]. Scripts [SBC^{+}22]. Sculpted [BJC^{+}22]. Sculptural [GMLLHFTT21, MSWS20]. Sculpture [BJC^{+}22]. Sculptures [GSB18, RAB^{+}16]. Sea [DSH^{+}15]. SeaLiT [FKD23]. Search [MRAB21, UA20]. Second [KA20, LLD15]. Segmentation [AVV21, HOGP17, MYS22, PLSL22, PGD^{+}22, Sap19, ZPS^{+}16, vBTS^{+}22]. Seismic [AAÇ22]. Selected [CD18]. Seljukid [GBAA22]. Semantic [BGP22, DMP^{+}17, DNS23, FPS^{+}21, FCWY21, GCM^{+}20, MRAB21, NMM^{+}17, PGD^{+}22, Tro22, USF08, BPLP23, BFJ^{+}23, HLT23, RJIH22, YH13]. semantically [SBL^{+}13]. Semantics [CGL^{+}22, TWZF24, HSB^{+}09]. Semi [HSMAK^{+}22, KML22, MDCS19]. Semi-automated [MDCS19]. Semi-automatic [KML22]. Semi-automatized [HSMAK^{+}22]. Semiautomatic [Che15, LDS16, PG15]. Sense [AFGG21, EKS^{+}22, JDDAB^{+}24]. Sensing [MR15, MBEB23]. Sentence [MPFC16]. Sentiment [JTL19]. Sergius [Erk22]. series [FWZ23]. Serious [ACF^{+}18, BSDG^{+}21, EMK^{+}23, ZOZ21, ALBA13, BBD^{+}12, CMN13, HH13, MBD^{+}13, MB13]. sets [MGM13]. Settings [TSB21]. Setups [MRB^{+}14]. Seventies [FCP24]. Seville [HSMAK^{+}22]. shading [PCC^{+}10]. ShadowPlay2.5D [ZM20]. Shape [DDGP20, BPV^{+}15, COGP16, ES16, NB17, PAZ^{+}23]. Shared [WSA^{+}18]. sharing [YH13]. sherds [MD12]. Shortening [MPFC16]. Showcase [HSB^{+}09]. SIAT [YN20]. Sign [RFW^{+}22]. Signals [MRFF24]. Signs [COGP18, PHB22]. SikuGPT [LWZ^{+}24]. Similar [SGMT21]. Similarities [KSSW24]. Similarity [BCFS15, PAZ^{+}23, WLZ19, ASC12, GBBS11]. simulating [SDF^{+}12]. Simulation [FBP20, HLYH14, HES09]. Single [AVV21, FHA^{+}21, YPGR17]. Site [CMGFOAFH23, CWR^{+}23, DMPC14, HAK21, KKK^{+}20, MBEB23, Pay23, PWM^{+}17, SKK^{+}20, XLL^{+}24, ALBA13, HES09]. Sites [Dim22, DCC^{+}20, DSP24, GWZ^{+}24, KOIC20, MRFF24, PPBH21, RFB^{+}19, Tai23, Ulv22, ZSV^{+}16, LDD08]. Situated [NPR^{+}20]. Slavery [LTP21]. Small [APH24]. Smart [Dim22, DSP24, Mas16, AMP^{+}17]. smartpen [OTBC11]. Smartphone [MR15]. Smartphone-Based [MR15]. SMO [GWZ^{+}24]. Social [ACH24, BSDG^{+}21, CYH^{+}22, CCFC23, DBYA18, FYYG22, FMG^{+}21, KOIC20, LSA^{+}24, OMRP22, VC11]. Socio [PDK^{+}22]. Socio-historical [PDK^{+}22]. Sofrim [WRKK^{+}22]. Software [CFC23, KRF^{+}14]. Solution [AFGG21, APH24, MBEB23]. Solving [EEG^{+}21]. Somalia [FBB^{+}24]. Song [LW16]. Songs [RSM23]. Sophia [Erk22]. Sound [YL24, VC11]. Source [SCDS15, ORD^{+}21]. sources [SBL^{+}13]. Southern [APHGG23, RKWC24]. Space [AY23, EMK^{+}21, Fra20, IEL21, KSSW24, Por21, RCB^{+}18, LDD08]. Space-adaptive [KSSW24]. Spaces [DBYA18, KSSW24]. Spain [HSMAK^{+}22]. Spanish [BZ24]. Sparse [GSL15, PAZ^{+}23]. Sparsely [WLZ19]. Spatial [FST24, JK22, JK23, KMMK22, RRL^{+}14, Tai23, VRC24]. spatialized [SBL^{+}13]. Spatio [MGB18]. Spatio-Temporal [MGB18]. Spatiotemporal [CSSC^{+}24]. Special [BFJ^{+}23, DDDK^{+}24, DNR17a, DNR17b, ERG^{+}19, Hac14, HMG22, LLL21, WHMW23, CL12, Fri11, MB13]. specific [SZGE22]. Specification [MSTG18]. specimen [MCV23]. Specimens [MCV23]. Spectral [CCAS23, KF23]. Spherical [PZF11, RRJBMM16]. SS [Erk22]. St. [PPBH21]. Standard [MGB18]. State [DAD^{+}22, HKP^{+}22]. States [KHO^{+}23]. Stock [NVCD22]. Stone [MPPB24]. stopping [AS23]. Stories [NVCD22, TV23]. Story [CN23]. Story-Based [CN23]. Storytelling [BEES15, KPV^{+}20, KMMK22, LC21, TLL^{+}24]. Strategies [CN23, OMRP22, FPMT10]. streaming [SRWK13]. Street [CWR^{+}23]. Structural [Cak22, Nic19a, RAB^{+}19]. Structure [Cak22, RTF22, WLZ19, ZLS^{+}22, LBM^{+}12]. Structure-Guided [ZLS^{+}22]. Structured [GMLLHFTT21, RTA22]. Structures [KHO^{+}23, NDN^{+}21, NVCD22, WM24]. Structuring [BGB24]. Studies [Eic21, MR15]. Studio [KF23]. Study [AASFB23, BKPW24, BSDG^{+}21, EEG^{+}21, EMK^{+}21, FMG^{+}21, FCL^{+}24, FHA^{+}21, GWZ^{+}24, GLS^{+}24, HDZM16, IEL21, JML24, KDBMG16, LLB24, NCP^{+}24, OZY^{+}21, RSM23, SGH^{+}21, SCDS15, SKD19, TS14, UY22, AL10, ALBA13, Che14, CMN13, HLZ12, LBM^{+}12, LLS^{+}12, PPBH21, SZMB11, MMGGA21]. Style [ES16, GSL15, MVCK^{+}23, VKS^{+}24]. Substitutional [IEL21]. Subtle [YL24]. Such [Sto22]. Supervised [WSHH24]. Supervision [MRFF24]. Support [BGP22, CCD^{+}21, CCAS23, PDK^{+}22, RRDC23, SZF^{+}10, DCC^{+}20]. Supported [CWR^{+}23]. Supporting [MCKC24, NMM^{+}17, NPR^{+}20, ORD^{+}21, ZKB^{+}20]. Surface [GWZ^{+}24, MBEB23, NWLZ23, SWZG23]. Survey [BPF^{+}18, DSH^{+}15, DSP24, MC22]. Surveying [FTM19]. Surveys [SDH^{+}16]. survival [HS08]. sustainable [HSB^{+}09]. Suzhou [DYW21]. SV [PAZ^{+}23]. SV-BRDF [PAZ^{+}23]. SVM [DCC^{+}20]. Syllabification [AD21]. Symbiosis [AVK^{+}22]. Symbolic [BGP22, PP23]. Symmetrical [SWZG23]. symmetry [DT08]. Syntactic [MPFC16]. Syntax [TWZF24]. System [BG14, CMGFOAFH23, CSS^{+}17, DWD^{+}23, DPG22, FF22, KRJ23, KY24, LLA^{+}19, LSA^{+}24, PTP23, SLRV23, MBD^{+}13, MSN^{+}13, NJDZ11, SLD^{+}12, VC11, YAGU09, Yan13]. Systematic [DH21, SZMB11]. Systems [CGL^{+}21, PSA^{+}17, VC21, STLL13, YN20].

Tablet [GLF^{+}20]. Tablets [PKO23]. Tagging [VC21]. Tailored [SBVVSL^{+}15]. Taiwan [CWR^{+}23]. Taiwanese [HH13]. Taking [FWZ23]. Talk [KPV^{+}20]. Tang [YZH^{+}24]. Tangible [DSP24, HGLS23, KKW22, NPR^{+}20, NCM^{+}19, RRC^{+}10]. Tango [SKD19]. Tanhuma [WRKK^{+}22]. taxonomy [FPMT10]. Teaching [CLP24, LTP21]. Technical [MGBM14, SPC^{+}23, BB08]. Technique [TSB21, Ulv22]. Techniques [GMLLHFTT21, KOIC20, LBC^{+}15, LMWP24, PP23, RTA22, HGP12]. Technological [SCCF22, CS08, Yan13]. Technologically [Dry24]. Technologies [LC21, SIS^{+}23, AL10, HLYH14]. Technology [DRH19, HMS20, KKRS22, PMG^{+}20, RKWC24, SOÇT23, YZH^{+}24, LBM^{+}12, OTBC11, SS13]. Technology-Enhanced [PMG^{+}20]. Teenage [CN23]. Tell [COGP18].

Temporal [MGB18, MQB22, Pay23, RAD^{+}23]. Tensions [OMRP22]. Terms [DMS23]. Terrestrial [RTA22]. Text [Fra22, FTM19, PKO23, PYR15]. Textile [NDN^{+}21]. Texts [CGG^{+}18, CMST22, FST24, LWZ^{+}24, SZGE22, TWZF24, ZJCN24]. Textual [FF22, JK23]. Texture [WLZ19, SRWK13]. Textures [SBS24]. Thai [RKWC24]. Theater [PZF11]. Theatre [BTIR^{+}23, IRHH^{+}22]. Their [Geo24]. Them [KPV^{+}20]. Thematic [HZ24, KSSW24]. Theologian [PPBH21]. Theoretical [GAKL24, LLD15]. Thera [RPA^{+}10]. Theran [SDF^{+}12]. Thesaurus [LDS16, ZLSD13]. these [SGMT21]. ThIATRO [FMAG13]. Things [LhL19]. Thirty [LLD15]. Thirty-One [LLD15]. Thirty-Second [LLD15]. Thos. [NVCD22]. thousand [BS12]. Three [SGH^{+}21, VC21]. Three-Dimensional [SGH^{+}21]. Thresholds [TMB^{+}20]. Throwing [MGBM14]. Tikkoun [WRKK^{+}22]. Time [CDV21, FWZ23, KT16, PP23, RCB^{+}18, SKK^{+}20, ZSV^{+}16, HLYH14, KY24, LDD08]. Time-Lapse [SKK^{+}20]. Time-series [FWZ23]. Timurid [MKM21]. Tinian [NVN^{+}22]. Tito [LRS^{+}21]. TLS [LXW24, ZA24]. Tomb [YZH^{+}24]. Tomography [LBC^{+}15]. Tonal [WM24]. Tool [CCAS23, DKY24, FMG^{+}21, GA16, KML22, WSA^{+}18, ZM20, DRAT11]. Tools [BTIR^{+}23, GHC21, SDH^{+}16, SCDS15]. Tophane [AAÇ22]. Tophane-i [AAÇ22]. Topic [MANC24]. Topics [Mün19]. Topography [EES^{+}20, GWZ^{+}24]. Toponyms [BPLP23]. Touchscreens [Dry24]. Tour [KKK^{+}20]. Tourism [FCWY21, FWZ23, GLS^{+}24]. Town [Che15, LTP21]. Track [VQP^{+}24]. Tracking [BKPW24, CAMR22, MQB22, ZJCN24]. Traditional [AY23, AKK^{+}24, KKRS22, NDN^{+}21, RKWC24, SPZ^{+}22, WKNC14, ZMD^{+}22, ZKB^{+}20, MWX^{+}12]. Train [Erk22]. Train-induced [Erk22]. trained [LWZ^{+}24]. Transcription [SBC^{+}22]. transdisciplinarity [LRD08]. Transfer [BDK23, KIM24, LC21, MVCK^{+}23]. Transferring [SPZ^{+}22]. transformation [PCC^{+}10]. Transforming [GB22]. transition [SZMB11]. Translation [LLB24]. Translations [LLB24]. Transnational [SFB^{+}22]. TransR [HBC24]. Travel [PT19]. treatise [DG11]. treatment [HS08]. Trends [SGH^{+}21]. Trinity [OZY^{+}21]. True [GCL^{+}19]. True-Color [GCL^{+}19]. truth [PLSL22]. TTProfiler [DMS23]. Tunisian [YN20]. Turkey [Cak22]. Turkish [AY23]. Twitter [FMMM22]. Two [CMST22, GMLLHFTT21, BS12]. Types [DMS23]. typology [HLB14].

UAV [Ulv22, ZA24]. Ubiquitous [SBO^{+}17]. UEQ [OAB^{+}24]. UEQ-IPA [OAB^{+}24]. UK [PLC^{+}20]. Ultra [SKK^{+}20]. Ultra-Long [SKK^{+}20]. Uncertainty [MQB22]. Understanding [JTL19]. Underwater [DSH^{+}15]. UNESCO [KKRS22]. unitttitle [MANC24]. University [GBG^{+}22]. Unorganized [DKC23]. Unsettling [DFSO21]. Urban [HSMAK^{+}22]. Usability [LSA^{+}24, PMGZ^{+}24]. Use [BSO22, CFRCDSC21, DNS23, GHC21]. User [ACH24, EMK^{+}23, FTM19, GCM^{+}20, GGN22, HKvF22, HGLS23, IAY15, JML24, JTL19, KCJP24, NCP^{+}24, NVN^{+}22, OZY^{+}21, OAB^{+}24, PLP^{+}17, PTP23, RDH15, RKG^{+}24, SSLdB24, SBC^{+}22, VQP^{+}24, VMH^{+}18, YK21, DRAT11]. User-Centered [RDH15]. User-Centred [PLP^{+}17]. User-Curated [SSLdB24]. User-Friendly [RKG^{+}24]. Using [AVV21, AFV16, ASC^{+}15, BA24, BPLP23, CKA13, CAMR22, DSH^{+}15, DS21, EES^{+}20, FYYG22, FKS22, GBBS11, GWZ^{+}24, HS08, HLT23, KRJ23, KT16, KOIC20, LBC^{+}15, LC21, MR15, NCP^{+}24, OKY15, OAAA21, OAB^{+}24, DKC23, PT22, RRJBMM16, SRH^{+}20, SCDS15, SF18, Ulv22, YZH^{+}24, ZA24, DT08, HLYH14, MM23, MSTG18, Pay23, RGRG09]. Utilization [HGLS23]. UV [DLI^{+}17]. UV-Induced [DLI^{+}17]. UX [GHC21, KC20].

V&A [TV23]. Validation [AS23, BGB24]. Valley [Pay23]. Valorization [FMG^{+}21]. Value [HMP17, JTL19, KHO^{+}23]. Vase [MVCK^{+}23]. Vector [AFV16, DCC^{+}20, SZF^{+}10]. Vernacular [ZDC21]. Versa [VQP^{+}24]. Very [CGP15, CCD^{+}11]. Vested [HKvF22]. via [CYH^{+}22, CCFC23, DCC^{+}10, HOGP17, MVCK^{+}23, MPFC16]. viable [CS08]. Vibrations [Erk22]. Vice [VQP^{+}24]. Video [OZY^{+}21, PP23, WSA^{+}18, ZM20]. Videogames [BCF^{+}21]. View [CCZ^{+}17, Mün19]. viewing [DRAT11]. Violins [DLI^{+}17]. Virgin [SOÇT23]. Virtual [BJC^{+}22, BPF^{+}18, CPR^{+}24, CGL^{+}21, CGP15, CCN^{+}24, FCL^{+}24, FBB^{+}24, HAK21, Hu24, HGLS23, IAY15, IA18, IRHH^{+}22, JML24, KSSW24, LSA^{+}24, MHJ22, PPF^{+}23, PA14, PT22, PTP23, PT19, SRH^{+}20, SCDS15, SPZ^{+}22, SOÇT23, TMB^{+}20, TLL^{+}24, VQP^{+}24, WHC^{+}19, GSGC08, MWX^{+}12, RRC^{+}10, WD13]. Virtual/Augmented [Hu24]. virtualization [BB08]. Virtually [LTP21]. Virtue [Nic21]. VISE [UA20]. Visible [GCL^{+}19]. Visible-Infrared [GCL^{+}19]. Visigothic [GMLLHFTT21]. Vision [VC21]. Visit [RBXM15, SLRV23]. Visitor [ADN21, CAMR22, HMS20, JZF23, KY24, Mas16, MR15, RBXM15, SGMT21, AL10, KSZ^{+}11]. Visitors [BBD^{+}24, CN23, CYH^{+}22, Dry24, NPR^{+}20, RFB^{+}19, RCB^{+}18, ZJCN24]. Visual [AFV16, BDK23, BGB24, BM22, KCJP24, MM16, MRB^{+}14, MNL^{+}21, PPBH21, RFA19, TF23, UA20]. Visualisation [HMS20, Vio23]. Visualization [GBG^{+}22, MKM21, RKG^{+}24, Sap19, SPZ^{+}22, FPMT10]. Visualizations [MHJ22]. Visualize [AFGG21]. Visualizing [COGP18, KTHP15]. Volumetric [OZY^{+}21]. VR [CLCL20, GLS^{+}24, KTHP15, KKW22, LSA^{+}24, PT19]. VR-based [KKW22]. VR-Mediated [PT19]. Vrouw [RDH15]. Vulnerabilities [AAÇ22].

Wall [SF18, RPA^{+}10, SDF^{+}12, DCC^{+}20]. wall-paintings [RPA^{+}10]. Walls [PGD^{+}22]. War [RJIH22, RNP21]. Warcbase [LMWZ17]. WarVictimSampo [RJIH22]. Weak [MRFF24]. Wearable [UY22]. Weaving [BCH^{+}15]. Web [BPLP23, BFJ^{+}23, BGP22, CCF20, DNS23, DRAT11, FKW17, HLT23, JVAP^{+}21, LMWZ17, OAB^{+}24, RJIH22, SBL^{+}13, YH13]. Web-Based [FKW17, DRAT11, JVAP^{+}21]. webGIS [CCF20]. WebGL [SRWK13]. WebGL-based [SRWK13]. Wedge [CTRG19]. Western [CLP24, VKS^{+}24, VKS^{+}24]. Western-style [VKS^{+}24]. Wheel [MGBM14]. Wheel-Throwing [MGBM14]. WiFi [JZF23]. Wiki [AGGD^{+}23]. Wikibase [KCRE23]. Wikipedia [PWM^{+}17]. Wikis [Tro22]. Wild [GK19, LXW24]. Wind [HES09]. Window [GNL20]. Winterreise [WZAM^{+}21]. Within [AVK^{+}22, AASFB23]. WMLS [UY22]. Woodblocks [NTM^{+}23]. Workflow [EEG^{+}21, HRS^{+}16, MA12]. Workflows [MSWS20]. Works [BPLP23, BG14]. Workshop [CD18]. World [GB22, KKRS22, OKY15, PMHdF24, SKK^{+}20, SPZ^{+}22, LKK^{+}20]. Worldwide [AGGD^{+}23]. Written [TF23].

X [LBM^{+}12]. X-ray [LBM^{+}12]. XML [CFC23]. years [BS12]. Young [NPR^{+}20]. Youngsters [BSDG^{+}21].

Zeri [DMP^{+}17]. Zernike [Haj21]. zither [SZF^{+}10]. ZoAM [PMHdF24]. zone [SHFS12].

References

Akturk:2022:DRS

[AAÇ22] F. Aktürk, F. Aras, and S. Çeribasi. Dynamic response and seismic vulnerabilities of the historic tophane-i amire: Dynamic and seismic analysis of a historic masonry building. Journal on Computing and Cultural Heritage (JOCCH), 15(3):52:1– 52:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3495225.

Antonini:2023:EOO

[AASFB23] Alessio Antonini, Alessandro Adamou, Mari Carmen Suárez-Figueroa, and Francesca Benatti. Experiential observations: an ontology pattern-based study on capturing the potential content within evidences of experiences. Journal on Computing and Cultural Heritage (JOCCH), 16 (3):58:1–58:??, September 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3586078.

Abdallah:2017:DML

[ABG^{+}17] Samer Abdallah, Emmanouil Benetos, Nicolas Gold, Steven Hargreaves, Tillman Weyde, and Daniel Wolff. The digital music lab: a big data infrastructure for digital musicology. Journal on Computing and Cultural Heritage (JOCCH), 10(1):2:1– 2:??, April 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Aliaga:2011:DFD

[ABV11] Daniel G. Aliaga, Elisa Bertino, and Stefano Valtolina. DECHO — a framework for the digital exploration of cultural heritage objects. Journal on Computing and Cultural Heritage (JOCCH), 3(3):12:1– 12:??, March 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Andreoli:2018:FDD

[ACF^{+}18] Roberto Andreoli, Angela Corolla, Armando Faggiano, Delfina Malandrino, Donato Pirozzi, Mirta Ranaldi, Gianluca Santangelo, and Vittorio Scarano. A framework to design, develop, and evaluate immersive and collaborative serious games in cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 11(1):4:1– 4:??, January 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ai:2024:SMA

[ACH24] Ziyan Ai, Dickson K. W.

Chiu, and Kevin K. W. Ho. Social media analytics of user evaluation for innovative digital cultural and creative products: Experiences regarding Dunhuang cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 17(3):42:1– 42:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3653307.

Asperti:2021:SDC

[AD21] Andrea Asperti and Stefano Dal Bianco. Syllabification of the Divine Comedy. Journal on Computing and Cultural Heritage (JOCCH), 14(3):27:1– 27:26, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3459011.

Asprino:2024:ONC

[ADD^{+}24] Luigi Asprino, Rossana Damiano, Marilena Daquino, Stefano De Giorgis, Aldo Gangemi, Antonio Lieto, Bruno Sartini, and Manuel Striani. An ontology network for citizen curation. Journal on Computing and Cultural Heritage (JOCCH), 17(4):72:1– 72:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3704729.

Almeshari:2021:MMG

[ADN21] Moneerah Almeshari, John Dowell, and Julianne Nyhan. Museum mobile guide preferences of different visitor personas. Journal on Computing and Cultural Heritage (JOCCH), 14(1): 9:1–9:13, February 2021. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3423186.

Addison:2015:E

[ADP15] Alonzo C. Addison, Livio De Luca, and Sofia Pescarin. Editorial. Journal on Computing and Cultural Heritage (JOCCH), 8(1):1:1– 1:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Amato:2015:FIC

[AFG15] Giuseppe Amato, Fabrizio Falchi, and Claudio Gennaro. Fast image classification for monument recognition. Journal on Computing and Cultural Heritage (JOCCH), 8(4):18:1– 18:??, August 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Apollonio:2021:HAP

[AFGG21] Fabrizio Ivan Apollonio, Riccardo Foschi, Marco Gaiani, and Simone Garagnani. How to analyze, preserve, and communicate Leonardo’s drawing? A solution to visualize in RTR Fine Art Graphics established from “the best sense”. Journal on Computing and Cultural Heritage (JOCCH), 14(3):36:1– 36:30, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3433606.

Amato:2016:VRA

[AFV16] Giuseppe Amato, Fabrizio Falchi, and Lucia Vadicamo. Visual recognition of ancient inscriptions using convolutional neural network and Fisher vector. Journal on Computing and Cultural Heritage (JOCCH), 9(4):21:1– 21:??, December 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Azizifard:2023:WLM

[AGGD^{+}23] Narges Azizifard, Lodewijk Gelauff, Jean-Olivier Gransard-Desmond, Miriam Redi, and Rossano Schifanella. Wiki loves monuments: Crowdsourcing the collective image of the worldwide built heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 20:1–20:??, March 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3569092.

Amrani:2024:PKC

[AKK^{+}24] Racha Amrani, Sabrina Kacher, Selma Khouri, Houda Oufaida, Safia Ouahab, and Mouna Cherrad. Proposal of a knowledge capitalization process to construct eco-diars: a knowledge-driven platform applied to traditional Algerian domestic architecture. Journal on Computing and Cultural Heritage (JOCCH), 17(1):4:1– 4:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3627166.

Antoniou:2010:MVP

[AL10] Angeliki Antoniou and George Lepouras. Modeling visitors’ profiles: a study to investigate adaptation aspects for museum learning technologies. Journal on Computing and Cultural Heritage (JOCCH), 3(2):7:1–7:??, September 2010. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Antoniou:2013:ASG

[ALBA13] Angeliki Antoniou, George Lepouras, Stavroula Bampatzia, and Hera Almpanoudi. An approach for serious game development for cultural heritage: Case study for an archaeological site and museum. Journal on Computing and Cultural Heritage (JOCCH), 6(4):17:1–17:??, November 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Amato:2017:BDM

[AMP^{+}17] Flora Amato, Vincenzo Moscato, Antonio Picariello, Francesco Colace, Massimo De Santo, Fabio A. Schreiber, and Letizia Tanca. Big data meets digital cultural heritage: Design and implementation of SCRABS, a Smart Context-awaRe Browsing Assistant for Cultural EnvironmentS. Journal on Computing and Cultural Heritage (JOCCH), 10(1):6:1–6:??, April 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Avgousti:2024:EOA

[APH24] A. Avgousti, G. Papaioannou, and S. Hermon. Enhancing online accessibility of digitized artifacts from small museum collections in Cyprus: an empirical evaluation of the CyprusArk solution. Journal on Computing and Cultural Heritage (JOCCH), 17(3):34:1– 34:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3648229.

Adamou:2023:FIH

[APHGG23] Alessandro Adamou, Davide Picca, Yumeng Hou, and Paula Loreto Granados-García. The facets of intangible heritage in southern Chinese martial arts: Applying a knowledge-driven cultural contact detection approach. Journal on Computing and Cultural Heritage (JOCCH), 16(3):63:1– 63:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3606702.

Arnold:2008:E

[Arn08a]  David Arnold.  Edito-

rial. Journal on Computing and Cultural Heritage (JOCCH), 1(2):7:1– 7:??, October 2008. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Arnold:2008:EII

[Arn08b] David Arnold. Editorial for inaugural issue of JOCCH: Pasteur’s Quadrant: Cultural heritage as inspiration for basic research in computer science. Journal on Computing and Cultural Heritage (JOCCH), 1(1): 1:1–1:??, June 2008. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ameryan:2023:HLL

[AS23] Mahya Ameryan and Lambert Schomaker. How to limit label dissipation in neural-network validation: Exploring label-free earlystopping heuristics. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 2:1–2:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3587168.

Abitbol:2021:MLB

[ASBD21] Roy Abitbol, Ilan Shimshoni, and Jonathan Ben-Dov. Machine learning based assembly of fragments of ancient papyrus. Journal on Computing and Cultural Heritage (JOCCH), 14(3): 33:1–33:21, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3460961.

Aletras:2012:CSB

[ASC12] Nikolaos Aletras, Mark Stevenson, and Paul Clough. Computing similarity between items in a digital library of cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 5(4):16:1– 16:??, December 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Aristidou:2015:FDE

[ASC^{+}15] Andreas Aristidou, Efstathios Stavrakis, Panayiotis Charalambous, Yiorgos Chrysanthou, and Stephania Loizidou Himona. Folk dance evaluation using Laban Movement Analysis. Journal on Computing and Cultural Heritage (JOCCH), 8(4):20:1–20:??, August 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Aristidou:2020:DDE

[ASC20] Andreas Aristidou, Ariel Shamir, and Yiorgos Chrysanthou. Digital dance ethnography: Organizing large dance collections. Journal on Computing and Cultural Heritage (JOCCH), 12(4):29:1–29:??, January 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3344383.

Alliata:2024:BEE

[ASY^{+}24] Giacomo Alliata, Kate Saccone, Yuchen Yang, Loïc Serafin, Samy Mannane, and Sarah Kenderdine. The BiographScope: Exploring early cinema in an immersive environment. Journal on Computing and Cultural Heritage (JOCCH), 17(4):70:1– 70:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3672084.

Ashley:2011:LHH

[ATP11] Michael Ashley, Ruth Tringham, and Cinzia Perlingieri. Last House on the Hill: Digitally remediating data and media for preservation and access. Journal on Computing and Cultural Heritage (JOCCH), 4(4):13:1–13:??, December 2011. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Antoniou:2022:RCC

[AVK^{+}22] Angeliki Antoniou, Maria Vayanou, Akrivi Katifori, Angeliki Chrysanthi, Filippia Cheilitsi, and Yannis Ioannidis. ‘Real change comes from within!”: Towards a symbiosis of human and digital guides in the museum. Journal on Computing and Cultural Heritage (JOCCH), 15(1):15:1– 15:19, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465557.

Ali:2021:SIF

[AVV21] Dilawar Ali, Steven Verstockt, and Nico Van De Weghe. Single image façade segmentation and computational rephotography of house images using deep learning. Journal on Computing and Cultural Heritage (JOCCH), 14(4):44:1– 44:17, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3461014.

Agirbas:2023:EMA

[AY23] Asli Agirbas and Enes Yildiz. The effect of muqarnas on acoustic quality of traditional Turkish bath interior space. Journal on Computing and Cultural Heritage (JOCCH), 16 (3):46:1–46:??, September 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3589248.

Bahrami:2024:DLI

[BA24] Mahdi Bahrami and Amir Albadvi. Deep learning for identifying Iran’s cultural heritage buildings in need of conservation using image classification and Grad-CAM. Journal on Computing and Cultural Heritage (JOCCH), 17(1):16:1– 16:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631130.

Bonardi:2008:PEM

[BB08] Alain Bonardi and Jérome Barthélemy. The preservation, emulation, migration, and virtualization of live electronics for performing arts: an overview of musical and technical issues. Journal on Computing and Cultural Heritage (JOCCH), 1(1):6:1– 6:??, June 2008. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Boukhelifa:2018:CI

[BBB^{+}18] Nadia Boukhelifa, Mike Bryant, Natasa Bulatović, Ivan Cukić, Jean-Daniel Fekete, Milica Knezević, Jörg Lehmann, David Stuart, and Carsten Thiel. The CENDARI infrastructure. Journal on Computing and Cultural Heritage (JOCCH), 11(2):8:1– 8:??, June 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Bordalo:2021:FDI

[BBC21] Rui Bordalo, Carlo Bottaini, and António Candeias. A framework design for information management in heritage science laboratories. Journal on Computing and Cultural Heritage (JOCCH), 14(1): 7:1–7:14, February 2021. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3417304.

Bellotti:2012:SGM

[BBD^{+}12] Francesco Bellotti, Riccardo Berta, Alessandro De Gloria, Annamaria D’ursi, and Valentina Fiore. A serious game model for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 5(4):17:1– 17:??, December 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Balducci:2024:EII

[BBD^{+}24] Fabrizio Balducci, Paolo Buono, Giuseppe Desolda, Rosa Lanzilotti, Antonio Piccinno, and Maria Francesca Costabile. Exploring the impact of an IoT-based game on the experience of visitors at a natural science museum. Journal on Computing and Cultural Heritage (JOCCH), 17(4):63:1– 63:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3652163.

Blanke:2017:EHR

[BBF^{+}17] Tobias Blanke, Michael Bryant, Michal Frankl, Conny Kristel, Reto Speck, Veerle Vanden Daelen, and René Van Horik. The European Holocaust Research Infrastructure portal. Journal on Computing and Cultural Heritage (JOCCH), 10(1):1:1–1:??, April 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Blecic:2021:FPC

[BCF^{+}21] Ivan Blecić, Sara Cuccu, Filippo Andrea Fanni, Vittoria Frau, Riccardo Macis, Valeria Saiu, Martina Senis, Lucio Davide Spano, and Alessandro Tola. First-person cinematographic videogames: Game model, authoring environment, and potential for creating affection for places. Journal on Computing and Cultural Heritage (JOCCH), 14(2):18:1– 18:29, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3446977.

Biasotti:2015:ASB

[BCFS15] Silvia Biasotti, Andrea Cerri, Bianca Falcidieno, and Michela Spagnuolo. 3D artifacts similarity based on the concurrent evaluation of heterogeneous properties. Journal on Computing and Cultural Heritage (JOCCH), 8(4):19:1– 19:??, August 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Brownlow:2015:OAC

[BCH^{+}15] Richard Brownlow, Stefano Capuzzi, Sven Helmer, Luciana Martins, Immanuel Normann, and Alex Poulovassilis. An ontological approach to creating an Andean weaving knowledge base. Journal on Computing and Cultural Heritage (JOCCH), 8(2):11:1– 11:??, March 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Banar:2023:TLV

[BDK23] Nikolay Banar, Walter Daelemans, and Mike Kestemont. Transfer learning for the visual arts: The multi-modal retrieval of iconclass codes. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 32:1–32:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3575865.

Barrientos:2023:RMP

[BED^{+}23] Francisco Barrientos, Aitziber Egusquiza, Claudia De Luca, Simona Tondelli, Pedro Martín-Lerones, David Olmedo, John Martin, Irina Pavlova, Jaime Gómez-García-Bermejo, and Eduardo Zalama Casanova. A robust monitoring platform for rural cultural and natural heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 37:1–37:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3593430.

Bounhas:2015:IRE

[BEES15] Ibrahim Bounhas, Bilel Elayeb, Fabrice Evrard, and Yahya Slimani. Information reliability evaluation: From Arabic storytelling to computer sciences. Journal on Computing and Cultural Heritage (JOCCH), 8(3):14:1– 14:??, May 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Bendito:2023:ACM

[Ben23] Petronio Bendito. Algorithmic color methods of media arts. Journal on Computing and Cultural Heritage (JOCCH), 16(1):10:1– 10:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3580602.

Bikakis:2023:ESI

[BFJ^{+}23] Antonis Bikakis, Roberta Ferrario, Stéphane Jean, Béatrice Markhoff, Alessandro Mosca, and Marianna Nicolosi Asmundo. Editorial: Special issue on Semantic Web and ontology design for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(3):54:1– 54:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3626254.

Baldissini:2014:IAP

[BG14] S. Baldissini and M. Gaiani. Interacting with the Andrea Palladio Works: The history of Palladian information system interfaces. Journal on Computing and Cultural Heritage (JOCCH), 7(2):11:1– 11:??, June 2014. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Blettery:2024:HIC

[BGB24] Emile Blettery and ValéRie Gouet-Brunet. Heritage iconographic content structuring: from automatic linking to visual validation. Journal on Computing and Cultural Heritage (JOCCH), 17(3):47:1– 47:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3666007.

Bobasheva:2022:LRC

[BGP22] Anna Bobasheva, Fabien Gandon, and Frederic Precioso. Learning and reasoning for cultural metadata quality: Coupling symbolic AI and machine learning over a semantic Web knowledge graph to support museum curators in improving the quality of cultural metadata and information retrieval. Journal on Computing and Cultural Heritage (JOCCH), 15(3):40:1– 40:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3485844.

Barreau:2022:DVR

[BJC^{+}22] Jean-Baptiste Barreau, Joëlle Jouneau, Christophe Charlet, Ghyslain Ferré, and Jérémie Robert. Digitization, virtual reality and robotic sculpture for the preservation and enhancement of the public heritage of the sculpted rocks of Rothéneuf. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 70:1–70:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3522595.

Ballatore:2024:TMO

[BKPW24] Andrea Ballatore, Valeri Katerinchuk, Alexandra Poulovassilis, and Peter T. Wood. Tracking museums’ online responses to the COVID-19 pandemic: a study in museum analytics. Journal on Computing and Cultural Heritage (JOCCH), 17(1):2:1– 2:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3627165.

Bogacz:2022:DAA

[BM22] Bartosz Bogacz and Hubert Mara. Digital Assyriology — advances in visual cuneiform analysis. Journal on Computing and Cultural Heritage (JOCCH), 15(2): 38:1–38:22, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3491239.

Bekele:2018:SAV

[BPF^{+}18] Mafkereseb Kassahun Bekele, Roberto Pierdicca, Emanuele Frontoni, Eva Savina Malinverni, and James Gain. A survey of augmented, virtual, and mixed reality for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 11(2): 7:1–7:??, June 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Bartalesi:2023:USW

[BPLP23] Valentina Bartalesi, Nicolo’ Pratelli, Emanuele Lenzi, and Paolo Pontari. Using Semantic Web to create and explore an index of toponyms cited in medieval geographical works. Journal on Computing and Cultural Heritage (JOCCH), 16(2):26:1– 26:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3582263.

Bettio:2015:MSE

[BPV^{+}15] Fabio Bettio, Ruggero Pintus, Alberto Jaspe Villanueva, Emilio Merella, Fabio Marton, and Enrico Gobbetti. Mont’e Scan: Effective shape and color digitization of cluttered 3D artworks. Journal on Computing and Cultural Heritage (JOCCH), 8(1):4:1– 4:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Bamman:2012:ETT

[BS12] David Bamman and David Smith. Extracting two thousand years of Latin from a million book library. Journal on Computing and Cultural Heritage (JOCCH), 5(1):2:1– 2:??, April 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Benito-Santos:2021:PDC

[BSDG^{+}21] Alejandro Benito-Santos, Amelie Dorn, Antonio G. Losada Gómez, Thomas Palfinger, Roberto Therón Sánchez, and Eveline Wandl-Vogt. Playing design: a case study on applying gamification to construct a serious game with youngsters at social risk. Journal on Computing and Cultural Heritage (JOCCH), 14(2): 13:1–13:19, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3427380.

Baron:2022:PME

[BSO22] Jason R. Baron, Mahmoud F. Sayed, and Douglas W. Oard. Providing more efficient access to government records: a use case involving application of machine learning to improve FOIA review for the deliberative process privilege. Journal on Computing and Cultural Heritage (JOCCH), 15(1):5:1– 5:19, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3481045.

Beale:2022:DCR

[BSW^{+}22] Gareth Beale, Nicole Smith, Theodore Wilkins, Guy Schofield, Jonathan Hook, and Anthony Masinton. Digital creativity and the regional museum: Experimental collaboration at the convergence of immersive media and exhibition design. Journal on Computing and Cultural Heritage (JOCCH), 15(4):78:1– 78:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3527620.

Boiko:2023:DTC

[BTIR^{+}23] Tetiana Boiko, Maryna Tatarenko, Kateryna Iudova-Romanova, Yuliya Tsyvata, and Yaroslav Lanchak. Digital tools in contemporary theatre practice. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 27:1–27:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3582265.

Baker:2017:CHR

[BV17] Kevin Baker and Steven Verstockt. Cultural heritage routing: a recreational navigation-based approach in exploring cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 10(4):24:1–24:??, October 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Brandsen:2022:CBD

[BVLW22] Alex Brandsen, Suzan Verberne, Karsten Lambers, and Milco Wansleeben. Can BERT dig it? Named entity recognition for information retrieval in the archaeology domain. Journal on Computing and Cultural Heritage (JOCCH), 15(3):51:1– 51:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3497842.

Barbosa:2024:SCP

[BZ24] Jennifer Carolina Cruz Barbosa and Herbert González Zymla. “The Spanish Cathedral”, a prototype of mobile application for the access to the religious cultural heritage of Gothic Hispanic cathedrals. Journal on Computing and Cultural Heritage (JOCCH), 17(2):24:1– 24:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3640456.

Cakir:2022:SPE

[Cak22] Ferit Cakir. Structural performance evaluation of reconstructed masonry structure: a case of Ephesus Celsus Library in Turkey. Journal on Computing and Cultural Heritage (JOCCH), 15(4):68:1– 68:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3517339.

Casano:2022:IBT

[CAMR22] Jonathan D. L. Casano, Jenilyn L. Agapito, Abigail S. Moreno, and Ma. Mercedes T. Rodrigo. INFbased tracking and characterization of museum visitor paths and behaviors using Bluetooth low energy beacons. Journal on Computing and Cultural Heritage (JOCCH), 15(2):26:1– 26:22, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3474830.

Colavizza:2022:AAO

[CBJN22] Giovanni Colavizza, Tobias Blanke, Charles Jeurgens, and Julia Noordegraaf. Archives and AI: an overview of current debates and future perspectives. Journal on Computing and Cultural Heritage (JOCCH), 15(1):4:1– 4:15, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3479010.

Cheng:2023:HCD

[CC23] Danzhao Cheng and Eugene Ch’ng. Harnessing collective differences in crowdsourcing behaviour for mass photogrammetry of 3D cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(1):19:1– 19:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3569090.

Cortea:2023:IAO

[CCAS23] I. M. Cortea, A. Chirosca, L. M. Angheluta, and G. Seritan. INFRA-ART: an open access spectral library of art-related materials as a digital support tool for cultural heritage science. Journal on Computing and Cultural Heritage (JOCCH), 16(2):40:1– 40:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3593427.

Callieri:2011:MAP

[CCD^{+}11] Marco Callieri, Antoni Chica, Matteo Dellepiane, Isaac Besora, Massimiliano Corsini, Jordi Moyés, Guido Ranzuglia, Roberto Scopigno, and Pere Brunet. Multiscale acquisition and presentation of very large artifacts: The case of portalada. Journal on Computing and Cultural Heritage (JOCCH), 3(4):14:1– 14:??, April 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Cisternino:2021:ARA

[CCD^{+}21] Doriana Cisternino, Laura Corchia, Valerio De Luca, Carola Gatto, Silvia Liaci, Liliana Scrivano, Anna Trono, and Lucio Tommaso De Paolis. Augmented reality applications to support the promotion of cultural heritage: The case of the Basilica of Saint Catherine of Alexandria in Galatina. Journal on Computing and Cultural Heritage (JOCCH), 14(4):47:1– 47:30, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3460657.

Cofrancesco:2020:MMC

[CCF20] Pacifico Cofrancesco, Consuelo Capolupo, and Alessia Frisetti. Mapping medieval Christianity. CARE-GIS: Reusing Web digital data in a webGIS application. Journal on Computing and Cultural Heritage (JOCCH), 12(4):26:1– 26:??, January 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3336125.

Chng:2023:SAR

[CCFC23] Eugene Ch’ng, Shengdan Cai, Pinyuan Feng, and Danzhao Cheng. Social augmented reality: Communicating via cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(2):29:1– 29:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3582266.

Cuperschmid:2024:VRD

[CCN^{+}24] Ana Regina Mizrahy Cuperschmid, Aline Prado Costa, Igor Neves, Flávio Raffaelli Fonseca, and Ethel Mizrahy Cuperschmid. Virtual reconstruction of the deconstructed heritage: The former building of the Belo Horizonte School of Medicine. Journal on Computing and Cultural Heritage (JOCCH), 17(2):21:1– 21:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3639931.

Chen:2017:MVF

[CCZ^{+}17] Long Chen, Jianda Chen, Qin Zou, Kai Huang, and Qingquan Li. Multi-view feature combination for ancient paintings chronological classification. Journal on Computing and Cultural Heritage (JOCCH), 10(2): 7:1–7:??, April 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Catalano:2018:BPS

[CD18] Chiara Eva Catalano and Livio De Luca. Best papers selected at the 14th Eurographics Workshop on Graphics and Cultural Heritage editorial. Journal on Computing and Cultural Heritage (JOCCH), 11(1): 1:1–1:??, January 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Carvalho:2024:ICE

[CDM^{+}24] Jason Carvalho, Enrico Daga, Paul Mulholland, Luigi Asprino, Chukwudi Uwasomba, Marilena Daquino, Aldo Gangemi, Mark Maguire, and Adam Stoneman. Integrating citizen experiences in cultural heritage archives with a linked non-open data hub. Journal on Computing and Cultural Heritage (JOCCH), 17(4):71:1– 71:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3706063.

Chandrasekar:2021:CCL

[CDV21] Krishna Kumar Thirukokaranam Chandrasekar, Emile Deman, and Steven Verstockt. Cross-collection linking of botanical imagery in Ghent altarpiece to learn more about Van Eyck’s masterpiece and to explore a region’s plant richness and diversity over time. Journal on Computing and Cultural Heritage (JOCCH), 14(3): 40:1–40:14, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3457184.

Clavaud:2023:RCS

[CFC23] Florence Clavaud, Thomas Francart, and Pauline Charbonnier. RiC-O converter: a software to convert EAC-CPF and EAD 2002 XML files to RDF datasets conforming to records in contexts ontology. Journal on Computing and Cultural Heritage (JOCCH), 16(3):42:1– 42:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3583592.

Costa:2022:EGD

[CFdS22] Lázaro Costa, Nuno Freitas, and João Rocha da Silva. An evaluation of graph databases and object-graph mappers in CIDOC CRM-compliant digital archives. Journal on Computing and Cultural Heritage (JOCCH), 15 (3):44:1–44:??, September 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3485847.

Cristobal-Fransi:2021:MDA

[CFRCDSC21] Eduard Cristobal-Fransi, José Ramón-Cardona, Natalia Daries, and Antoni Serra-Cantallops. Museums in the Digital Age: an analysis of online communication and the use of e-commerce. Journal on Computing and Cultural Heritage (JOCCH), 14(4): 56:1–56:21, December 2021. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3464977.

Christy:2018:MDE

[CGG^{+}18] Matthew Christy, Anshul Gupta, Elizabeth Grumbach, Laura Mandell, Richard Furuta, and Ricardo Gutierrez-Osuna. Mass digitization of early modern texts with optical character recognition. Journal on Computing and Cultural Heritage (JOCCH), 11(1):6:1– 6:25, January 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ceccacci:2021:RHF

[CGL^{+}21] Silvia Ceccacci, Andrea Generosi, Alma Leopardi, Maura Mengoni, and Ferruccio Mandorli. The role of haptic feedback and gamification in virtual museum systems. Journal on Computing and Cultural Heritage (JOCCH), 14(3):38:1– 38:14, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3453074.

Colla:2022:BSH

[CGL^{+}22] Davide Colla, Annamaria Goy, Marco Leontino, Diego Magro, and Claudia Picardi. Bringing semantics into historical archives with computeraided rich metadata generation. Journal on Computing and Cultural Heritage (JOCCH), 15(3):39:1– 39:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3484398.

Cignoni:2015:VEV

[CGP15] Giovanni A. Cignoni, Fabio Gadducci, and Stefano Paci. A virtual experience on the very first Italian computer. Journal on Computing and Cultural Heritage (JOCCH), 7(4):21:1– 21:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Chen:2014:ICH

[Che14] Gen-Fang Chen. Intangible cultural heritage preservation: an exploratory study of digitization of the historical literature of Chinese Kunqu opera librettos. Journal on Computing and Cultural Heritage (JOCCH), 7(1):4:1– 4:??, April 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Chevrier:2015:SPM

[Che15] Christine Chevrier. Semiautomatic parametric modelling of the buildings on town scale models. Journal on Computing and Cultural Heritage (JOCCH), 7(4): 20:1–20:??, February 2015.

CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Calogero:2013:UPM

[CKA13] Erica Calogero, Jaime Kaminski, and David Arnold. Using procedural modeling to explore alternative designs for the Louvre. Journal on Computing and Cultural Heritage (JOCCH), 6(4):19:1–19:??, November 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Chantas:2018:POF

[CKNK18] Giannis Chantas, Sotiris Karavarsamis, Spiros Nikolopoulos, and Ioannis Kompatsiaris. A probabilistic, ontological framework for safeguarding the intangible cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 11(3):12:1–12:??, September 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Crane:2012:ISI

[CL12] Gregory Crane and Anke Lüdeling. Introduction to the special issue on corpus and computational linguistics, philology, and the linguistic heritage of humanity. Journal on Computing and Cultural Heritage (JOCCH), 5(1):1:1– 1:??, April 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Chng:2020:EVE

[CLCL20] Eugene Ch’ng, Yue Li, Shengdan Cai, and Fui-Theng Leow. The effects of VR environments on the acceptance, experience, and expectations of cultural heritage learning. Journal on Computing and Cultural Heritage (JOCCH), 13(1):7:1–7:21, February 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3352933.

Chenot:2018:SOP

[CLN18] Jean-Hugues Chenot, Louis Laborelli, and Jean-Étienne Noiré. Saphir: Optical playback of damaged and delaminated analogue audio disc records. Journal on Computing and Cultural Heritage (JOCCH), 11(3):14:1–14:??, September 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Cheng:2024:ARB

[CLP24] Lee Cheng, Leung Kwok Prudence Lau, and Wing Yan Jasman Pang. Augmented reality book design for teaching and learning architectural heritage: Educational heritage in Hong Kong Central and Western District. Journal on Computing and Cultural Heritage (JOCCH), 17(4):65:1– 65:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3655628.

Calzado-Martinez:2023:IIS

[CMGFOAFH23] Alberto Calzado-Martínez, Ángel-Luis García-Fernández, Lidia M. Ortega-Alvarado, and Francisco-Ramón Feito-Higueruela. Integrated information system for 3D interactive reconstruction of an archaeological site. Journal on Computing and Cultural Heritage (JOCCH), 16(3):44:1– 44:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3586077.

Coenen:2013:MCS

[CMN13] Tanguy Coenen, Lien Mostmans, and Kris Naessens. MuseUs: Case study of a pervasive cultural heritage serious game. Journal on Computing and Cultural Heritage (JOCCH), 6(2): 8:1–8:??, May 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Corbara:2022:MMT

[CMST22] Silvia Corbara, Alejandro Moreo, Fabrizio Sebastiani, and Mirko Tavoni. Med-LatinEpi and MedLatin-Lit: Two datasets for the computational authorship analysis of medieval Latin texts. Journal on Computing and Cultural Heritage (JOCCH), 15(3):57:1– 57:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3485822.

Cesario:2023:LLE

[CN23] Vanessa Cesário and Valentina Nisi. Lessons learned on engaging teenage visitors in museums with story-based and game-based strategies. Journal on Computing and Cultural Heritage (JOCCH), 16(2):28:1– 28:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3575867.

Cacciari:2015:DMC

[CNS15] I. Cacciari, P. Nieri, and S. Siano. 3D digital microscopy for characterizing punchworks on medieval panel paintings. Journal on Computing and Cultural Heritage (JOCCH), 7(4):19:1–19:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Can:2016:ESR

[COGP16] Gülcan Can, Jean-Marc Odobez, and Daniel Gatica-Perez. Evaluating shape representations for Maya glyph classification. Journal on Computing and Cultural Heritage (JOCCH), 9(3):14:1–14:??, November 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Can:2018:HTA

[COGP18] Gülcan Can, Jean-Marc Odobez, and Daniel Gatica-Perez. How to tell ancient signs apart? Recognizing and visualizing Maya glyphs with CNNs. Journal on Computing and Cultural Heritage (JOCCH), 11(4):20:1–20:??, December 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3230670.

Cannavo:2024:PHC

[CPR^{+}24] Alberto Cannavò, Simona Pacchiotti, Nicola Retta, Martina Terzoli, Roberta Spallone, and Fabrizio Lamberti. Passive haptics and conversational avatars for interacting with ancient Egypt remains in highfidelity virtual reality experiences. Journal on Computing and Cultural Heritage (JOCCH), 17(2):29:1– 29:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3648003.

Correia:2014:DIE

[CRR^{+}14] N. Correia, T. Romão, A. Ricardo, T. Mota, M. J. Melo, R. Castro, R. Carvalho, and A. Miranda. Design of an interactive experience with medieval illuminations: A journey into the beauty and meaning of medieval Portuguese manuscripts. Journal on Computing and Cultural Heritage (JOCCH), 7(2): 13:1–13:??, July 2014. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Cignoni:2008:SMC

[CS08] Paolo Cignoni and Roberto Scopigno. Sampled 3D models for CH applications: a viable and enabling new medium or just a technological exercise? Journal on Computing and Cultural Heritage (JOCCH), 1(1): 2:1–2:??, June 2008. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Celesti:2017:ICB

[CSS^{+}17] Antonio Celesti, Grazia Salamone, Anna Sapienza, Marianna Spinelli, Mariangela Puglisi, and Maria Caltabiano. An innovative cloud-based system for the diachronic analysis in numismatics. Journal on Computing and Cultural Heritage (JOCCH), 10(4):23:1–23:??, October 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Casanova-Salas:2024:EAN

[CSSC^{+}24] Pablo Casanova-Salas, Javier Sevilla, Inmaculada Coma, Pedro Morillo, Mar Gaitán, and Cristina Portalés. Enhancing accessibility and navigation of heritage collections through interactive spatiotemporal maps. Journal on Computing and Cultural Heritage (JOCCH), 17(3):39:1– 39:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3652521.

Crespel:2019:WCM

[CTRG19] Thomas Crespel, Adrian Travis, Patrick Reuter, and Xavier Granier. Wedge cameras for minimally invasive archaeology. Journal on Computing and Cultural Heritage (JOCCH), 12(2): 14:1–14:??, June 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3284425.

Chiang:2023:ARS

[CWR^{+}23] Kai-Chun Chiang, Cathy Weng, Abirami Rathinasabapathi, Hsuan Chen, and Jin-Hwei Su. Augmented reality supported learning for cultural heritage of Taiwan in on-site and off-site environments: The case of a daxi old street. Journal on Computing and Cultural Heritage (JOCCH), 16(3):50:1– 50:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3589231.

Chang:2022:IMV

[CYH^{+}22] Mi Chang, Taeha Yi, Sukjoo Hong, Po Yan Lai, Ji Young Jun, and Ji-Hyun Lee. Identifying museum visitors via social network analysis of Instagram. Journal on Computing and Cultural Heritage (JOCCH), 15(3):56:1– 56:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3505635.

Daga:2022:ICE

[DAD^{+}22] Enrico Daga, Luigi Asprino, Rossana Damiano, Marilena Daquino, Belen Diaz Agudo, Aldo Gangemi, Tsvi Kuflik, Antonio Lieto, Mark Maguire, Anna Maria Marras, Delfina Martinez Pandiani, Paul Mulholland, Silvio Peroni, Sofia Pescarin, and Alan Wecker. Integrating citizen experiences in cultural heritage archives: Requirements, state of the art, and challenges. Journal on Computing and Cultural Heritage (JOCCH), 15(1):11:1– 11:35, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3477599.

Diaz:2018:AEC

[DBYA18] Paloma Díaz, Andrea Bellucci, Chien-Wen Yuan, and Ignacio Aedo. Augmented experiences in cultural spaces through social participation. Journal on Computing and Cultural Heritage (JOCCH), 11(4):19:1–19:??, December 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3230675.

Dellepiane:2010:ICA

[DCC^{+}10] M. Dellepiane, M. Callieri, M. Corsini, P. Cignoni, and R. Scopigno. Improved color acquisition and mapping on 3D models via flash-based photography. Journal on Computing and Cultural Heritage (JOCCH), 2(4):9:1– 9:??, February 2010. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Du:2020:DAE

[DCC^{+}20] Yumin Du, Wenwu Chen, Kai Cui, Jingke Zhang, Zhuo Chen, and Qiyong Zhang. Damage assessment of earthen sites of the Ming Great Wall in Qinghai Province: a comparison between Support Vector Machine (SVM) and BP neural network. Journal on Computing and Cultural Heritage (JOCCH), 13(2): 14:1–14:18, June 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3376120.

Daga:2024:ISI

[DDDK^{+}24] Enrico Daga, Rossana Damiano, Lily Díaz-Kommonen, Tsvi Kuflik, and Manuel Striani. Introduction to the special issue on extended intelligence for cultural engagement. Journal on Computing and Cultural Heritage (JOCCH), 17(4):62:1– 62:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3706102.

Angelo:2020:ASF

[DDGP20] Luca Di Angelo, Paolo Di Stefano, Emanuele Guardiani, and Caterina Pane. Automatic shape feature recognition for ceramic finds. Journal on Computing and Cultural Heritage (JOCCH), 13(3):20:1– 20:21, October 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3386730.

DeAngeli:2021:UPP

[DFSO21] Daniela De Angeli, Daniel J. Finnegan, Lee Scott, and Eamonn O’Neill. Unsettling play: Perceptions of agonistic games. Journal on Computing and Cultural Heritage (JOCCH), 14(2): 15:1–15:25, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3431925.

Das:2011:DRP

[DG11] Vinay Mohan Das and Yogesh K. Garg. Digital reconstruction of pavilions described in an ancient Indian architectural treatise. Journal on Computing and Cultural Heritage (JOCCH), 4 (1):1:1–1:??, August 2011. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Davis:2021:LDC

[DH21] Edie Davis and Bahareh Heravi. Linked data and cultural heritage: a systematic review of participation, collaboration, and motivation. Journal on Computing and Cultural Heritage (JOCCH), 14(2):21:1– 21:18, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3429458.

Doerr:2008:DGK

[DI08] Martin Doerr and Dolores Iorizzo. The dream of a global knowledge network

— a new approach. Journal on Computing and Cultural Heritage (JOCCH), 1(1): 5:1–5:??, June 2008. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Dima:2022:DFS

[Dim22]  Mariza Dima.  A de-

sign framework for smart glass augmented reality experiences in heritage sites. Journal on Computing and Cultural Heritage (JOCCH), 15(4):66:1– 66:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3490393.

Doerr:2011:FAC

[DKB11] Martin Doerr, Athina Kritsotaki, and Katerina Boutsika. Factual argumentation — a core model for assertions making. Journal on Computing and Cultural Heritage (JOCCH), 3(3): 8:1–8:??, March 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Paiva:2023:MOU

[DKC23] Pedro Victor Vieira De Paiva and Eloisa Dezen-Kempter and Marco Carvalho. Morphological operations on unorganized point clouds using octree graphs. Journal on Computing and Cultural Heritage (JOCCH), 16(1):12:1– 12:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3534930.

Dincer:2024:CAD

[DKY24] Sevde Gülizar Dinçer, Mustafa Korumaz, and Tugrul Yazar. A computeraided design tool for muqarnas. Journal on Computing and Cultural Heritage (JOCCH), 17(2):31:1– 31:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3648000.

Dondi:2017:AAU

[DLI^{+}17] Piercarlo Dondi, Luca Lombardi, Claudia Invernizzi, Tommaso Rovetta, Marco Malagodi, and Maurizio Licchelli. Automatic analysis of UV-induced fluorescence imagery of historical violins. Journal on Computing and Cultural Heritage (JOCCH), 10(2):12:1– 12:??, April 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Debruyne:2022:CKG

[DMK^{+}22] Christophe Debruyne, Gary Munnelly, Lynn Kilgallon, Declan O’Sullivan, and Peter Crooks. Creating a knowledge graph for Ireland’s lost history: Knowledge engineering and curation in the Beyond 2022 Project. Journal on Computing and Cultural Heritage (JOCCH), 15(2):25:1– 25:25, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3474829.

deMooij:2022:CFO

[dMKBD22] Jan de Mooij, Can Kurtan, Jurian Baas, and Mehdi Dastani. A computational framework for organizing and querying cultural heritage archives. Journal on Computing and Cultural Heritage (JOCCH), 15 (3):45:1–45:??, September 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3485843.

Daquino:2017:ESE

[DMP^{+}17] Marilena Daquino, Francesca Mambelli, Silvio Peroni, Francesca Tomasi, and Fabio Vitali. Enhancing semantic expressivity in the cultural heritage domain: Exposing the Zeri Photo Archive as linked open data. Journal on Computing and Cultural Heritage (JOCCH), 10(4):21:1– 21:??, October 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Durand:2014:RSP

[DMPC14] Emmanuel Durand, Frederic Merienne, Christian Pere, and Patrick Callet. Ray-on, an on-site photometric augmented reality device. Journal on Computing and Cultural Heritage (JOCCH), 7(2):7:1– 7:??, July 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Diop:2023:TTT

[DMS23] Lamine Diop, Béatrice Markhoff, and Arnaud Soulet. TTProfiler: Types and terms profile building for online cultural heritage knowledge graphs. Journal on Computing and Cultural Heritage (JOCCH), 16 (3):61:1–61:??, September 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3588443.

DeSanto:2017:ESI

[DNR17a] Massimo De Santo, Franco Niccolucci, and Julian Richards. Editorial: Special issue on digital infrastructures for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 10(1):1:1– 1:??, April 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

DeSanto:2017:JSI

[DNR17b] Massimo De Santo, Franco Niccolucci, and Julian Richards. JOCCH special issue on digital infrastructures for cultural heritage (2nd part). Guest Editors: Massimo De Santo, Franco Niccolucci, and Julian Richards. Journal on Computing and Cultural Heritage (JOCCH), 10(3): 13:1–13:??, August 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Derudas:2023:NAA

[DNS23] Paola Derudas, Federico Nurra, and Andreas Svensson. New AIR for the archaeological process? The use of 3D Web semantic for publishing archaeological reports. Journal on Computing and Cultural Heritage (JOCCH), 16(3):57:1– 57:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3594722.

DeJong:2008:ARI

[DOHO08] Franciska De Jong, Douglas W. Oard, Willemijn Heeren, and Roeland Ordelman. Access to recorded interviews: a research agenda. Journal on Computing and Cultural Heritage (JOCCH), 1(1):3:1– 3:??, June 2008. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Dutailly:2022:RRI

[DPG22] Bruno Dutailly, Jean-Christophe Portais, and Xavier Granier. RIS3D: a referenced information system in 3D. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 73:1–73:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3517043.

Diaz:2011:IWB

[DRAT11] Lily Díaz, Markku Reunanen, Blanca Acuña, and Atte Timonen. ImaNote: a Web-based multi-user image map viewing and annotation tool. Journal on Computing and Cultural Heritage (JOCCH), 3(4): 13:1–13:??, April 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Damala:2019:MMC

[DRH19] Areti Damala, Ian Ruthven, and Eva Hornecker. The MUSETECH model: a comprehensive evaluation framework for museum technology. Journal on Computing and Cultural Heritage (JOCCH), 12(1): 7:1–7:??, February 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3297717.

Drygalska:2024:TME

[Dry24] Ewa Drygalska. Technologically mediated experiences of visitors to the Frederic Chopin Museum: Postphenomenological perspective on touchscreens affordances. Journal on Computing and Cultural Heritage (JOCCH), 17(1):14:1– 14:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3631127.

DeGotzen:2009:MHF

[DS09] Amalia De Götzen and Stefania Serafin. The musical heritage of futurism: a digital reconstruction of the intonarumori family. Journal on Computing and Cultural Heritage (JOCCH), 2(2): 5:1–5:??, November 2009. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Drosopoulos:2021:DDC

[DS21] Georgios A. Drosopoulos and Georgios E. Stavroulakis. Data-driven computational homogenization using neural networks: FE 2-NN application on damaged masonry. Journal on Computing and Cultural Heritage (JOCCH), 14(1):6:1– 6:19, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3423154.

Drap:2015:RPD

[DSH^{+}15] Pierre Drap, Julien Seinturier, Bilal Hijazi, Djamal Merad, Jean-Marc Boi, Bertrand Chemisky, Emmanuelle Seguin, and Luc Long. The ROV 3D Project: Deep-sea underwater survey using photogrammetry: Applications for underwater archaeology. Journal on Computing and Cultural Heritage (JOCCH), 8(4):21:1– 21:??, August 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Duranti:2024:SOR

[DSP24] Daniele Duranti, Davide Spallazzo, and Daniela Petrelli. Smart objects and replicas: a survey of tangible and embodied interactions in museums and cultural heritage sites. Journal on Computing and Cultural Heritage (JOCCH), 17(1): 7:1–7:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631132.

Djibril:2008:IGP

[DT08] Mohamed Ould Djibril and Rachid Oulad Haj Thami. Islamic geometrical patterns indexing and classification using discrete symmetry groups. Journal on Computing and Cultural Heritage (JOCCH), 1(2): 10:1–10:??, October 2008. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Dragoni:2017:KMA

[DTM17] Mauro Dragoni, Sara Tonelli, and Giovanni Moretti. A knowledge management architecture for digital cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 10(3): 15:1–15:??, August 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Daquino:2023:CLO

[DWD^{+}23] Marilena Daquino, Mari Wigham, Enrico Daga, Lucia Giagnolini, and Francesca Tomasi. CLEF. A linked open data native system for crowdsourcing. Journal on Computing and Cultural Heritage (JOCCH), 16(3):41:1– 41:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3594721.

Ding:2021:ELE

[DYW21] Ding Ding, Xianjun Yu, and Zhonglin Wang. The evolution of the living environment in Suzhou in the Ming and Qing dynasties based on historical paintings. Journal on Computing and Cultural Heritage (JOCCH), 14(2):20:1– 20:14, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3430700.

Edwards:2021:BCD

[EEG^{+}21] B. Edwards, B. B. Edwards, S. Griffiths, F. F. Reynolds, A. Stanford, and M. Woods. The Bryn Celli Ddu Minecraft experience: a workflow and problem-solving case study in the creation of an archaeological reconstruction in Minecraft for cultural heritage education. Journal on Computing and Cultural Heritage (JOCCH), 14(2): 23:1–23:16, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3427913.

Elkhuizen:2020:GCT

[EES^{+}20] Willemijn Elkhuizen, Tessa Essers, Yu Song, Jo Geraedts, Clemens Weijkamp, Joris Dik, and Sylvia Pont. Gloss, color, and topography scanning for reproducing a painting’s appearance using 3D printing. Journal on Computing and Cultural Heritage (JOCCH), 12(4): 27:1–27:??, January 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Eichert:2021:DMM

[Eic21]  Stefan Eichert.  Digi-

tal mapping of medieval cemeteries: Case studies from Austria and Czechia. Journal on Computing and Cultural Heritage (JOCCH), 14(1):3:1– 3:15, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3406535.

ElRaheb:2022:DAE

[EKS^{+}22] Katerina El Raheb, Lori Kougioumtzian, Marina Stergiou, Dimitra Petousi, Akrivi Katifori, Katerina Servi, Vera Kriezi, Valia Vraka, Stefania Merakos, Alexandros Charkiolakis, Foteini Venieri, Maria Boile, and Yannis Ioannidis. Designing an augmented experience for a music archive: What does the audience need beyond the sense of hearing? Journal on Computing and Cultural Heritage (JOCCH), 15(4):80:1– 80:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3528366.

Erdbrink:2021:LSE

[EMK^{+}21] A. Erdbrink, J. Michael, R. Kortmann, M. Hamel, K. Van Eijck, and A. Verbraeck. Listening space: an exploratory case study on a persuasive game designed to enrich the experience of classical music concerts. Journal on Computing and Cultural Heritage (JOCCH), 14(4):49:1– 49:20, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3458677.

Elesini:2023:MSG

[EMK^{+}23] Urska Stanković Elesini, Hlede Miha, David Kristan, Andreja Korosec, Eni Protić, Andrej Ucakar, Urska Vrabic Brodnjak, and Joze Rugelj. Mobile serious game for enhancing user experience in museum. Journal on Computing and Cultural Heritage (JOCCH), 16(1):15:1– 15:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3569088.

Economou:2019:ENS

[ERG^{+}19] Maria Economou, Ian Ruthven, Areti Galani, Milena Dobreva, and Marco de Niet. Editorial note for special issue on the Evaluation of Digital Cultural Resources — January 2019. Journal on Computing and Cultural Heritage (JOCCH), 12(1):1:1– 1:??, February 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3307313.

Erkal:2022:RLH

[Erk22] A. Erkal. Response of Little Hagia Sophia (Church of SS Sergius and Bacchus) to adjacent train-induced vibrations. Journal on Computing and Cultural Heritage (JOCCH), 15(3):54:1– 54:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3495224.

Echavarria:2016:ADS

[ES16] Karina Rodriguez Echavarria and Ran Song. Analyzing the decorative style of 3D heritage collections based on shape saliency. Journal on Computing and Cultural Heritage (JOCCH), 9(4):20:1– 20:??, December 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Echavarria:2022:CEE

[ESD^{+}22] Karina Rodriguez Echavarria, Myrsini Samaroudi, Laurie Dibble, Edward Silverton, and Sophie Dixon. Creative experiences for engaging communities with cultural heritage through place-based narratives. Journal on Computing and Cultural Heritage (JOCCH), 15(2):33:1–33:19, June 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3479007.

ElRaheb:2021:MCM

[ESKI21] Katerina El Raheb, Marina Stergiou, Akrivi Katifori, and Yannis Ioannidis. Moving in the cube: a motionbased playful experience for introducing Labanotation to beginners. Journal on Computing and Cultural Heritage (JOCCH), 14(2): 14:1–14:26, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3427379.

Echavarria:2020:FAP

[ESW20] Karina Rodriguez Echavarria, Myrsini Samaroudi, and Tim Weyrich. Fracturing artefacts into 3D printable puzzles to enhance audience engagement with heritage collections. Journal on Computing and Cultural Heritage (JOCCH), 13(1):6:1– 6:22, February 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3351343.

Frinchillucci:2024:EDV

[FBB^{+}24] Gianluca Frinchillucci, Eugenio Bortolini, Stefano Biagetti, Alberto Urcia, Alessia Brucato, Sonkor Geyre, Ali Abdullah Warsame, Federico Prizzi, Francesco Spigonardi, Thomas Conci, Luca Natali, and Marco Madella. Emergency documentation and virtual model of a rock art palimpsest from Garowe, Nugaal Region, Somalia. Journal on Computing and Cultural Heritage (JOCCH), 17(4):59:1– 59:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3679022.

Fita:2020:ESA

[FBP20] Josep Lluis Fita, Gonzalo Besuievsky, and Gustavo Patow. Earthquake simulation on ancient masonry buildings. Journal on Computing and Cultural Heritage (JOCCH), 13(2):11:1– 11:18, June 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3372421.

Franke:2024:PAV

[FCL^{+}24] Isabel F. Franke, Stefania D. Conte, Claudia A. Libbi, Victor De Boer, and Tilo Hartmann. A polyvocal approach to virtual heritage: an immersive case study. Journal on Computing and Cultural Heritage (JOCCH), 17(4):68:1– 68:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3678177.

Fiordelmondo:2024:RPI

[FCP24] Alessandro Fiordelmondo, Sergio Canazza, and Niccoló Pretto. Reactivating and preserving interactive multimedia artworks: an analog performance from the Seventies. Journal on Computing and Cultural Heritage (JOCCH), 17(2): 27:1–27:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3647995.

Fang:2021:FFE

[FCWY21] Hui Fang, Chongcheng Chen, Xiaozu Wu, and Xiaoyan Ye. FESRCT: a framework for the event semantic retrieval of cultural tourism. Journal on Computing and Cultural Heritage (JOCCH), 14(3):37:1– 37:23, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3451994.

Fugini:2022:DPQ

[FF22] Mariagrazia Fugini and Jacopo Finocchi. Data and process quality evaluation in a textual big data archiving system. Journal on Computing and Cultural Heritage (JOCCH), 15(1): 2:1–2:19, February 2022.

CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3461015.

Flagg:2023:RAD

[FF23] Cristopher Flagg and Ophir Frieder. Reconstruction of artifacts from digital image repositories. Journal on Computing and Cultural Heritage (JOCCH), 16(1):14:1– 14:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3552298.

Frisky:2021:ISI

[FHA^{+}21] Aufaclav Zatu Kusuma Frisky, Agus Harjoko, Lukman Awaludin, Sebastian Zambanini, and Robert Sablatnig. Investigation of single image depth prediction under different lighting conditions: a case study of ancient Roman coins. Journal on Computing and Cultural Heritage (JOCCH), 14(4):54:1– 54:17, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465742.

Fafalios:2023:SOE

[FKD23] Pavlos Fafalios, Athina Kritsotaki, and Martin Doerr. The SeaLiT ontology — an extension of CIDOC-CRM for the modeling and integration of maritime history information. Journal on Computing and Cultural Heritage (JOCCH), 16(3):60:1– 60:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3586080.

Finkel:2022:ACM

[FKS22] Raphael Finkel, Daniel Kaufman, and Ahmed Shamim. Analyzing codemixing in linguistic corpora using Kratylos. Journal on Computing and Cultural Heritage (JOCCH), 15(1): 3:1–3:15, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3480238.

Foley:2017:WBI

[FKW17] Jacob Foley, Paul Kwan, and Mitchell Welch. A Web-based infrastructure for the assisted annotation of heritage collections. Journal on Computing and Cultural Heritage (JOCCH), 10(3):14:1– 14:??, August 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ferreira-Lopes:2020:DDA

[FL20] Prof. Patricia Ferreira-Lopes. A data-driven approach for architectural history knowledge. Capturing buildings’ construction events for historical research collaboration. Journal on Computing and Cultural Heritage (JOCCH), 13(2):15:1– 15:22, June 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3376925.

Froschauer:2013:AHC

[FMAG13] Josef Froschauer, Dieter Merkl, Max Arends, and Doron Goldfarb. Art history concepts at play with ThIATRO. Journal on Computing and Cultural Heritage (JOCCH), 6(2): 7:1–7:??, May 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Fontanella:2021:HSG

[FMG^{+}21] F. Fontanella, M. Molinara, A. Gallozzi, M. Cigola, L. J. Senatore, R. Florio, P. Clini, and F. Celis D’amico. HeGO, a social game as a tool for cultural heritage valorization: The case study of the Atina Historical Center. Journal on Computing and Cultural Heritage (JOCCH), 14(2): 12:1–12:16, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3431926.

Furini:2022:PMI

[FMMM22] Marco Furini, Federica Mandreoli, Riccardo Martoglia, and Manuela Montangero. A predictive method to improve the effectiveness of Twitter communication in a cultural heritage scenario. Journal on Computing and Cultural Heritage (JOCCH), 15(2): 21:1–21:18, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3470786.

Foni:2010:TVS

[FPMT10] Alessandro E. Foni, George Papagiannakis, and Nadia Magnenat-Thalmann. A taxonomy of visualization strategies for cultural heritage applications. Journal on Computing and Cultural Heritage (JOCCH), 3(1): 1:1–1:??, June 2010. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Fafalios:2021:FCC

[FPS^{+}21] Pavlos Fafalios, Kostas Petrakis, Georgios Samaritakis, Korina Doerr, Athina Kritsotaki, Yannis Tzitzikas, and Martin Doerr. FAST CAT: Collaborative data entry and curation for semantic interoperability in digital humanities. Journal on Computing and Cultural Heritage (JOCCH), 14(4):45:1– 45:20, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3461460.

Franceschet:2020:AS

[Fra20] Massimo Franceschet. Art for space. Journal on Computing and Cultural Heritage (JOCCH), 13(3):24:1– 24:9, October 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3402443.

Franks:2022:TCR

[Fra22] Jason Franks. Text classification for records management. Journal on Computing and Cultural Heritage (JOCCH), 15(3):42:1– 42:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3485846.

Friedlander:2011:ISI

[Fri11] A. Friedlander. Introduction to the special issue on eHeritage. Journal on Computing and Cultural Heritage (JOCCH), 4(3):8:1– 8:??, December 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Fan:2024:SRB

[FST24] Zhixiang Fan, Tianrui Song, and Xiaolan Tang. The spatial relationship between cities and mountains in Jiangnan Region in the Qing Dynasty based on historical texts. Journal on Computing and Cultural Heritage (JOCCH), 17(3):37:1– 37:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3648002.

Franzini:2019:DET

[FTM19] Greta Franzini, Melissa Terras, and Simon Mahony. Digital editions of text: Surveying user requirements in the digital humanities. Journal on Computing and Cultural Heritage (JOCCH), 12(1): 1:1–1:??, February 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3230671.

Flores:2019:CAA

[FUD^{+}19] Fernando Castillo Flores, Francisco García Ugalde, José Luis Punzo Díaz, Jesús Zarco Navarro, Alfonso Gastelum-Strozzi, María Del Pilar Angeles, and Mariko Nakano Miyatake. Computer algorithm for archaeological projectile points automatic classification. Journal on Computing and Cultural Heritage (JOCCH), 12(3):19:1– 19:??, October 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3300972.

Feng:2023:REE

[FWZ23] Xin Feng, Xu Wang, and Yue Zhang. Research on the effect evaluation and the time-series evolution of public culture’s Internet communication under the background of new media: Taking the information dissemination of red tourism culture as an example. Journal on Computing and Cultural Heritage (JOCCH), 16(1):7:1– 7:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3530999.

Fan:2022:UML

[FYYG22] Lizhou Fan, Zhanyuan Yin, Huizi Yu, and Anne J. Gilliland. Using machine learning to enhance archival processing of social media archives. Journal on Computing and Cultural Heritage (JOCCH), 15(3):46:1– 46:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3547146.

Guidi:2016:DMM

[GA16] Gabriele Guidi and Davide Angheleddu. Displacement mapping as a metric tool for optimizing mesh models originated by 3D digitization. Journal on Computing and Cultural Heritage (JOCCH), 9(2):9:1– 9:??, May 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Gil:2024:MDL

[GAKL24] Aleksander Gil, Yusuf Arayici, Bimal Kumar, and Richard Laing. Machine and deep learning implementations for heritage building information modelling: a critical review of theoretical and applied research. Journal on Computing and Cultural Heritage (JOCCH), 17(3):36:1– 36:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3649442.

Giannini:2022:CCT

[GB22] Tula Giannini and Jonathan P.

Bowen. Computational culture: Transforming archives practice and education for a post-covid world. Journal on Computing and Cultural Heritage (JOCCH), 15(3):47:1– 47:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3493342.

Gokmen:2022:CMA

[GBAA22] Sabri Gokmen, Altan Basik, Yusuf Aykin, and Sema Alacam. Computational modeling and analysis of Seljukid Muqarnas in Kayseri. Journal on Computing and Cultural Heritage (JOCCH), 15(2):27:1– 27:19, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3477399.

Grieser:2011:UOD

[GBBS11] Karl Grieser, Timothy Baldwin, Fabian Bohnert, and Liz Sonenberg. Using ontological and document similarity to estimate museum exhibit relatedness. Journal on Computing and Cultural Heritage (JOCCH), 3(3):10:1– 10:??, March 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Grifoni:2022:HDV

[GBG^{+}22] Emanuela Grifoni, Letizia Bonizzoni, Marco Gargano, Jacopo Melada, Nicola Ludwig, Silvia Bruni, and Ilaria Mignani. Hyperdimensional visualization of cultural heritage: a novel multi-analytical approach on 3D pomological models in the collection of the University of Milan. Journal on Computing and Cultural Heritage (JOCCH), 15(2): 34:1–34:15, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3477398.

Grifoni:2019:NIT

[GCL^{+}19] Emanuela Grifoni, Beatrice Campanella, Stefano Legnaioli, Giulia Lorenzetti, Luciano Marras, Stefano Pagnotta, Vincenzo Palleschi, Francesco Poggialini, Emanuele Salerno, and Anna Tonazzini. A new infrared true-color approach for visible-infrared multispectral image analysis. Journal on Computing and Cultural Heritage (JOCCH), 12(2):8:1– 8:??, June 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3241065.

Goy:2020:BSM

[GCM^{+}20] Annamaria Goy, Davide Colla, Diego Magro, Cristina Accornero, Fabrizio Loreto, and Daniele Paolo Radicioni. Building semantic metadata for historical archives through an ontology-driven user interface. Journal on Computing and Cultural Heritage (JOCCH), 13(3):25:1– 25:36, October 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3402440.

Georgoulaki:2024:CIP

[Geo24] Kristina Georgoulaki. Classification of impressionist and pointillist paintings based on their brushstrokes characteristics. Journal on Computing and Cultural Heritage (JOCCH), 17 (3):46:1–46:??, September 2024. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3665501.

Gil-Fuentetaja:2019:CMC

[GFE19] Ion Gil-Fuentetaja and Maria Economou. Communicating museum collections information online: Analysis of the philosophy of communication extending the constructivist approach. Journal on Computing and Cultural Heritage (JOCCH), 12(1):3:1– 3:??, February 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3283253.

Gagnon:2024:EIL

[GFZ24] Michel Gagnon, Ludovic Font, and Amal Zouaq. An exploration of IFLA LRM for literature data representation. Journal on Computing and Cultural Heritage (JOCCH), 17(3):50:1– 50:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3687486.

Guedes:2022:DUI

[GGN22] Cláudia Guedes, Bruno Giesteira, and Sérgio Nunes. Designing user interaction with linked data in historical archives. Journal on Computing and Cultural Heritage (JOCCH), 15 (3):43:1–43:??, September 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3485731.

Green:2021:UTU

[GHC21] Daniel Green, Charlie Hargood, and Fred Charles. Use of tools: UX principles for interactive narrative authoring tools. Journal on Computing and Cultural Heritage (JOCCH), 14(3): 41:1–41:25, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3458769.

Galani:2019:EDC

[GK19] Areti Galani and Jenny Kidd. Evaluating digital cultural heritage ‘in the wild’: The case for reflexivity. Journal on Computing and Cultural Heritage (JOCCH), 12(1):5:1– 5:??, February 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3287272.

Gaugne:2020:ETD

[GLF^{+}20] Ronan Gaugne, Françoise Labaune, Dominique Fontaine, Gaétan Le Cloirec, and Valérie Gouranton. From the engraved tablet to the digital tablet, history of a 15th-century music score. Journal on Computing and Cultural Heritage (JOCCH), 13(3):17:1– 17:18, October 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3383782.

Guo:2024:DHT

[GLS^{+}24] Yuqing Guo, Shizhu Lu, Min Shen, Wei Huang, Xiaolie Yi, and Jifa Zhang. Differences in heritage tourism experience between VR and AR: a comparative experimental study based on presence and authenticity. Journal on Computing and Cultural Heritage (JOCCH), 17(2):30:1– 30:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3648001.

Garcia-Molina:2021:DDT

[GMLLHFTT21] Diego F. García-Molina, Samuel López-Lago, Rafael E. Hidalgo-Fernandez, and Paula Triviño-Tarradas. Digitalization and 3D documentation techniques applied to two pieces of visigothic sculptural heritage in merida through structured light scanning. Journal on Computing and Cultural Heritage (JOCCH), 14(4): 43:1–43:19, December 2021. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3427381.

Goncalves:2009:HDR

[GMMC09] Alexandrino Gonçalves, Luís Magalhães, João Moura, and Alan Chalmers. High dynamic range — a gateway for predictive ancient lighting. Journal on Computing and Cultural Heritage (JOCCH), 2(1): 3:1–3:??, July 2009. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Grobe:2020:DDM

[GNL20] Lars Oliver Grobe, Andreas Noback, and Franziska Lang. Data-driven modelling of daylight scattering by Roman window glass. Journal on Computing and Cultural Heritage (JOCCH), 13(1):2:1– 2:20, February 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3350428.

Gomes:2018:ERD

[GSB18] Leonardo Gomes, Luciano Silva, and Olga Regina Pereira Bellon. Exploring RGB-D cameras for 3D reconstruction of cultural heritage: a new approach applied to Brazilian Baroque sculptures. Journal on Computing and Cultural Heritage (JOCCH), 11(4):21:1–21:??, December 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3230674.

Gutierrez:2008:MLS

[GSGC08] Diego Gutierrez, Veronica Sundstedt, Fermin Gomez, and Alan Chalmers. Modeling light scattering for virtual heritage. Journal on Computing and Cultural Heritage (JOCCH), 1 (2):8:1–8:??, October 2008. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Gao:2015:ASR

[GSL15] Zhi Gao, Mo Shan, and Qingquan Li. Adaptive sparse representation for analyzing artistic style of paintings. Journal on Computing and Cultural Heritage (JOCCH), 8(4):22:1– 22:??, August 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Guo:2024:CFR

[GWZ^{+}24] Zhiqian Guo, Chong Wu, Shuai Zhang, Wenwu Chen, Qiang Qi, and Haiyu Wu. Characterization on flaking of rammed earthen sites using SMO algorithm and surface topography analysis: a case study of Jiaohe Ruins. Journal on Computing and Cultural Heritage (JOCCH), 17(3):49:1– 49:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3673763.

Hachet:2014:ISI

[Hac14]  Martin Hachet.  Intro-

duction to special issue on interacting with the past. Journal on Computing and Cultural Heritage (JOCCH), 7(2):6e:1– 6e:??, July 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hajebi:2021:IRH

[Haj21]  Bita Hajebi.  Intelli-

gent restoration of historical parametric geometric patterns by Zernike moments and neural networks. Journal on Computing and Cultural Heritage (JOCCH), 14(4):42:1– 42:27, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3459607.

Hiebel:2021:OME

[HAK21] Gerald Hiebel, Edeltraud Aspöck, and Karin Kopetzky. Ontological modeling for excavation documentation and virtual reconstruction of an Ancient Egyptian site. Journal on Computing and Cultural Heritage (JOCCH), 14(3):32:1– 32:14, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3439735.

Hou:2024:CTK

[HBC24] Wenjun Hou, Bing Bai, and Chenyang Cai. CR-TransR: a knowledge graph embedding model for cultural domain. Journal on Computing and Cultural Heritage (JOCCH), 17(1):12:1– 12:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3625299.

Hincapie:2016:MFD

[HDZM16] Mauricio Hincapie, Christian Diaz, Maria Zapata, and Camilo Mesias. Methodological framework for the design and development of applications for reactivation of cultural heritage: Case study Cisneros Marketplace at Medellin, Colombia. Journal on Computing and Cultural Heritage (JOCCH), 9(2):8:1– 8:??, May 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hussein:2009:WFM

[HES09] Ashraf S. Hussein and Hisham El-Shishiny. Wind flow modeling and simulation over the Giza Plateau cultural heritage site in Egypt. Journal on Computing and Cultural Heritage (JOCCH), 2(2):6:1– 6:??, November 2009. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hurtut:2011:ALD

[HGCS11] Thomas Hurtut, Yann Gousseau, Farida Cheriet, and Francis Schmitt. Artistic line-drawings retrieval based on the pictorial content. Journal on Computing and Cultural Heritage (JOCCH), 4(1):3:1– 3:??, August 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hulusic:2023:TUI

[HGLS23] Vedad Hulusic, Linda Gusia, Nita Luci, and Michael Smith. Tangible user interfaces for enhancing user experience of virtual reality cultural heritage applications for utilization in educational environment. Journal on Computing and Cultural Heritage (JOCCH), 16(2):38:1– 38:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3593429.

Hug:2012:QEC

[HGP12] Charlotte Hug and Cesar Gonzalez-Perez. Qualitative evaluation of cultural heritage information modeling techniques. Journal on Computing and Cultural Heritage (JOCCH), 5(2): 8:1–8:??, July 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Huang:2013:ASB

[HH13] Chih-Hong Huang and Yi-Ting Huang. An Annales School-based serious game creation framework for Taiwanese indigenous cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 6(2):9:1– 9:??, May 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hodgson:2024:LLP

[HHJ^{+}24] Isabella Hodgson, Dennis Hoffmann, Luca Junge, Frederic Theis, and Nicolas Dittgen. Learning to listen properly: Participatory digital projects at the German Maritime Museum based on iterative cocreation. Journal on Computing and Cultural Heritage (JOCCH), 17(3):51:1– 51:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3656585.

Hou:2022:DIC

[HKP^{+}22] Yumeng Hou, Sarah Kenderdine, Davide Picca, Mattia Egloff, and Alessandro Adamou. Digitizing intangible cultural heritage embodied: State of the art. Journal on Computing and Cultural Heritage (JOCCH), 15(3):55:1– 55:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3494837.

Hoekstra:2022:VAE

[HKvF22] Rik Hoekstra, Marijn Koolen, and Marijke van Faassen. Vested authorities, emergent brokers and user archivists: Power and legitimacy in information provision. Journal on Computing and Cultural Heritage (JOCCH), 15(3):48:1– 48:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3484481.

Horr:2014:MLB

[HLB14] Christian Hörr, Elisabeth Lindinger, and Guido Brunnett. Machine learning based typology development in archaeology. Journal on Computing and Cultural Heritage (JOCCH), 7(1):2:1–2:??, April 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hyvonen:2023:LHL

[HLT23] Eero Hyvönen, Petri Leskinen, and Jouni Tuominen. LetterSampo-Historical letters on the Semantic Web: a framework and its application to publishing and using epistolary data. Journal on Computing and Cultural Heritage (JOCCH), 16(1):21:1– 21:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3569372.

Hsieh:2014:IPC

[HLYH14] Chun-Ko Hsieh, Wen-Ching Liao, Meng-Chieh Yu, and Yi-Ping Hung. Interacting with the past: Creating a time perception journey experience using kinect-based breath detection and deterioration and recovery simulation technologies. Journal on Computing and Cultural Heritage (JOCCH), 7(1):1:1– 1:??, February 2014. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hirschmann:2012:MCL

[HLZ12] Hagen Hirschmann, Anke Lüdeling, and Amir Zeldes. Measuring and coding language change: an evolving study in a multilayer corpus architecture. Journal on Computing and Cultural Heritage (JOCCH), 5(1): 4:1–4:??, April 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Harris:2018:FPP

[HLZL18] Martyn Harris, Mark Levene, Dell Zhang, and Dan Levene. Finding parallel passages in cultural heritage archives. Journal on Computing and Cultural Heritage (JOCCH), 11(3):15:1–15:??, September 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hedges:2022:ISI

[HMG22] Mark Hedges, Richard Marciano, and Eirini Goudarouli. Introduction to the special issue on computational archival science. Journal on Computing and Cultural Heritage (JOCCH), 15(1): 1:1–1:2, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3495004.

Heitman:2017:ITL

[HMP17] Carrie Heitman, Worthy Martin, and Stephen Plog. Innovation through largescale integration of legacy records: Assessing the “Value Added” in cultural heritage resources. Journal on Computing and Cultural Heritage (JOCCH), 10(3):17:1–17:??, August 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hammady:2020:AIV

[HMS20] Ramy Hammady, Minhua Ma, and Carl Strathearn. Ambient information visualisation and visitors’ technology acceptance of mixed reality in museums. Journal on Computing and Cultural Heritage (JOCCH), 13(2): 9:1–9:22, June 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3359590.

Hu:2017:EMG

[HOGP17] Rui Hu, Jean-Marc Odobez, and Daniel Gatica-Perez. Extracting Maya glyphs from degraded ancient documents via image segmentation. Journal on Computing and Cultural Heritage (JOCCH), 10(2):10:1– 10:??, April 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hess:2016:DIP

[HRS^{+}16] Mona Hess, Stuart Robson, Margaret Serpico, Giancarlo Amati, Ivor Pridden, and Tonya Nelson. Developing 3D imaging programmes–workflow and quality control. Journal on Computing and Cultural Heritage (JOCCH), 9(1): 1:1–1:??, February 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hutchings:2008:USA

[HS08] Jeremy Hutchings and Jonathan Ashley Smith. Using survival analysis on conservation metadata to benchmark treatment frequency. Journal on Computing and Cultural Heritage (JOCCH), 1(2):9:1– 9:??, October 2008. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Havemann:2009:ASR

[HSB^{+}09] Sven Havemann, Volker Settgast, René Berndt, Øyvind Eide, and Dieter W. Fellner. The Arrigo Showcase reloaded — towards a sustainable link between 3D and semantics. Journal on Computing and Cultural Heritage (JOCCH), 2(1):4:1– 4:??, July 2009. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hidalgo-Sanchez:2022:GSA

[HSMAK^{+}22] Francisco M. Hidalgo-Sánchez, Emilio J. Mascort-Albea, Martin Kada, Rocío Romero-Hernández, Jacinto Canivell, and Francisco López-Larrínaga. 3D GIS semi-automatized modelling procedure for the conservation of the PHiM: Heritage municipal buildings of Seville (Spain). A new dimension for urban cultural data management. Journal on Computing and Cultural Heritage (JOCCH), 15(1):17:1– 17:25, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3467976.

Hu:2024:IIV

[Hu24] Jia Hu. Individually integrated virtual/augmented reality environment for interactive perception of cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 17(1): 9:1–9:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631145.

Heerlien:2015:NHP

[HVS^{+}15] Maarten Heerlien, Joost Van Leusen, Stephanie Schnörr, Suzanne De Jong-Kole, Niels Raes, and Kirsten Van Hulsen. The natural history production line: an industrial approach to the digitization of scientific collections. Journal on Computing and Cultural Heritage (JOCCH), 8(1): 3:1–3:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Holowko:2014:CBA

[HWSK14] Elwira Holowko, Jerzy Wojsz, Robert Sitnik, and Maciej Karaszewski. Colorbased algorithm for automatic merging of multiview 3D point clouds. Journal on Computing and Cultural Heritage (JOCCH), 7(3): 16:1–16:??, June 2014. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Hou:2024:ETD

[HZ24] Jingrui Hou and Shitou Zhang. Exploring thematic diversity in classical Chinese poetry: a novel dataset and a BERTenhanced ensemble learning approach. Journal on Computing and Cultural Heritage (JOCCH), 17(4):60:1– 60:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3685679.

Isemann:2014:OAI

[IA14] Daniel Isemann and Khurshid Ahmad. Ontological access to images of fine art. Journal on Computing and Cultural Heritage (JOCCH), 7(1):3:1– 3:??, April 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ibrahim:2018:CFD

[IA18] Nazrita Ibrahim and Nazlena Mohamad Ali. A conceptual framework for designing virtual heritage environment for cultural learning. Journal on Computing and Cultural Heritage (JOCCH), 11(2):11:1– 11:??, June 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ibrahim:2015:FFC

[IAY15] Nazrita Ibrahim, Nazlena Mohamad Ali, and Noor Faezah Mohd Yatim. Factors facilitating cultural learning in virtual architectural heritage environments: End user perspective. Journal on Computing and Cultural Heritage (JOCCH), 8(2):8:1– 8:??, March 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ioannidis:2021:WDY

[IEL21] Petros Ioannidis, Lina Eklund, and Anders Sundnes Løvlie. We dare you: a lifecycle study of a substitutional reality installation in a museum space. Journal on Computing and Cultural Heritage (JOCCH), 14(3): 34:1–34:21, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3439862.

Iudova-Romanova:2022:VRC

[IRHH^{+}22] Kateryna Iudova-Romanova, Tetiana Humenyuk, Serhii Horevalov, Serhii Honcharuk, and Volodymyr Mykhalov. Virtual reality in contemporary theatre. Journal on Computing and Cultural Heritage (JOCCH), 15(4):75:1– 75:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3524024.

Jimenez-Diaz:2024:ICR

[JDDAB^{+}24] Guillermo Jiménez-Díaz, Belen Diaz-Agudo, Luis Emilio Bruni, Nele Kadastik, Anna Follo, Rossana Damiano, Manuel Striani, Angel Sanchez-Martin, and Antonio Lieto. Interpretable clusters for representing citizens’ sense of belonging through interaction with cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 17(4):74:1– 74:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3665142.

Jang:2022:CCS

[JK22] Sun-Young Jang and Sung-Ah Kim. Content curation for spatial experience of architectural heritage. Journal on Computing and Cultural Heritage (JOCCH), 15(4):69:1– 69:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3513136.

Jang:2023:AEL

[JK23] Sun-Young Jang and Sung-Ah Kim. Automatic extraction and linkage between textual and spatial data for architectural heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(3):48:1– 48:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3586158.

Jin:2024:CSH

[JML24] Yunshui Jin, Minhua Ma, and Yun Liu. Comparative study of HMDbased virtual and augmented realities for immersive museums: User acceptance, medium, and learning. Journal on Computing and Cultural Heritage (JOCCH), 17(1):13:1– 13:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3627164.

Jones:2019:EIS

[JTL19] Catherine Emma (Kate)

Jones, Stathis Theodosis, and Ioanna Lykourentzou. The enthusiast, the interested, the sceptic, and the cynic: Understanding user experience and perceived value in locationbased cultural heritage games through qualitative and sentiment analysis. Journal on Computing and Cultural Heritage (JOCCH), 12(1):6:1– 6:??, February 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3297716.

Jaspe-Villanueva:2021:WBE

[JVAP^{+}21] Alberto Jaspe-Villanueva, Moonisa Ahsan, Ruggero Pintus, Andrea Giachetti, Fabio Marton, and Enrico Gobbetti. Web-based exploration of annotated multi-layered relightable image models. Journal on Computing and Cultural Heritage (JOCCH), 14(2): 24:1–24:29, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3430846.

Jiang:2023:TMA

[JZF23] Yicheng Jiang, Xia Zheng, and Chao Feng. Toward multi-area contactless museum visitor counting with commodity WiFi. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 8:1–8:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3530694.

Kramer:2020:PAC

[KA20] Madison Kramer and Ergun Akleman. A procedural approach to creating American second empire houses. Journal on Computing and Cultural Heritage (JOCCH), 13(1):1:1– 1:19, February 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3343196.

Khamjane:2018:GIQ

[KB18] Aziz Khamjane and Rachid Benslimane. Generating Islamic quasi-periodic patterns: a new method. Journal on Computing and Cultural Heritage (JOCCH), 11(3):13:1–13:??, September 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Konstantakis:2020:ACU

[KC20] Markos Konstantakis and George Caridakis. Adding culture to UX: UX research methodologies and applications in cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 13(1):4:1– 4:17, February 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3354002.

Kim:2024:ARB

[KCJP24] Eunhoo Kim, Junyeop Cha, Dahye Jeong, and Eunil Park. Artwork recommendations based on user preferences: Integrating clustering analysis with visual features. Journal on Computing and Cultural Heritage (JOCCH), 17(3):38:1– 38:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3649901.

Koho:2023:WMP

[KCRE23] Mikko Koho, L. P. Coladangelo, Lynn Ransom, and Doug Emery. Wikibase model for premodern manuscript metadata harmonization, linked data integration, and discovery. Journal on Computing and Cultural Heritage (JOCCH), 16(3):56:1– 56:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3594723.

Kenderdine:2014:PLF

[KCS14] Sarah Kenderdine, Leith K. Y. Chan, and Jeffrey Shaw. Pure Land: Futures for embodied museography. Journal on Computing and Cultural Heritage (JOCCH), 7(2):8:1– 8:??, July 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Kroher:2016:CCR

[KDBMG16] Nadine Kroher, José-Miguel Díaz-Báñez, Joaquin Mora, and Emilia Gómez. Corpus COFLA: a research corpus for the computational study of flamenco music. Journal on Computing and Cultural Heritage (JOCCH), 9(2): 10:1–10:??, May 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Kuzio:2023:CPS

[KF23] Olivia Kuzio and Susan Farnand. Comparing practical spectral imaging methods for cultural heritage studio photography. Journal on Computing and Cultural Heritage (JOCCH), 16(1):11:1– 11:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3531019.

Koller:2009:RCD

[KFH09] David Koller, Bernard Frischer, and Greg Humphreys. Research challenges for digital archives of 3D cultural heritage models. Journal on Computing and Cultural Heritage (JOCCH), 2(3): 7:1–7:??, December 2009. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Koolen:2023:VPS

[KHO^{+}23] Marijn Koolen, Rik Hoekstra, Joris Oddens, Ronald Sluijter, Rutger Van Koert, Gijsjan Brouwer, and Hennie Brugman. The value of preexisting structures for digital access: Modelling the resolutions of the Dutch States General. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 1:1–1:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3575864.

Kim:2022:NMM

[Kim22] Minseo Kim. New method for museum archiving: “Quantitative Analysis Meets Art History”. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 63:1–63:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3531018.

Kim:2024:ODH

[KIM24] Yongho Kim, Chanjong Im, and Thomas Mandl. Object detection in historical images: Transfer learning and pseudo labelling. Journal on Computing and Cultural Heritage (JOCCH), 17(4):58:1– 58:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3699963.

Koo:2020:DAR

[KKK^{+}20] Seungbum Koo, Jinyoung Kim, Changhyuk Kim, Jeeyeop Kim, and Hee Sung Cha. Development of an augmented reality tour guide for a cultural heritage site. Journal on Computing and Cultural Heritage (JOCCH), 12(4):24:1– 24:??, January 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3317552.

Kumar:2022:RTU

[KKRS22] Abhishek Kumar, Ankit Kumar, Linesh Raja, and Kamred Udham Singh. Rediscovering the traditional UNESCO World Heritage Hawamahal through 3D animation and immersive technology. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 61:1–61:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3524023.

Krumpen:2022:TTC

[KKW22] Stefan Krumpen, Reinhard Klein, and Michael Weinmann. Towards tangible cultural heritage experiences-enriching VRbased object inspection with haptic feedback. Journal on Computing and Cultural Heritage (JOCCH), 15(1):19:1–19:17, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3470470.

Kolomenkin:2013:RRO

[KLST13] Michael Kolomenkin, George Leifman, Ilan Shimshoni, and Ayellet Tal. Reconstruction of relief objects from archeological line drawings. Journal on Computing and Cultural Heritage (JOCCH), 6(1):3:1– 3:??, March 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

King:2022:SAR

[KML22] Elise King, Katie Pierce Meyer, and King-Ip (David) Lin. Semi-automatic residential floor plan detection: Developing a tool for humanities research. Journal on Computing and Cultural Heritage (JOCCH), 15(4):67:1– 67:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3503046.

Kocaturk:2022:GIE

[KMMK22] Tuba Kocaturk, Domenico Mazza, Malcolm McKinnon, and Sofija Kaljevic. GDOM: an immersive experience of intangible heritage through spatial storytelling. Journal on Computing and Cultural Heritage (JOCCH), 15(4):59:1– 59:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3498329.

Kumar:2020:DDA

[KOIC20] Pakhee Kumar, Ferda Ofli, Muhammad Imran, and Carlos Castillo. Detection of disaster-affected cultural heritage sites from social media images using deep learning techniques. Journal on Computing and Cultural Heritage (JOCCH), 13(3):23:1– 23:31, October 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3383314.

Katifori:2020:LTT

[KPV^{+}20] Akrivi Katifori, Sara Perry, Maria Vayanou, Angeliki Antoniou, Ioannis-Panagiotis Ioannidis, Sierra McKinney, Angeliki Chrysanthi, and Yannis Ioannidis. “Let Them Talk!”: Exploring guided group interaction in digital storytelling experiences. Journal on Computing and Cultural Heritage (JOCCH), 13(3): 21:1–21:30, October 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3382773.

Kim:2014:HGS

[KRF^{+}14] Min H. Kim, Holly Rushmeier, John Ffrench, Irma Passeri, and David Tidmarsh. Hyper3D: 3D graphics software for examining cultural artifacts. Journal on Computing and Cultural Heritage (JOCCH), 7(3):1:1–1:??, February 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Kale:2023:DDL

[KRJ23] Manjeeta R. Kale, Priti P. Rege, and Radhika D. Joshi. Designing a duallevel facial expression evaluation system for performers using geometric features and Petri nets. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 31:1–31:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3583557.

Kim:2024:SAA

[KSSW24] Hayun Kim, Maryam Shakeri, Jae-Eun Shin, and Woontack Woo. Spaceadaptive artwork placement based on content similarities for curating thematic spaces in a virtual museum. Journal on Computing and Cultural Heritage (JOCCH), 17(1):3:1– 3:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631134.

Kuflik:2011:VGA

[KSZ^{+}11] Tsvi Kuflik, Oliviero Stock, Massimo Zancanaro, Ariel Gorfinkel, Sadek Jbara, Shahar Kats, Julia Sheidin, and Nadav Kashtan. A visitor’s guide in an active museum: Presentations, communications, and reflection. Journal on Computing and Cultural Heritage (JOCCH), 3(3):11:1– 11:??, March 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Knight:2016:UGI

[KT16] Gregory P. Knight and Nasseh Tabrizi. Using n-grams to identify time periods of cultural influence. Journal on Computing and Cultural Heritage (JOCCH), 9(3):15:1– 15:??, November 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Katsouri:2015:VAH

[KTHP15] Irene Katsouri, Aimilia Tzanavari, Kyriakos Herakleous, and Charalambos Poullis. Visualizing and assessing hypotheses for marine archaeology in a VR CAVE environment. Journal on Computing and Cultural Heritage (JOCCH), 8 (2):10:1–10:??, March 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Kwon:2024:RTM

[KY24] Do Hyung Kwon and Jeong Min Yu. Real-time multi-CNN-based emotion recognition system for evaluating museum visitors’ satisfaction. Journal on Computing and Cultural Heritage (JOCCH), 17(1): 15:1–15:??, March 2024. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3631123.

Kavelar:2014:RLR

[KZK14] Albert Kavelar, Sebastian Zambanini, and Martin Kampel. Reading the legends of Roman Republican coins. Journal on Computing and Cultural Heritage (JOCCH), 7(1):5:1– 5:??, April 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Liu:2024:EPE

[LA24] Shao Liu and Sos S. Agaian. 3DEmo: For portrait emotion recognition with new dataset. Journal on Computing and Cultural Heritage (JOCCH), 17(2): 17:1–17:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631133.

Laycock:2015:UCM

[LBC^{+}15] S. D. Laycock, G. D. Bell, N. Corps, D. B. Mortimore, G. Cox, S. May, and I. Finkel. Using a combination of micro–computed tomography, CAD and 3D printing techniques to reconstruct incomplete 19th-Century Cantonese chess pieces. Journal on Computing and Cultural Heritage (JOCCH), 7(4):25:1– 25:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Laycock:2012:CXR

[LBM^{+}12] Stephen D. Laycock, Graham D. Bell, David B. Mortimore, Mark K. Greco, Nick Corps, and Irving Finkle. Combining X-ray micro-CT technology and 3D printing for the digital preservation and study of a 19th Century Cantonese chess piece with intricate internal structure. Journal on Computing and Cultural Heritage (JOCCH), 5(4):13:1–13:??, December 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Lopez:2021:EKT

[LC21] Gonzalo Andrés López and David Checa Cruz. Experiences of knowledge transfer on industrial heritage using games, storytelling, and new technologies: “A History of Enterprises”. Journal on Computing and Cultural Heritage (JOCCH), 14(2):19:1– 19:26, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3424951.

Leoni:2015:DCS

[LCD^{+}15] Chiara Leoni, Marco Callieri, Matteo Dellepiane, Daniel Paul O’Donnell, Roberto Rosselli Del Turco, and Roberto Scopigno. The dream and the cross: a 3D scanning project to bring 3D content in a digital edition. Journal on Computing and Cultural Heritage (JOCCH), 8(1):5:1– 5:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Laycock:2008:ECH

[LDD08] R. G. Laycock, D. Drinkwater, and A. M. Day. Exploring cultural heritage sites through space and time. Journal on Computing and Cultural Heritage (JOCCH), 1(2):11:1– 11:??, October 2008. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Liebeskind:2016:SCC

[LDS16] Chaya Liebeskind, Ido Dagan, and Jonathan Schler. Semiautomatic construction of cross-period thesaurus. Journal on Computing and Cultural Heritage (JOCCH), 9(4):22:1– 22:??, December 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Lee:2019:CHI

[LhL19] Woosik Lee and Dong hoon Lee. Cultural heritage and the intelligent Internet of Things. Journal on Computing and Cultural Heritage (JOCCH), 12(3):21:1– 21:??, October 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3316414.

Lange:2020:PPD

[LKK^{+}20] Vanessa Lange, Philipp Kurth, Benjamin Keinert, Martin Boss, Marc Stamminger, and Frank Bauer. Proxy painting: Digital colorization of real-world objects. Journal on Computing and Cultural Heritage (JOCCH), 13(3):16:1– 16:20, October 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3377145.

Lee:2019:NPB

[LLA^{+}19] Yong Yi Lee, Jong Hun Lee, Bilal Ahmed, Moon Gu Son, and Kwan H. Lee. A new projection-based exhibition system for a museum. Journal on Computing and Cultural Heritage (JOCCH), 12(2):10:1– 10:??, June 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3275522.

Liebeskind:2024:MTH

[LLB24] Chaya Liebeskind, Shmuel Liebeskind, and Dan Bouhnik. Machine translation for historical research: a case study of Aramaic– Ancient Hebrew translations. Journal on Computing and Cultural Heritage (JOCCH), 17(2):20:1– 20:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3627168.

LoBuglio:2015:WDT

[LLD15] David Lo Buglio, Vanessa Lardinois, and Livio De Luca. What do thirtyone columns say about a “Theoretical” thirtysecond? Journal on Computing and Cultural Heritage (JOCCH), 8(1):6:1– 6:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Lepouras:2021:ISI

[LLL21] George Lepouras, Ioanna Lykourentzou, and Antonios Liapis. Introduction to the special issue on “Culture Games”. Journal on Computing and Cultural Heritage (JOCCH), 14(2): 11:1–11:3, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3453690.

Liu:2024:PDH

[LLL^{+}24] Liu Liu, Chufei Liu, Wenqi Li, Dongbo Wang, and Shuiqing Huang. Perspective of digital humanities on person names in Chinese pre-Qin classic. Journal on Computing and Cultural Heritage (JOCCH), 17(2): 32:1–32:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3647997.

Lu:2012:IBA

[LLS^{+}12] Zheng Lu, Wei Luo, Zhijun Sun, Moshe Ben-Ezra, and Michael S. Brown. Imaging Buddhist art with a digital large-format camera: a field study report from the Dunhuang Caves. Journal on Computing and Cultural Heritage (JOCCH), 5 (3):9:1–9:??, October 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Liu:2024:HCG

[LMWP24] Jiangfeng Liu, Xueliang Ma, Lanyu Wang, and Lei Pei. How can generative artificial intelligence techniques facilitate intelligent research into ancient books? Journal on Computing and Cultural Heritage (JOCCH), 17(4):57:1– 57:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3690391.

Lin:2017:WSA

[LMWZ17] Jimmy Lin, Ian Milligan, Jeremy Wiebe, and Alice Zhou. Warcbase: Scalable analytics infrastructure for exploring Web archives. Journal on Computing and Cultural Heritage (JOCCH), 10(4):22:1– 22:??, October 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Lombardo:2016:SAD

[LPD16] Vincenzo Lombardo, Antonio Pizzo, and Rossana Damiano. Safeguarding and accessing drama as intangible cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 9(1):5:1– 5:??, February 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Lampe:2008:RBN

[LRD08] Karl-Heinz Lampe, Klaus Riede, and Martin Doerr. Research between natural and cultural history information: Benefits and ITrequirements for transdisciplinarity. Journal on Computing and Cultural Heritage (JOCCH), 1(1):4:1– 4:??, June 2008. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Lovlie:2021:PGT

[LRS^{+}21] Anders Sundnes Løvlie, Karin Ryding, Jocelyn Spence, Paulina Rajkowska, Annika Waern, Tim Wray, Steve Benford, William Preston, and Emily Clare-Thorn. Playing games with Tito: Designing hybrid museum experiences for critical play. Journal on Computing and Cultural Heritage (JOCCH), 14(2):16:1– 16:26, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3446620.

Lucifora:2024:VEU

[LSA^{+}24] Chiara Lucifora, Massimiliano Schembri, Luigi Asprino, Anna Follo, and Aldo Gangemi. Virtual empathy: Usability and immersivity of a VR system for enhancing social cohesion through cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 17(4):69:1– 69:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3679019.

Lewis:2021:HTT

[LTP21] Richard Lewis and Molly Taylor-Poleskey. Hidden town in 3D: Teaching and reinterpreting slavery virtually at a living history museum. Journal on Computing and Cultural Heritage (JOCCH), 14(2):26:1– 26:14, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3431924.

Liu:2016:RBG

[LW16] Jin Liu and Zhong-Ke Wu.

Rule-based generation of ancient Chinese architecture from the Song Dynasty. Journal on Computing and Cultural Heritage (JOCCH), 9(2):7:1– 7:??, May 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Liu:2024:SGP

[LWZ^{+}24] Chang Liu, Dongbo Wang, Zhixiao Zhao, Die Hu, Mengcheng Wu, Litao Lin, Jiangfeng Liu, Hai Zhang, Si Shen, Bin Li, and Lianzhen Zhao. SikuGPT: a generative pre-trained model for intelligent information processing of ancient texts from the perspective of digital humanities. Journal on Computing and Cultural Heritage (JOCCH), 17(4):53:1– 53:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3676969.

Lin:2024:DRA

[LXW24] Xiaohu Lin, Bei Xue, and Xiqi Wang. Digital 3D reconstruction of ancient Chinese Great Wild Goose Pagoda by TLS point cloud hierarchical registration. Journal on Computing and Cultural Heritage (JOCCH), 17(2):23:1– 23:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3639932.

Matthews:2012:IME

[MA12] Paul Matthews and Judith Aston. Interactive multimedia ethnography: Archiving workflow, interface aesthetics and metadata. Journal on Computing and Cultural Heritage (JOCCH), 5(4):14:1–14:??, December 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Musso:2024:WUC

[MANC24] Marta Musso, Kerstin Arnold, Federico Nanni, and Beatrice Cannelli. What is in a < unitttitle>? cross-lingual topic detection & information retrieval in archives portal Europe. Journal on Computing and Cultural Heritage (JOCCH), 17(2):25:1– 25:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3494572.

Mason:2016:MGP

[Mas16] Marco Mason. The MIT Museum Glassware prototype: Visitor experience exploration for designing smart glasses. Journal on Computing and Cultural Heritage (JOCCH), 9(3):

12:1–12:??, November 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Mortara:2013:ISI

[MB13] Michela Mortara and Francesco Bellotti. Introduction to special issue on serious games for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 6(2):6:1– 6:??, May 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Mori:2013:EAD

[MBD^{+}13] Daniele Mori, Riccardo Berta, Alessandro De Gloria, Valentina Fiore, and Lauto Magnani. An easy to author dialogue management system for serious games. Journal on Computing and Cultural Heritage (JOCCH), 6(2):10:1– 10:??, May 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Meklati:2023:SDI

[MBEB23] Safia Meklati, Kenza Boussora, Mohamed El Hafedh Abdi, and Sid-Ahmed Berrani. Surface damage identification for heritage site protection: a mobile crowd-sensing solution based on deep learning. Journal on Computing and Cultural Heritage (JOCCH), 16(2):25:1– 25:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3569093.

Michalakis:2022:CAC

[MC22] Konstantinos Michalakis and George Caridakis. Context awareness in cultural heritage applications: a survey. Journal on Computing and Cultural Heritage (JOCCH), 15(2):32:1– 32:31, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3480953.

Mallik:2011:NPI

[MCG11] Anupama Mallik, Santanu Chaudhury, and Hiranmay Ghosh. Nrityakosha: Preserving the intangible heritage of Indian classical dance. Journal on Computing and Cultural Heritage (JOCCH), 4(3):11:1– 11:??, December 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Moraitou:2024:OBF

[MCKC24] Efthymia Moraitou, Yannis Christodoulou, Konstantinos Kotis, and George Caridakis. An ontology-based framework for supporting decision-making in conservation and restoration interventions for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 17(3):41:1– 41:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3653977.

Milleville:2023:AES

[MCV23] Kenzo Milleville, Krishna Kumar Thirukokaranam Chandrasekar, and Steven Verstockt. Automatic extraction of specimens from multi-specimen herbaria. Journal on Computing and Cultural Heritage (JOCCH), 16(1):4:1–4:??, March 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3575862.

Makridis:2012:ACA

[MD12] Michael Makridis and Petros Daras. Automatic classification of archaeological pottery sherds. Journal on Computing and Cultural Heritage (JOCCH), 5(4): 15:1–15:??, December 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Marais:2019:SAC

[MDCS19] Patrick Marais, Matteo Dellepiane, Paolo Cignoni, and Roberto Scopigno.

Semi-automated cleaning of laser scanning campaigns with machine learning. Journal on Computing and Cultural Heritage (JOCCH), 12(3):16:1– 16:??, October 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3292027.

Milani:2021:DCM

[MF21] Federico Milani and Piero Fraternali. A dataset and a convolutional model for iconography classification in paintings. Journal on Computing and Cultural Heritage (JOCCH), 14(4): 46:1–46:18, December 2021. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3458885.

Muller:2010:PDO

[MFRG10] Florian Müller, Peter Fornaro, Lukas Rosenthaler, and Rudolf Gschwind. PE-VIAR: Digital originals. Journal on Computing and Cultural Heritage (JOCCH), 3(1):2:1–2:??, June 2010. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Migliorini:2018:IST

[MGB18] Sara Migliorini, Piergiovanna Grossi, and Alberto Belussi. An interoperable spatio-temporal model for archaeological data based on ISO Standard 19100. Journal on Computing and Cultural Heritage (JOCCH), 11(1):5:1– 5:??, January 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Manitsaris:2014:CMR

[MGBM14] S. Manitsaris, A. Glushkova, F. Bevilacqua, and F. Moutarde. Capture, modeling, and recognition of expert technical gestures in wheelthrowing art of pottery. Journal on Computing and Cultural Heritage (JOCCH), 7(2):10:1–10:??, June 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Merry:2013:FPB

[MGM13] Bruce Merry, James Gain, and Patrick Marais. Fast inplace binning of laser rangescanned point sets. Journal on Computing and Cultural Heritage (JOCCH), 6(3): 14:1–14:??, July 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Meinecke:2022:TEV

[MHJ22] Christofer Meinecke, Chris Hall, and Stefan Jänicke. Towards enhancing virtual museums by contextualizing art through interactive visualizations. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 62:1–62:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3527619.

Mimno:2012:CHD

[Mim12] David Mimno. Computational historiography: Data mining in a century of classics journals. Journal on Computing and Cultural Heritage (JOCCH), 5(1): 3:1–3:??, April 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Milosz:2021:SVL

[MKM21] Marek Milosz, Jacek Kesik, and Jerzy Montusiewicz. 3D scanning and visualization of large monuments of Timurid architecture in Central Asia — a methodical approach. Journal on Computing and Cultural Heritage (JOCCH), 14(1):8:1– 8:31, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3425796.

May:2021:CMM

[MKZ21] Michael J. May, Efrat Kantor, and Nissim Zror. CemoMemo: Making more out of gravestones (with help from the crowd). Journal on Computing and Cultural Heritage (JOCCH), 14(4):57:1–57:22, December 2021. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3467888.

Morse:2022:MAD

[MLL^{+}22] Christopher Morse, Blandine Landau, Carine Lallemand, Lars Wieneke, and Vincent Koenig. From #MuseumAtHome to #AtHomeAtTheMuseum: Digital museums and dialogical engagement beyond the COVID-19 pandemic. Journal on Computing and Cultural Heritage (JOCCH), 15(2):30:1– 30:29, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3480955.

Madsen:2016:HVR

[MM16] Jacob B. Madsen and Claus B. Madsen. Handheld visual representation of a castle chapel ruin. Journal on Computing and Cultural Heritage (JOCCH), 9(1): 6:1–6:??, February 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Montusiewicz:2021:AJL

[MM21] Jerzy Montusiewicz and Marek Milosz. Architectural jewels of Lublin: a modern computerized board game in cultural heritage education. Journal on Computing and Cultural Heritage (JOCCH), 14(3): 35:1–35:21, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3446978.

Maalek:2023:ARD

[MM23] Reza Maalek and Shahrokh Maalek. Automatic recognition and digital documentation of cultural heritage hemispherical domes using images. Journal on Computing and Cultural Heritage (JOCCH), 16(1):6:1– 6:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3528412.

Mesanza-Moraza:2021:MLB

[MMGGA21] Amaia Mesanza-Moraza, Ismael García-Gómez, and Agustín Azkarate. Machine learning for the Built Heritage Archaeological Study. Journal on Computing and Cultural Heritage (JOCCH), 14(1):10:1– 10:21, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3422993.

Morse:2021:CLR

[MNL^{+}21] Christopher Morse, Jasmin Niess, Carine Lallemand, Lars Wieneke, and Vincent Koenig. Casual leisure in rich-prospect: Advancing visual information behavior for digital museum collections. Journal on Computing and Cultural Heritage (JOCCH), 14(3):30:1– 30:23, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3437257.

Moritz:2016:SSM

[MPFC16] Maria Moritz, Barbara Pavlek, Greta Franzini, and Gregory Crane. Sentence shortening via morphosyntactic annotated data in historical language learning. Journal on Computing and Cultural Heritage (JOCCH), 9(1):3:1– 3:??, February 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Munoz-Pandiella:2024:FBA

[MPPB24] Imanol Munoz-Pandiella, Xavier Pueyo, and Carles Bosch. A 3D feature-based approach for mapping scaling effects on stone monuments. Journal on Computing and Cultural Heritage (JOCCH), 17(2):33:1– 33:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3651988.

Mirri:2017:HNH

[MPRS17] Silvia Mirri, Catia Prandi, Marco Roccetti, and Paola Salomoni. Handmade narrations: Handling digital narrations on food and gastronomic culture. Journal on Computing and Cultural Heritage (JOCCH), 10(4): 20:1–20:??, October 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Migliorini:2022:TDP

[MQB22] Sara Migliorini, Elisa Quintarelli, and Alberto Belussi. Tracking data provenance of archaeological temporal information in presence of uncertainty. Journal on Computing and Cultural Heritage (JOCCH), 15(2):31:1– 31:32, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3480956.

Murano:2023:DIA

[MQD^{+}23] Francesca Murano, Valeria Quochi, Angelo Mario Del Grosso, Luca Rigobianco, and Mariarosaria Zinzi. Describing inscriptions of Ancient Italy. The ItAnt project and its information encoding process. Journal on Computing and Cultural Heritage (JOCCH), 16(3):53:1– 53:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3606703.

Moussouri:2015:CVS

[MR15] Theano Moussouri and George Roussos. Conducting visitor studies using Smartphone-based location sensing. Journal on Computing and Cultural Heritage (JOCCH), 8(3):12:1– 12:??, May 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Maree:2021:MMS

[MRAB21] Mohammed Maree, Amjad Rattrout, Muhanad Altawil, and Mohammed Belkhatir. Multi-modality search and recommendation on Palestinian cultural heritage based on the Holy-Land ontology and extrinsic semantic resources. Journal on Computing and Cultural Heritage (JOCCH), 14(3): 29:1–29:23, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3447523.

Marton:2014:IIV

[MRB^{+}14] Fabio Marton, Marcos Balsa Rodriguez, Fabio Bettio, Marco Agus, Alberto Jaspe Villanueva, and Enrico Gobbetti. IsoCam: Interactive visual exploration of massive cultural heritage models on large projection setups. Journal on Computing and Cultural Heritage (JOCCH), 7(2):12:1– 12:??, July 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Mazzamuto:2024:LDA

[MRFF24] Michele Mazzamuto, Francesco Ragusa, Antonino Furnari, and Giovanni Maria Farinella. Learning to detect attended objects in cultural sites with gaze signals and weak object supervision. Journal on Computing and Cultural Heritage (JOCCH), 17(3):35:1– 35:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3647999.

Mv:2013:CFP

[MSN^{+}13] Rohith Mv, Gowri Somanath, Debra Norris, Jennifer Gutierrez, and Chandra Kambhamettu. A camera flash projector-based reconstruction system for digital preservation of artifacts. Journal on Computing and Cultural Heritage (JOCCH), 6(1):5:1– 5:??, March 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Meghini:2017:ARI

[MSR^{+}17] Carlo Meghini, Roberto Scopigno, Julian Richards, Holly Wright, Guntram Geser, Sebastian Cuy, Johan Fihn, Bruno Fanini, Hella Hollander, Franco Niccolucci, Achille Felicetti, Paola Ronzino, Federico Nurra, Christos Papatheodorou, Dimitris Gavrilis, Maria Theodoridou, Martin Doerr, Douglas Tudhope, Ceri Binding, and Andreas Vlachidis. ARIADNE: a research infrastructure for archaeology. Journal on Computing and Cultural Heritage (JOCCH), 10(3):18:1– 18:??, August 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Milotta:2018:MCS

[MSTG18] Filippo L. M. Milotta, Filippo Stanco, Davide Tanasi, and Anna M. Gueli. Munsell color specification using ARCA (Automatic Recognition of Color for Archaeology). Journal on Computing and Cultural Heritage (JOCCH), 11(4): 17:1–17:??, December 2018.

CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3216463.

Mccarthy:2020:OWP

[MSWS20] J. Mccarthy, E. Sebo, B. Wilkinson, and F. Sheehan. Open workflows for polychromatic reconstruction of historical sculptural monuments in 3D. Journal on Computing and Cultural Heritage (JOCCH), 13(3): 18:1–18:16, October 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3386314.

Munster:2019:DHS

[Mün19] Sander Münster. Digital heritage as a scholarly fieldtopics, researchers, and perspectives from a bibliometric point of view. Journal on Computing and Cultural Heritage (JOCCH), 12(3):22:1–22:??, October 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3310012.

Madhu:2023:EHP

[MVCK^{+}23] Prathmesh Madhu, Angel Villar-Corrales, Ronak Kosti, Torsten Bendschus, Corinna Reinhardt, Peter Bell, Andreas Maier, and Vincent Christlein. Enhancing human pose estimation in ancient vase paintings via perceptuallygrounded style transfer learning. Journal on Computing and Cultural Heritage (JOCCH), 16(1):16:1– 16:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3569089.

Miles:2016:ARR

[MWL^{+}16] Helen C. Miles, Andrew T. Wilson, Frédéric Labrosse, Bernard Tiddeman, Seren Griffiths, Ben Edwards, Panagiotis D. Ritsos, Joseph W. Mearman, Katharina Möller, Raimund Karl, and Jonathan C. Roberts. Alternative representations of 3Dreconstructed heritage data. Journal on Computing and Cultural Heritage (JOCCH), 9(1):4:1–4:??, February 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ma:2012:ATC

[MWX^{+}12] Wei Ma, Yizhou Wang, Ying-Qing Xu, Qiong Li, Xin Ma, and Wen Gao. Annotating traditional Chinese paintings for immersive virtual exhibition. Journal on Computing and Cultural Heritage (JOCCH), 5(2):6:1– 6:??, July 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Melnik:2022:DSC

[MYS22] Gil Melnik, Yuval Yekutieli, and Andrei Sharf. Deep segmentation of corrupted glyphs. Journal on Computing and Cultural Heritage (JOCCH), 15(1):13:1– 13:24, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3465629.

Nasri:2017:PSG

[NB17] Abdelbar Nasri and Rachid Benslimane. Parametric shape grammar formalism for Moorish geometric design analysis and generation. Journal on Computing and Cultural Heritage (JOCCH), 10(4):19:1– 19:??, October 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Not:2019:DAH

[NCM^{+}19] Elena Not, Dario Cavada, Stefano Maule, Anna Pisetti, and Adriano Venturini. Digital augmentation of historical objects through tangible interaction. Journal on Computing and Cultural Heritage (JOCCH), 12(3):18:1–18:??, October 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3297764.

Neamtu:2024:EUE

[NCP^{+}24] Calin Neamtu, Radu Comes, Dorin-Mircea Popovici, Elena Bautu, Mateescu-Suciu Liliana, Adrien Syrotnik, and Matei-Ioan Popovici. Evaluating user experience in the context of cultural heritage dissemination using extended reality: a case study of the Dacian bronze matrix with hollow design. Journal on Computing and Cultural Heritage (JOCCH), 17(2):19:1– 19:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3639933.

Ngo:2021:ECA

[NDN^{+}21] Vuong M. Ngo, Thuy-Van T. Duong, Tat-Bao-Thien Nguyen, Phuong T. Nguyen, and Owen Conlan. An efficient classification algorithm for traditional textile patterns from different cultures based on structures. Journal on Computing and Cultural Heritage (JOCCH), 14(4):53:1– 53:22, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465381.

Niccolucci:2018:E

[Nic18] Franco Niccolucci. Editorial. Journal on Computing and Cultural Heritage (JOCCH), 11(4):17:1– 17:??, December 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3292026.

Niccolucci:2019:APC

[Nic19a] Franco Niccolucci. Addendum to “From Point Cloud Data to Structural Analysis by a Geometrical hBIM-Oriented Model”. Journal on Computing and Cultural Heritage (JOCCH), 12(3):23:1–23:??, October 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3359625.

See [RAB^{+}19].

Niccolucci:2019:MMD

[Nic19b] Karina Rodriguez Echavarria / Franco Niccolucci. In memoriam, Matteo Dellepiane (1979–2019). Journal on Computing and Cultural Heritage (JOCCH), 12(1): 1:1–1:??, February 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3308556.

Niccolucci:2021:EPV

[Nic21] Franco Niccolucci. Editorial: For pursuit of virtue and of knowledge. Journal on Computing and Cultural Heritage (JOCCH), 14(3): 27e:1–27e:3, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465627.

Ni:2011:HIS

[NJDZ11] Enzhi Ni, Minjun Jiang, Xiaojun Ding, and Changle Zhou. Handwriting input system of Chinese guqin notation. Journal on Computing and Cultural Heritage (JOCCH), 3(3):9:1– 9:??, March 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Noll:2015:FAO

[NKRS15] Tobias Nöll, Johannes Köhler, Gerd Reis, and Didier Stricker. Fully automatic, omnidirectional acquisition of geometry and appearance in the context of cultural heritage preservation. Journal on Computing and Cultural Heritage (JOCCH), 8(1):2:1– 2:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Niang:2017:SSI

[NMM^{+}17] Cheikh Niang, Claudia Marinica, Beatrice Markhoff, Elise Leboucher, Olivier Malavergne, Luc Bouiller, Claude Darrieumerlou, and Francois Laissus. Supporting semantic interoperability in conservationrestoration domain: The PARCOURS project. Journal on Computing and Cultural Heritage (JOCCH), 10(3):16:1–16:??, August 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Nofal:2020:STG

[NPR^{+}20] Eslam Nofal, Georgia Panagiotidou, Rabee M. Reffat, Hendrik Hameeuw, Vanessa Boschloos, and Andrew Vande Moere. Situated tangible gamification of heritage for supporting collaborative learning of young museum visitors. Journal on Computing and Cultural Heritage (JOCCH), 13(1):3:1– 3:24, February 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3350427.

Ngo:2023:QAC

[NTM^{+}23] Duyen Thi Ngo, Cuong Viet Ta, Chau Thi Ma, Hoa Minh Nguyen, Hung Xuan Nguyen, and Ha Thanh Le. Quality assessment criteria and methods for 3D digital replica of historical printing woodblocks. Journal on Computing and Cultural Heritage (JOCCH), 16 (3):43:1–43:??, September 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/doi/
10.1145/3532854.

Noble:2022:NAS

[NVCD22] Laura Noble, Valentina Vavassori, Alan Crookham, and Stuart Dunn. Networking the archive: The stories and structures of Thos. Agnew’s stock books. Journal on Computing and Cultural Heritage (JOCCH), 15(1): 12:1–12:14, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3479009.

Nikolakopoulou:2022:DUE

[NVN^{+}22] Vasiliki Nikolakopoulou, Spyros Vosinakis, Giorgos Nikopoulos, Modestos Stavrakis, Nikolaos Politopoulos, Labros Fragkedis, and Panayiotis Koutsabasis. Design and user experience of a hybrid mixed reality installation that promotes Tinian marble crafts heritage. Journal on Computing and Cultural Heritage (JOCCH), 15(4):60:1– 60:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3522743.

Niu:2023:ACF

[NWLZ23] Xiaolei Niu, Qifeng Wang, Bin Liu, and Jianxin Zhang. An automatic chinaware fragments reassembly method framework based on linear feature of fracture surface contour. Journal on Computing and Cultural Heritage (JOCCH), 16(1):22:1– 22:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3569091.

Othman:2021:ICC

[OAAA21] Mohd Kamal Othman, Shaziti Aman, Nurfarahani Norman Anuar, and Ikram Ahmad. Improving children’s cultural heritage experience using gamebased learning at a living museum. Journal on Computing and Cultural Heritage (JOCCH), 14(3):39:1– 39:24, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3453073.

Othman:2024:CUE

[OAB^{+}24] Mohd Kamal Othman, Nurfarahani Norman Anuar, Mohamad Hardyman Barawi, Abdulrazak Saleh Al-Hababi Yahya, and Ahmad Azaini Abdul Manaf. A comprehensive user experience analysis of cultural heritage progressive Web app using a hybrid UEQ-IPA approach. Journal on Computing and Cultural Heritage (JOCCH), 17(2): 26:1–26:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3647998.

Okura:2015:MRW

[OKY15] Fumio Okura, Masayuki Kanbara, and Naokazu Yokoya. Mixed-reality world exploration using image-based rendering. Journal on Computing and Cultural Heritage (JOCCH), 8 (2):9:1–9:??, March 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Otero:2022:BCH

[OMRP22] David Otero, Patricia Martin-Rodilla, and Javier Parapar. Building cultural heritage reference collections from social media through pooling strategies: The case of 2020’s tensions over race and heritage. Journal on Computing and Cultural Heritage (JOCCH), 15(1):9:1– 9:13, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3477604.

Origlia:2021:MSD

[ORD^{+}21] Antonio Origlia, Silvia Rossi, Sergio Di Martino, Francesco Cutugno, and Maria Laura Chiacchio. Multiple-source data collection and processing into a graph database supporting cultural heritage applications. Journal on Computing and Cultural Heritage (JOCCH), 14(4):55:1– 55:27, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3465741.

Obonyo:2011:DST

[OTBC11] Victor Obonyo, Douglas Troy, Daryl Baldwin, and James Clarke. Digital smartpen technology and revitalization of the Myaamia language. Journal on Computing and Cultural Heritage (JOCCH), 4(4):12:1–12:??, December 2011. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

ODwyer:2021:VVA

[OZY^{+}21] Néill O’Dwyer, Emin Zerman, Gareth W. Young, Aljosa Smolic, Siobhán Dunne, and Helen Shenton. Volumetric video in augmented reality applications for museological narratives: a user study for the Long Room in the Library of Trinity College Dublin. Journal on Computing and Cultural Heritage (JOCCH), 14(2):22:1– 22:20, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3425400.

Pietroni:2014:IVR

[PA14] Eva Pietroni and Andrea Adami. Interacting with virtual reconstructions in museums: The Etruscanning Project. Journal on Computing and Cultural Heritage (JOCCH), 7(2): 9:1–9:??, June 2014. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Payntar:2023:MTA

[Pay23] Nicole D. Payntar. A multitemporal analysis of archaeological site destruction using Landsat satellite data and machine learning, Moche Valley, Peru. Journal on Computing and Cultural Heritage (JOCCH), 16(3):47:1– 47:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3586079.

Pintus:2023:ELS

[PAZ^{+}23] Ruggero Pintus, Moonisa Ahsan, Antonio Zorcolo, Fabio Bettio, Fabio Marton, and Enrico Gobbetti. Exploiting local shape and material similarity for effective SV-BRDF reconstruction from sparse multi-light image collections. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 39:1–39:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3593428.

Pinilla-Buitrago:2019:IHR

[PBCOMTRR19] Laura Alejandra Pinilla-Buitrago, Jesús Ariel Carrasco-Ochoa, José Fco Martínez-Trinidad, and Edgar Román-Rangel. Improved hieroglyph representation for image retrieval. Journal on Computing and Cultural Heritage (JOCCH), 12(2): 13:1–13:??, June 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3284388.

Palma:2010:DSE

[PCC^{+}10] Gianpaolo Palma, Massimiliano Corsini, Paolo Cignoni, Roberto Scopigno, and Mark Mudge. Dynamic shading enhancement for reflectance transformation imaging. Journal on Computing and Cultural Heritage (JOCCH), 3(2):6:1– 6:??, September 2010. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Partarakis:2022:RSH

[PDK^{+}22] Nikolaos N. P. Partarakis, Paraskevi P. D. Doulgeraki, Effie E. K. Karuzaki, Ilia I. A. Adami, Stavroula S. N. Ntoa, Daniele D. M. Metilli, Valentina V. B. Bartalesi, Carlo C. M. Meghini, Yannis Y. M. Marketakis, Danai D. M. Kaplanidi, Maria M. T. Theodoridou, and Xenophon X. Z. Zabulis. Representation of socio-historical context to support the authoring and presentation of multimodal narratives: The Mingei online platform. Journal on Computing and Cultural Heritage (JOCCH), 15(1):16:1– 16:26, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465556.

Pio:2015:DNP

[PFF^{+}15] Gianvito Pio, Fabio Fumarola, Antonio E. Felle, Donato Malerba, and Michelangelo Ceci. Discovering novelty patterns from the ancient Christian inscriptions of Rome. Journal on Computing and Cultural Heritage (JOCCH), 7(4):22:1–22:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Piche:2023:CHR

[PFZG23] Dominique Piché, Ludovic Font, Amal Zouaq, and Michel Gagnon. Comparing heuristic rules and masked language models for entity alignment in the literature domain. Journal on Computing and Cultural Heritage (JOCCH), 16(3):62:1– 62:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3606699.

Pintus:2015:FRF

[PG15] Ruggero Pintus and Enrico Gobbetti. A fast and robust framework for semiautomatic and automatic registration of photographs to 3D geometry. Journal on Computing and Cultural Heritage (JOCCH), 7(4): 23:1–23:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Pavoni:2022:AAS

[PGD^{+}22] Gaia Pavoni, Francesca Giuliani, Anna De Falco, Massimiliano Corsini, Federico Ponchio, Marco Callieri, and Paolo Cignoni. On assisting and automatizing the semantic segmentation of masonry walls. Journal on Computing and Cultural Heritage (JOCCH), 15(2): 22:1–22:17, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3477400.

Pedersen:2017:MTM

[PGMBR17] Isabel Pedersen, Nathan Gale, Pejman Mirza-Babaei, and Samantha Reid. More than meets the eye: The benefits of augmented reality and holographic displays for digital cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 10(2):11:1– 11:??, April 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Polig:2022:RQI

[PHB22] Martina Polig, Sorin Hermon, and Joachim Bretschneider. Resolution and quality issues in 3D analysis of inscribed signs: an example from Cypro–Minoan inscriptions. Journal on Computing and Cultural Heritage (JOCCH), 15(1): 14:1–14:12, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3465334.

Papavassileiou:2023:GMM

[PKO23] Katerina Papavassileiou, Dimitrios I. Kosmopoulos, and Gareth Owens. A generative model for the Mycenaean Linear B script and its application in infilling text from ancient tablets. Journal on Computing and Cultural Heritage (JOCCH), 16(3):52:1– 52:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3593431.

Poulovassilis:2020:CKB

[PLC^{+}20] Alexandra Poulovassilis, Nick Larsson, Fiona Candlin, Jamie Larkin, and Andrea Ballatore. Creating a knowledge base to research the history of UK Museums through rapid application development. Journal on Computing and Cultural Heritage (JOCCH), 12(4):30:1–30:??, January 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Power:2017:IAO

[PLP^{+}17] Christopher Power, Andrew Lewis, Helen Petrie, Katie Green, Julian Richards, Mark Eramian, Brittany Chan, Ekta Walia, Isaac Sijaranamual, and Maarten De Rijke. Improving archaeologists’ online archive experiences through usercentred design. Journal on Computing and Cultural Heritage (JOCCH), 10(1): 3:1–3:??, April 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Pack:2022:ABP

[PLSL22] Chulwoo Pack, Yi Liu, Leen-Kiat Soh, and Elizabeth Lorang. Augmentationbased pseudo-ground truth generation for deep learning in historical document segmentation for greater levels of archival description and access. Journal on Computing and Cultural Heritage (JOCCH), 15(3):41:1– 41:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3485845.

Pretto:2020:TEI

[PMG^{+}20] Niccoló Pretto, Edoardo Micheloni, Silvia Gasparotto, Carlo Fantozzi, Giovanni De Poli, and Sergio Canazza. Technologyenhanced interaction with cultural heritage: an antique Pan flute from Egypt. Journal on Computing and Cultural Heritage (JOCCH), 13(2):8:1– 8:20, June 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3355395.

Paneva-Marinova:2024:FFA

[PMGZ^{+}24] Desislava Paneva-Marinova, Maxim Goynov, Yanislav Zhelev, Mariya Monova-Zheleva, Emanuela Mitreva, Detelin Luchev, Radoslav Pavlov, and Lilia Pavlova. Full-fledged access and usability of content in a digital cultural heritage library: Approaches, paradigms, and implementation. Journal on Computing and Cultural Heritage (JOCCH), 17(1):8:1– 8:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3631135.

Pessoa:2024:ZGJ

[PMHdF24] Larissa Pessoa, Lia Martins, Meng Hsu, and Rosiane de Freitas. ZoAM GameBot: a journey to the lost computational world in the Amazonia. Journal on Computing and Cultural Heritage (JOCCH), 17(3):48:1– 48:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3657303.

Portnova:2021:EEC

[Por21] Tatiana V. Portnova. Exploring the experience of contemporary dance practices in the context of global art choreography in the museum space. Journal on Computing and Cultural Heritage (JOCCH), 14(4):50:1– 50:6, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3460456.

Parthasarathy:2023:NVB

[PP23] Niveditha Parthasarathy and Yogesh Palanichamy. Novel video benchmark dataset generation and real-time recognition of symbolic hand gestures in Indian dance applying deep learning techniques. Journal on Computing and Cultural Heritage (JOCCH), 16(3):49:1– 49:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3586207.

Polig:2021:AVP

[PPBH21] Martina Polig, Despina G.

Papacharalambous, Nikolas Bakirtzis, and Sorin Hermon. Assessing visual perception in heritage sites with visual acuity: Case study of the cathedral of St. John the Theologian in Nicosia, Cyprus. Journal on Computing and Cultural Heritage (JOCCH), 14(1):1:1– 1:18, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3417710.

Paolanti:2023:ELO

[PPF^{+}23] Marina Paolanti, Mariapaola Puggioni, Emanuele Frontoni, Lorella Giannandrea, and Roberto Pierdicca. Evaluating learning outcomes of virtual reality applications in education: a proposal for digital cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 36:1–36:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3593432.

Perez-Pascual:2024:CII

[PPGLPMR24] Asun Pérez-Pascual, Jose Luis Giménez-López, Daniel Palacio, and José Marín-Roig. O-City: Implementation of an innovative multimedia platform for promoting orange economy. Journal on Computing and Cultural Heritage (JOCCH), 17(1):1:1– 1:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3631121.

Pessanha:2022:CLO

[PS22] Francisca Pessanha and Almila Akdag Salah. A computational look at oral history archives. Journal on Computing and Cultural Heritage (JOCCH), 15(1): 6:1–6:16, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3477605.

Papaioannou:2017:ROC

[PSA^{+}17] Georgios Papaioannou, Tobias Schreck, Anthousis Andreadis, Pavlos Mavridis, Robert Gregor, Ivan Sipiran, and Konstantinos Vardis. From reassembly to object completion: a complete systems pipeline. Journal on Computing and Cultural Heritage (JOCCH), 10(2):8:1– 8:??, April 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Puglisi:2015:AEP

[PSBM15] Giovanni Puglisi, Filippo Stanco, Germana Barone, and Paolo Mazzoleni. Automatic extraction of petrographic features from pottery of archaeological interest. Journal on Computing and Cultural Heritage (JOCCH), 8(3):13:1– 13:??, May 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Pujol-Tost:2019:DWJ

[PT19]  Laia Pujol-Tost.  Did

we just travel to the past? Building and evaluating with cultural presence different modes of VRmediated experiences in virtual archaeology. Journal on Computing and Cultural Heritage (JOCCH), 12(1): 2:1–2:??, February 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3230678.

Pirbazari:2022:RIC

[PT22] Alireza Gholinejad Pirbazari and Sina Kamali Tabrizi. RecorDIM of Iran’s cultural heritage using an online virtual museum, considering the coronavirus pandemic. Journal on Computing and Cultural Heritage (JOCCH), 15(2): 37:1–37:14, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3500925.

Pratisto:2023:VRP

[PTP23] Eko Harry Pratisto, Nik Thompson, and Vidyasagar Potdar. Virtual reality at a prehistoric museum: Exploring the influence of system quality and personality on user intentions. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 33:1–33:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3585425.

Pages-Vila:2024:DRP

[PVPMP24] Anna Pagès-Vilà, Xavier Pueyo, and Imanol Muñoz-Pandiella. Digital reconstruction of partially lost altarpieces. The case of the Rosary’s altarpiece of Sant Pere Màrtir de Manresa. Journal on Computing and Cultural Heritage (JOCCH), 17(3):40:1– 40:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3652860.

Pentzold:2017:DWO

[PWM^{+}17] Christian Pentzold, Esther Weltevrede, Michele Mauri, David Laniado, Andreas Kaltenbrunner, and Erik Borra. Digging Wikipedia: The online encyclopedia as a digital cultural heritage gateway and site. Journal on Computing and Cultural Heritage (JOCCH), 10(1): 5:1–5:??, April 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Pintus:2015:AAT

[PYR15] Ruggero Pintus, Ying Yang, and Holly Rushmeier. ATHENA: Automatic text height extraction for the analysis of text lines in old handwritten manuscripts. Journal on Computing and Cultural Heritage (JOCCH), 8(1): 1:1–1:??, February 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Pisa:2011:SPC

[PZF11] Cecilia Pisa, Fabiana Zeppa, and Gabriele Fangi. Spherical photogrammetry for cultural heritage — San Galgano Abbey and the Roman Theater, Sabratha. Journal on Computing and Cultural Heritage (JOCCH), 4(3):9:1–9:??, December 2011. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Rodriguez:2016:DMP

[RAB^{+}16] Marcos Balsa Rodríguez, Marco Agus, Fabio Bettio, Fabio Marton, and Enrico Gobbetti. Digital Mont’e Prama: Exploring large collections of detailed 3D models of sculptures. Journal on Computing and Cultural Heritage (JOCCH), 9(4):18:1–18:??, December 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Rolin:2019:PCD

[RAB^{+}19] Raphael Rolin, Eduard Antaluca, Jean-Louis Batoz, Fabien Lamarque, and Mathieu Lejeune. From point cloud data to structural analysis through a geometrical hBIM-oriented model. Journal on Computing and Cultural Heritage (JOCCH), 12(2):9:1– 9:??, June 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3242901. See ad-

dendum [Nic19a].

Rezaei:2023:MTI

[RAD^{+}23] Alireza Rezaei, Emanuel Aldea, Piercarlo Dondi, Sylvie Le Hégarat-Mascle, and Marco Malagodi. Multitemporal image analysis for preventive conservation of historical musical instruments. Journal on Computing and Cultural Heritage (JOCCH), 16(2):35:1– 35:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3575866.

Rubino:2015:ILB

[RBXM15] Irene Rubino, Claudia Barberis, Jetmir Xhembulla, and Giovanni Malnati. Integrating a location-based mobile game in the museum visit: Evaluating visitors’ behaviour and learning. Journal on Computing and Cultural Heritage (JOCCH), 8(3):15:1– 15:??, May 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ress:2018:MHO

[RCB^{+}18] S. Ress, F. Cafaro, D. Bora, D. Prasad, and D. Soundarajan. Mapping history: Orienting museum visitors across time and space. Journal on Computing and Cultural Heritage (JOCCH), 11(3):16:1– 16:??, September 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Reunanen:2015:HUC

[RDH15] Markku Reunanen, Lily Díaz, and Tommi Horttana. A holistic user-centered approach to immersive digital cultural heritage installations: Case Vrouw Maria. Journal on Computing and Cultural Heritage (JOCCH), 7(4):24:1– 24:??, February 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Rennick-Egglestone:2016:FMD

[REBK^{+}16] S. Rennick-Egglestone, P. Brundell, B. Koleva, S. Benford, M. Roussou, and C. Chaffardon. Families and mobile devices in museums: Designing for integrated experiences. Journal on Computing and Cultural Heritage (JOCCH), 9(2):11:1– 11:??, May 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Raptis:2019:DGD

[RFA19] George E. Raptis, Christos Fidas, and Nikolaos Avouris. Do game designers’ decisions related to visual activities affect knowledge acquisition in cultural heritage games? An evaluation from a human cognitive processing perspective. Journal on Computing and Cultural Heritage (JOCCH), 12(1):4:1– 4:??, February 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3292057.

Ragusa:2019:EVL

[RFB^{+}19] Francesco Ragusa, Antonino Furnari, Sebastiano Battiato, Giovanni Signorello, and Giovanni Maria Farinella. Egocentric visitors localization in cultural sites. Journal on Computing and Cultural Heritage (JOCCH), 12(2):11:1– 11:??, June 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3276772.

Rest:2022:IBA

[RFW^{+}22] Christopher Rest, Denis Fisseler, Frank Weichert, Turna Somel, and Gerfrid G. W. Müller. Illumination-based augmentation for cuneiform deep neural sign classification. Journal on Computing and Cultural Heritage (JOCCH), 15(3):50:1– 50:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3495263.

Remondino:2009:MCD

[RGRG09] Fabio Remondino, Stefano Girardi, Alessandro Rizzi, and Lorenzo Gonzo. 3D modeling of complex and detailed cultural heritage using multi-resolution data. Journal on Computing and Cultural Heritage (JOCCH), 2(1):2:1– 2:??, July 2009. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Rantala:2022:WNW

[RJIH22] Heikki Rantala, Ilkka Jokipii, Esko Ikkala, and Eero Hyvönen. WarVictim-Sampo 1914–1922: a National War Memorial on the Semantic Web for digital humanities research and applications. Journal on Computing and Cultural Heritage (JOCCH), 15(1): 8:1–8:18, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3477606.

Righetto:2024:EUF

[RKG^{+}24] Leonardo Righetto, Mohammad Khademizadeh, Andrea Giachetti, Federico Ponchio, Davit Gigilashvili, Fabio Bettio, and Enrico Gobbetti. Efficient and user-friendly visualization of neural relightable images for cultural heritage applications. Journal on Computing and Cultural Heritage (JOCCH), 17(4):54:1– 54:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3690390.

Rattanarungrot:2024:PST

[RKWC24] Sasithorn Rattanarungrot, Kosin Kalarat, Martin White, and Kannattha Chaisriya. Preserving southern Thai traditional manora dance through mobile role-playing game technology. Journal on Computing and Cultural Heritage (JOCCH), 17(1):11:1– 11:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3631122.

Rovera:2021:EBA

[RNP21] Marco Rovera, Federico Nanni, and Simone Paolo Ponzetto. Event-based access to historical Italian war memoirs. Journal on Computing and Cultural Heritage (JOCCH), 14(1):2:1– 2:23, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3406210.

Roussopoulos:2010:IPA

[RPA^{+}10] Panayiotis Roussopoulos, Constantin Papaodysseus, Dimitris Arabadjis, Mihalis Exarhos, and Michail Panagopoulos. Image and pattern analysis for the determination of the method of drawing celebrated Thera wall-paintings circa 1650 B.C. Journal on Computing and Cultural Heritage (JOCCH), 3(2): 3:1–3:??, September 2010. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Reuter:2010:ADV

[RRC^{+}10] Patrick Reuter, Guillaume Riviere, Nadine Couture, Stephanie Mahut, and Loic Espinasse. ArcheoTUI — driving virtual reassemblies with tangible 3D interaction. Journal on Computing and Cultural Heritage (JOCCH), 3(2):4:1– 4:??, September 2010. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Rinaldo:2023:CLA

[RRDC23] Constance Rinaldo, Diane Rielinger, Joseph Deveer, and Danielle Castronovo. Connecting libraries, archives, and museums: Collections in support of natural history science. Journal on Computing and Cultural Heritage (JOCCH), 16(1):3:1– 3:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3570905.

Roman-Rangel:2016:CRA

[RRJBMM16] Edgar Roman-Rangel, Diego Jimenez-Badillo, and Stephane Marchand-Maillet. Classification and retrieval of archaeological potsherds using histograms of spherical orientations. Journal on Computing and Cultural Heritage (JOCCH), 9(3):17:1–17:??, November 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ridel:2014:RFI

[RRL^{+}14] Brett Ridel, Patrick Reuter, Jeremy Laviole, Nicolas Mellado, Nadine Couture, and Xavier Granier. The revealing flashlight: Interactive spatial augmented reality for detail exploration of cultural heritage artifacts. Journal on Computing and Cultural Heritage (JOCCH), 7(2):6:1– 6:??, June 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Rosenzweig:2023:CAA

[RSM23] Sebastian Rosenzweig, Frank Scherbaum, and Meinard Müller. Computer-assisted analysis of field recordings: a case study of Georgian funeral songs. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 13:1–13:??, March 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3551645.

Ruiz:2022:CAB

[RTA22] Rafael Melendreras Ruiz, Ma Teresa Marín Torres, and Paloma Sánchez Allegue. Comparative analysis between the main 3D scanning techniques: Photogrammetry, terrestrial laser scanner, and structured light scanner in religious imagery: The case of the Holy Christ of the Blood. Journal on Computing and Cultural Heritage (JOCCH), 15(1):18:1– 18:23, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3469126.

Ronzino:2022:DSA

[RTF22] Paola Ronzino, Anna Toth, and Bianca Falcidieno. Documenting the structure and adaptive reuse of Roman amphitheatres through the CIDOC CRMba model. Journal on Computing and Cultural Heritage (JOCCH), 15(2):36:1– 36:23, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3485466.

Sapirstein:2019:SRV

[Sap19] Philip Sapirstein. Segmentation, reconstruction, and visualization of ancient inscriptions in 2.5D. Journal on Computing and Cultural Heritage (JOCCH), 12(2): 15:1–15:??, June 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3286977.

Senhaji:2022:RMA

[SB22] Mohammed Senhaji and Rachid Benslimane. 3D reconstruction of Moroccan– Andalusian Muqarnas domes. Journal on Computing and Cultural Heritage (JOCCH), 15(2):23:1–23:27, June 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3479712.

Souibgui:2022:UPH

[SBC^{+}22] Mohamed Ali Souibgui, Asma Bensalah, Jialuo Chen, Alicia Fornés, and Michelle Waldispühl. A user perspective on HTR methods for the automatic transcription of rare scripts: The case of Codex Runicus. Journal on Computing and Cultural Heritage (JOCCH), 15(4):72:1– 72:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3519306.

Stefani:2013:WPC

[SBL^{+}13] Chiara Stefani, Chawee Busayarat, Julie Lombardo, Livio De Luca, and Philippe Véron. A Web platform for the consultation of spatialized and semantically enriched iconographic sources on cultural heritage buildings. Journal on Computing and Cultural Heritage (JOCCH), 6(3): 13:1–13:??, July 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Simou:2022:GBM

[SBN22] Sana Simou, Khadija Baba, and Abderrahman Nounah. A GIS-based methodology to explore and manage the historical heritage of Rabat City (Morocco). Journal on Computing and Cultural Heritage (JOCCH), 15(4):74:1– 74:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3517142.

Seifert:2017:UAD

[SBO^{+}17] Christin Seifert, Werner Bailer, Thomas Orgel, Louis Gantner, Roman Kern, Hermann Ziak, Albin Petit, Jörg Schlötterer, Stefan Zwicklbauer, and Michael Granitzer. Ubiquitous access to digital cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 10(1): 4:1–4:??, April 2017. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Sepulveda:2024:RPR

[SBS24] Sebastian Sepulveda, Benjamin Bustos, and Ivan Sipiran. Repetitive patterns recognition in textures of ancient Peruvian pottery. Journal on Computing and Cultural Heritage (JOCCH), 17(4):55:1– 55:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3694789.

Sartini:2023:IOC

[SBvE^{+}23] Bruno Sartini, Sofia Baroncini, Marieke van Erp, Francesca Tomasi, and Aldo Gangemi. ICON: an ontology for comprehensive artistic interpretations. Journal on Computing and Cultural Heritage (JOCCH), 16(3):59:1– 59:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3594724.

Sanchez-Belenguer:2015:APT

[SBVVSL^{+}15] Carlos Sánchez-Belenguer, Eduardo Vendrell-Vidal, Miguel Sánchez-López, Carmen Díaz-Marín, and Elvira Aura-Castro. Automatic production of tailored packaging for fragile archaeological artifacts. Journal on Computing and Cultural Heritage (JOCCH), 8(3):17:1– 17:??, May 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Silva:2022:TPD

[SCCF22] Felipe Bruno Silva, Ana Regina Cuperschmid, Ana Lúcia Cerávolo, and Márcio Fabrício. A technological prospect for a diagnostic model in HBIM. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 79:1–79:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3526091.

Siotto:2015:APS

[SCDS15] Eliana Siotto, Marco Callieri, Matteo Dellepiane, and Roberto Scopigno. Ancient polychromy: Study and virtual reconstruction using open source tools. Journal on Computing and Cultural Heritage (JOCCH), 8(3):16:1– 16:??, May 2015. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Setzu:2024:EAI

[SCM^{+}24] Mattia Setzu, Silvia Corbara, Anna Monreale, Alejandro Moreo, and Fabrizio Sebastiani. Explainable authorship identification in cultural heritage applications. Journal on Computing and Cultural Heritage (JOCCH), 17(3):44:1– 44:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3654675.

Scopigno:2012:E

[Sco12] Roberto Scopigno. Editorial. Journal on Computing and Cultural Heritage (JOCCH), 5(2):5:1– 5:??, July 2012. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Scopigno:2013:E

[Sco13] Roberto Scopigno. Editorial. Journal on Computing and Cultural Heritage (JOCCH), 6(1):1:1– 1:??, March 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Shin:2012:ASF

[SDF^{+}12] Hijung Shin, Christos Doumas, Thomas Funkhouser, Szymon Rusinkiewicz, Kenneth Steiglitz, Andreas Vlachopoulos, and Tim Weyrich. Analyzing and simulating fracture patterns of Theran wall paintings. Journal on Computing and Cultural Heritage (JOCCH), 5(3):10:1– 10:??, October 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Samaan:2016:CRP

[SDH^{+}16] Mariam Samaan, Marc Pierrot Deseilligny, Raphaële Heno, Etienne De La Vaissière, and Jacques Roger. Close-range photogrammetric tools for epigraphic surveys. Journal on Computing and Cultural Heritage (JOCCH), 9(3):16:1–16:??, November 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Scopigno:2017:MDA

[SE17] Roberto Scopigno and Karina Rodriguez Echavarria. In memoriam: David Arnold (1951–2016). Journal on Computing and Cultural Heritage (JOCCH), 10(1):1:1–1:??, April 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Sizikova:2018:WPR

[SF18] Elena Sizikova and Thomas Funkhouser. Wall painting reconstruction using a genetic algorithm. Journal on Computing and Cultural Heritage (JOCCH), 11(1): 3:1–3:??, January 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Smithies:2022:MTA

[SFB^{+}22] James Smithies, Pascal Flohr, Fadi Bala’awi, Sahar Idwan, Carol Palmer, Alessandra Esposito, Shatha Mubaideen, and Shaher Rababeh. MaDiH(): a transnational approach to building digital cultural heritage capacity. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 71:1–71:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3513261.

Sero:2021:STD

[SGH^{+}21] Dzemila Sero, Isabelle Garachon, Erma Hermens, Robert Van Liere, and Kees Joost Batenburg. The study of three-dimensional fingerprint recognition in cultural heritage: Trends and challenges. Journal on Computing and Cultural Heritage (JOCCH), 14(4):51:1– 51:20, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3461341.

Sprugnoli:2021:ASA

[SGMT21] Rachele Sprugnoli, Marco Guerini, Giovanni Moretti, and Sara Tonelli. Are these artworks similar? analysing visitors’ judgements on aesthetic perception with a digital game. Journal on Computing and Cultural Heritage (JOCCH), 14(4):52:1– 52:14, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3461663.

Sichani:2022:CHB

[SH22] Anna-Maria Sichani and David Hendy. Connected histories of the BBC: Opening up the BBC Oral History Archive to the digital domain. Journal on Computing and Cultural Heritage (JOCCH), 15(1):10:1– 10:16, February 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3480954.

Schreiber:2012:EPL

[SHFS12] Stephan Schreiber, Klaus G.

Hinzen, Claus Fleischer, and Sven Schütte. Excavationparallel laser scanning of a medieval cesspit in the archaeological zone Cologne, Germany. Journal on Computing and Cultural Heritage (JOCCH), 5(3):12:1– 12:??, October 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Sovhyra:2023:PID

[SIS^{+}23] Tetiana Sovhyra, Iryna Ivashchenko, Viktoriia Strelchuk, Kateryna Pyvovarova, and Anatolii Tykhomyrov. The problem of introduction of digital technologies in the performing arts. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 23:1–23:??, March 2023. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3587169.

Stepputat:2019:DMI

[SKD19] Kendra Stepputat, Wolfgang Kienreich, and Christopher S. Dick. Digital methods in intangible cultural heritage research: a case study in Tango Argentino. Journal on Computing and Cultural Heritage (JOCCH), 12(2):12:1– 12:??, June 2019. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3279951.

Shishido:2020:AOM

[SKK^{+}20] Hidehiko Shishido, Emi Kawasaki, Youhei Kawamura, Toshiya Matsui, and Itaru Kitahara. Accurate overlapping method of ultra-long interval timelapse images for world heritage site investigation. Journal on Computing and Cultural Heritage (JOCCH), 13(2):10:1– 10:18, June 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3373357.

Semeraro:2012:FBR

[SLD^{+}12] Giovanni Semeraro, Pasquale Lops, Marco De Gemmis, Cataldo Musto, and Fedelucio Narducci. A folksonomy-based recommender system for personalized access to digital artworks. Journal on Computing and Cultural Heritage (JOCCH), 5(3):11:1– 11:??, October 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Serres:2023:VCM

[SLRV23] Barthélémy Serres, Damien Létienne, Olivier Roussey, and Gilles Venturini. VISIT: a content management and exploration system for mobile augmented reality in the context of digital humanities and cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 16(2):30:1– 30:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3582264.

Setty:2018:RIB

[SM18] Shankar Setty and Uma Mudenagudi. Region of interest-based 3D inpainting of cultural heritage artifacts. Journal on Computing and Cultural Heritage (JOCCH), 11(2):9:1– 9:??, June 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Suvari:2023:VRA

[SOÇT23] Anil Süvari, S. Ebru Okuyucu, Gamze Çoban, and Emine Eren Tarakci. Virtual reconstruction with the augmented reality technology of the cultural heritage components that have disappeared: The Ayazini Virgin Mary Church. Journal on Computing and Cultural Heritage (JOCCH), 16(1):17:1– 17:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579361.

Sapirstein:2016:PMA

[SP16] Philip Sapirstein and Eric Psota. Pattern matching and the analysis of damaged ancient objects: The case of the column drum. Journal on Computing and Cultural Heritage (JOCCH), 9(3):13:1– 13:??, November 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Sarti:2023:FOD

[SPC^{+}23] Beatrice Sarti, Alice Plutino, Arianna Crespi, Giulia Morabito, and Alessandro Rizzi. FiRe 2: an online database for photographic and cinematographic film technical data. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 9:1–9:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3532520.

Stefanidi:2022:TTC

[SPZ^{+}22] Evropi Stefanidi, Nikolaos Partarakis, Xenophon Zabulis, Ilia Adami, Stavroula Ntoa, and George Papagiannakis. Transferring traditional crafts from the physical to the virtual world: an authoring and visualization method and platform. Journal on Computing and Cultural Heritage (JOCCH), 15(2):35:1– 35:24, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3484397.

Selmanovic:2020:IAI

[SRH^{+}20] Elmedin Selmanović, Selma Rizvic, Carlo Harvey, Dusanka Boskovic, Vedad Hulusic, Malek Chahin, and Sanda Sljivo. Improving accessibility to intangible cultural heritage preservation using virtual reality. Journal on Computing and Cultural Heritage (JOCCH), 13(2):13:1– 13:19, June 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3377143.

Schwartz:2013:WBS

[SRWK13] Christopher Schwartz, Roland Ruiters, Michael Weinmann, and Reinhard Klein. WebGL-based streaming and presentation of objects with bidirectional texture functions. Journal on Computing and Cultural Heritage (JOCCH), 6(3):11:1– 11:??, July 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Suominen:2013:GLF

[SS13] Jaakko Suominen and Anna Sivula. Gaming legacy? Four approaches to the relation between cultural heritage and digital technology. Journal on Computing and Cultural Heritage (JOCCH), 6(3):12:1– 12:??, July 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Sartini:2024:MEE

[SSLdB24] Bruno Sartini, Sarah Binta Alam Shoilee, Claudia A. Libbi, and Victor de Boer. Multivocal exhibition: Exploring cultural perspectives through user-curated art exhibitions. Journal on Computing and Cultural Heritage (JOCCH), 17(4): 73:1–73:??, December 2024. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3679020.

Shamir:2012:CAA

[ST12] Lior Shamir and Jane A.

Tarakhovsky. Computer analysis of art. Journal on Computing and Cultural Heritage (JOCCH), 5(2): 7:1–7:??, July 2012. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Soler:2013:DCH

[STLL13] F. Soler, J. C. Torres, A. J.

León, and M. V. Luzón. Design of cultural heritage information systems based on information layers. Journal on Computing and Cultural Heritage (JOCCH), 6(4):15:1–15:??, November 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Stork:2022:ACM

[Sto22] David G. Stork. Automatic computation of meaning in authored images such as artworks: a grand challenge for AI. Journal on Computing and Cultural Heritage (JOCCH), 15(4):65:1– 65:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3471619.

Savasta:2022:ECT

[SU22] Daniele Savasta and Zeynep Tuna Ultav. Exposing collections through interaction ecologies: a prototype for architectural ephemera. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 64:1–64:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3531528.

Schneider:2024:PPA

[SV24] Stefanie Schneider and Ricarda Vollmer. Poses of people in art: a dataset for human pose estimation in digital art history. Journal on Computing and Cultural Heritage (JOCCH), 17(4):61:1– 61:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3696455.

Shui:2023:LFA

[SWZG23] Wuyang Shui, Pianpian Wei, Xia Zheng, and Shengling Geng. A landmark-free approach for surface asymmetry detection and profile drawings from bilaterally symmetrical geometry. Journal on Computing and Cultural Heritage (JOCCH), 16(2): 34:1–34:??, June 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3589247.

Sun:2010:ACZ

[SZF^{+}10] Qing Sun, Deyun Zhang, Yifeng Fan, Kaizhong Zhang, and Bin Ma. Ancient Chinese zither (guqin) music recovery with support vector machine. Journal on Computing and Cultural Heritage (JOCCH), 3(2):5:1–5:??, September 2010. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Suissa:2022:TPS

[SZGE22] Omri Suissa, Maayan Zhitomirsky-Geffet, and Avshalom Elmalech. Toward a periodspecific optimized neural network for OCR error correction of historical Hebrew texts. Journal on Computing and Cultural Heritage (JOCCH), 15(2):24:1– 24:20, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3479159.

Suissa:2023:AGN

[SZGE23] Omri Suissa, Maayan Zhitomirsky-Geffet, and Avshalom Elmalech. Around the GLOBE: Numerical aggregation question-answering on heterogeneous genealogical knowledge graphs with deep neural networks. Journal on Computing and Cultural Heritage (JOCCH), 16(3):51:1– 51:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3586081.

Seidl:2011:GTD

[SZMB11] Markus Seidl, Matthias Zeppelzauer, Dalibor Mitrović, and Christian Breiteneder. Gradual transition detection in historic film material

— a systematic study. Journal on Computing and Cultural Heritage (JOCCH), 4(3):10:1–10:??, December 2011. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Tai:2023:DAS

[Tai23] Nan-Ching Tai. Digital archiving of the spatial experience of cultural heritage sites with ancient lighting. Journal on Computing and Cultural Heritage (JOCCH), 16(3):45:1– 45:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3589229.

Torsi:2020:INI

[TAR20] Silvia Torsi, Carmelo Ardito, and Cristina Rebek. An interactive narrative to improve cultural heritage experience in elementary school children. Journal on Computing and Cultural Heritage (JOCCH), 13(3): 22:1–22:14, October 2020.

CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3382771.

Trentin:2023:BWV

[TF23] Mia Trentin and Achille Felicetti. Between written and visual communication: CIDOC CRM ontology for Medieval and Early Modern European graffiti. Journal on Computing and Cultural Heritage (JOCCH), 16(3):55:1– 55:??, September 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3589230.

Thuswaldner:2009:DAO

[TFK^{+}09] Barbara Thuswaldner, Simon Flöry, Robert Kalasek, Michael Hofer, Qi-Xing Huang, and Hilke Thür. Digital anastylosis of the Octagon in Ephesos. Journal on Computing and Cultural Heritage (JOCCH), 2(1):1:1–1:??, July 2009. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Trumpy:2015:ODD

[TG15] Giorgio Trumpy and Rudolf Gschwind. Optical detection of dust and scratches on photographic film. Journal on Computing and Cultural Heritage (JOCCH), 8 (2):7:1–7:??, March 2015. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Turner-Jones:2024:DDP

[TJTHW24] Richard N. Turner-Jones, Gervase Tuxworth, Robert A. Haubt, and Lynley Wallis. Digitising the deep past: Machine learning for rock art motif classification in an educational citizen science application. Journal on Computing and Cultural Heritage (JOCCH), 17(4):64:1– 64:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3665796.

Tong:2024:ACV

[TLL^{+}24] Lingwei Tong, Robert W. Lindeman, Heide Lukosch, Rory Clifford, and Holger Regenbrecht. Applying cinematic virtual reality with adaptability to indigenous storytelling. Journal on Computing and Cultural Heritage (JOCCH), 17(2): 28:1–28:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3647996.

Tennent:2020:TEV

[TMB^{+}20] Paul Tennent, Sarah Martindale, Steve Benford, Dimitrios Darzentas, Pat Brundell, and Mat Collishaw. Thresholds: Embedding virtual reality in the museum. Journal on Computing and Cultural Heritage (JOCCH), 13(2): 12:1–12:35, June 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3369394.

Troncoso:2022:OBA

[Tro22] Alvaro R. Ortiz Troncoso. Ontology-based approach to creating semantic Wikis. Journal on Computing and Cultural Heritage (JOCCH), 15(2):28:1– 28:7, June 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3479012.

Trevino:2014:ANP

[TS14] Jeffrey Treviño and Craig Sapp. Automated notation of piano recordings for historic performance practice study. Journal on Computing and Cultural Heritage (JOCCH), 7(3):17:1– 17:??, August 2014. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Tzima:2021:ARO

[TSB21] Stavroula Tzima, Georgios Styliaras, and Athanasios Bassounas. Augmented reality in outdoor settings: Evaluation of a hybrid image recognition technique. Journal on Computing and Cultural Heritage (JOCCH), 14(3):31:1– 31:17, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3439953.

Tsiviltidou:2023:DSO

[TV23] Zoi Tsiviltidou and Giasemi Vavoula. Digital stories with the online collection of the V&A for inquiry-based learning. Journal on Computing and Cultural Heritage (JOCCH), 16(1):18:1– 18:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3570330.

Tan:2024:JER

[TWZF24] Yuyao Tan, Hao Wang, Zibo Zhao, and Tao Fan. A joint entity-relation detection and generalization method based on syntax and semantics for Chinese intangible cultural heritage texts. Journal on Computing and Cultural Heritage (JOCCH), 17(1):5:1– 5:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631124.

Usman:2020:VIV

[UA20] Mahmud Ahmed Usman and Apostolos Antonacopoulos. VISE: an interface for visual search and exploration of museum collections. Journal on Computing and Cultural Heritage (JOCCH), 12(4):25:1– 25:??, January 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ulvi:2022:UUP

[Ulv22]  Ali Ulvi.  Using UAV

photogrammetric technique for monitoring, change detection, and analysis of archeological excavation sites. Journal on Computing and Cultural Heritage (JOCCH), 15(3):58:1– 58:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3522742.

Ullrich:2008:SFR

[USF08] Torsten Ullrich, Volker Settgast, and Dieter W. Fellner. Semantic fitting and reconstruction. Journal on Computing and Cultural Heritage (JOCCH), 1(2): 12:1–12:??, October 2008. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Ulvi:2022:CWM

[UY22] Ali Ulvi and Abdurahman Yasin Yigit. Comparison of the Wearable Mobile Laser Scanner (WMLS) with other point cloud data collection methods in cultural heritage: a case study of Diokaisareia. Journal on Computing and Cultural Heritage (JOCCH), 15(4): 81:1–81:??, December 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3551644.

vonBuelow:2022:DFS

[vBTS^{+}22] Max von Buelow, Reimar Tausch, Martin Schurig, Volker Knauthe, Tristan Wirth, Stefan Guthe, Pedro Santos, and Dieter W. Fellner. Depth-of-field segmentation for nearlossless image compression and 3D reconstruction. Journal on Computing and Cultural Heritage (JOCCH), 15(3):49:1– 49:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3500924.

Volpe:2011:SES

[VC11] Gualtiero Volpe and Antonio Camurri. A system for embodied social active listening to sound and music content. Journal on Computing and Cultural Heritage (JOCCH), 4(1):2:1– 2:??, August 2011. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Villaespesa:2021:CVT

[VC21] Elena Villaespesa and Seth Crider. Computer vision tagging the Metropolitan Museum of Art’s collection: a comparison of three systems. Journal on Computing and Cultural Heritage (JOCCH), 14(3):28:1– 28:17, July 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3446621.

Viola:2023:NMN

[Vio23]  Lorella Viola.  Net-

works of migrants’ narratives: a post-authentic approach to heritage visualisation. Journal on Computing and Cultural Heritage (JOCCH), 16(1):5:1– 5:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3575863.

Vinayavekhin:2024:IRC

[VKS^{+}24] Phongtharin Vinayavekhin, Banphatree Khomkham, Vorapong Suppakitpaisarn, Phillippe Codognet, Torahiko Terada, and Atsushi Miura. Identifying relationships and classifying western-style paintings: Machine learning approaches for artworks by Western artists and Meiji-era Japanese artists. Journal on Computing and Cultural Heritage (JOCCH), 17(1):6:1– 6:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631136.

Volioti:2018:NUI

[VMH^{+}18] Christina Volioti, Sotiris Manitsaris, Edgar Hemery, Stelios Hadjidimitriou, Vasileios Charisis, Leontios Hadjileontiadis, Eleni Katsouli, Fabien Moutarde, and Athanasios Manitsaris. A natural user interface for gestural expression and emotional elicitation to access the musical intangible cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 11(2): 10:1–10:??, June 2018. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Vasic:2024:VVV

[VQP^{+}24] Iva Vasic, Ramona Quattrini, Roberto Pierdicca, Adriano Mancini, and Bata Vasic. 3VR: Vice versa virtual reality algorithm to track and map user experience. Journal on Computing and Cultural Heritage (JOCCH), 17(3):43:1– 43:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3656346.

Vilar:2024:ERP

[VRC24] João Vilar, Armanda Rodrigues, and Nuno Correia. An extended reality platform for cultural gaming: Enabling interactive narratives in spatial contexts. Journal on Computing and Cultural Heritage (JOCCH), 17(4):66:1– 66:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3702484.

Vendrell-Vidal:2014:DAP

[VVSB14] Eduardo Vendrell-Vidal and Carlos Sánchez-Belenguer. A discrete approach for pairwise matching of archaeological fragments. Journal on Computing and Cultural Heritage (JOCCH), 7 (3):15:1–15:??, June 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Venkata:2022:AHR

[VYBG22] Santhilata Kuppili Venkata, Paul Young, Mark Bell, and Alex Green. Alexa, is this a historical record? Journal on Computing and Cultural Heritage (JOCCH), 15(1): 7:1–7:20, February 2022. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3479008.

Woolford:2013:EAG

[WD13] Kirk Woolford and Stuart Dunn. Experimental archaeology and games: Challenges of inhabiting virtual heritage. Journal on Computing and Cultural Heritage (JOCCH), 6(4):16:1–16:??, November 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Wang:2019:VID

[WHC^{+}19] Haiyan Wang, Zhongshi He, Dingding Chen, Yongwen Huang, and Yiman He. Virtual inpainting for Dazu rock carvings based on a sample dataset. Journal on Computing and Cultural Heritage (JOCCH), 12(3):20:1–20:??, October 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3303767.

Weber:2023:ISI

[WHMW23] Andreas Weber, Maarten Heerlien, Eulàlia Gassó Miracle, and Katherine Wolstencroft. Introduction to the special issue on digital natural and cultural heritage: Opportunities and challenges. Journal on Computing and Cultural Heritage (JOCCH), 16(1): 1:1–1:??, March 2023. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3597459.

Wang:2014:ITC

[WKNC14] Xuan Wang, Eng Tat Khoo, Ryohei Nakatsu, and Adrian Cheok. Interacting with traditional Chinese culture through natural language. Journal on Computing and Cultural Heritage (JOCCH), 7(3): 18:1–18:??, June 2014. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Wang:2019:IDM

[WLZ19] Huan Wang, Qingquan Li, and Qin Zou. Inpainting of Dunhuang murals by sparsely modeling the texture similarity and structure continuity. Journal on Computing and Cultural Heritage (JOCCH), 12(3):17:1–17:??, October 2019. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3280790.

Weiss:2024:MSA

[WM24] Christof Weiß and Meinard Müller. From music scores to audio recordings: Deep pitch-class representations for measuring tonal structures. Journal on Computing and Cultural Heritage (JOCCH), 17(3):45:1– 45:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3659103.

Wilczek:2018:CAO

[WMJ^{+}18] Josef Wilczek, Fabrice Monna, Ahmed Jébrane, Catherine Labruère Chazal, Nicolas Navarro, Sébastien Couette, and Carmela Chateau Smith. Computer-assisted orientation and drawing of archaeological pottery. Journal on Computing and Cultural Heritage (JOCCH), 11(4):22:1–22:??, December 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3230672.

Wecker:2022:TSM

[WRKK^{+}22] Alan J. Wecker, Vered Raziel-Kretzmer, Benjamin Kiessling, Daniel Stökl Ben Ezra, Moshe Lavee, Tsvi Kuflik, Dror Elovits, Moshe Schorr, Uri Schor, and Pawel Jablonski. Tikkoun Sofrim: Making ancient manuscripts digitally accessible: The case of Midrash Tanhuma. Journal on Computing and Cultural Heritage (JOCCH), 15(2):20:1– 20:20, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3476776.

Wang:2018:COT

[WSA^{+}18] Zeyu Wang, Weiqi Shi, Kiraz Akoglu, Eleni Kotoula, Ying Yang, and Holly Rushmeier. CHER-Ob: a tool for shared analysis and video dissemination. Journal on Computing and Cultural Heritage (JOCCH), 11(4):18:1–18:??, December 2018. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3230673.

Winarno:2024:HFS

[WSHH24] Edy Winarno, Anindita Septiarini, Wiwien Hadikurniawati, and Hamdani Hamdani. The hybrid features and supervised learning for batik pattern classification. Journal on Computing and Cultural Heritage (JOCCH), 17(2):18:1– 18:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3631131.

Willard:2022:CDF

[WWS^{+}22] Charlie Willard, Nancy Wade, Matija Strlic, John R.

Gilchrist, Tim Weyrich, and Adam Gibson. Correction of dropped frames in high-resolution push-broom hyperspectral images for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 15(2): 29:1–29:19, June 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3479011.

Weiss:2021:SWD

[WZAM^{+}21] Christof Weiß, Frank Zalkow, Vlora Arifi-Müller, Meinard Müller, Hendrik Vincent Koops, Anja Volk, and Harald G. Grohganz. Schubert Winterreise Dataset: a multimodal scenario for music analysis. Journal on Computing and Cultural Heritage (JOCCH), 14(2):25:1– 25:18, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3429743.

Wang:2021:PMF

[WZL^{+}21] Haiping Wang, Yufu Zang, Fuxun Liang, Zhen Dong, Hongchao Fan, and Bisheng Yang. A probabilistic method for fractured cultural relics automatic reassembly. Journal on Computing and Cultural Heritage (JOCCH), 14(1):5:1– 5:25, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3417711.

Xu:2024:HAD

[XLL^{+}24] Ningning Xu, Yue Li, Jiachen Liang, Kexiang Shuai, Yuwen Li, Jiaqi Yan, Cheng Zhang, and Yiping Dong. Heritage-Site AR: Design and evaluation of a mobile augmented reality exploration game for a Chinese heritage site. Journal on Computing and Cultural Heritage (JOCCH), 17(4):67:1– 67:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3700881.

Yalniz:2009:OAE

[YAGU09] Ismet Zeki Yalniz, Ismail Sengor Altingovde, Ugur Güdükbay, and Özgür Ulusoy. Ottoman Archives Explorer: a retrieval system for digital Ottoman archives. Journal on Computing and Cultural Heritage (JOCCH), 2(3):8:1– 8:??, December 2009. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Yannopoulos:2013:DAT

[Yan13] Angelos Yannopoulos. DirectorNotation: Artistic and technological system for professional film directing. Journal on Computing and Cultural Heritage (JOCCH), 6(1):2:1– 2:??, March 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Yu:2013:DSC

[YH13] Chih-Hao Yu and Jane Hunter. Documenting and sharing comparative analyses of 3D digital museum artifacts through Semantic Web annotations. Journal on Computing and Cultural Heritage (JOCCH), 6(4):18:1–18:??, November 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Yi:2021:UER

[YK21] Ji Hyun Yi and Hae Sun Kim. User experience research, experience design, and evaluation methods for museum mixed reality experience. Journal on Computing and Cultural Heritage (JOCCH), 14(4):48:1– 48:28, December 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3462645.

Yates:2024:SSD

[YL24] Mia F. Yates and Anders S. Løvlie. Subtle sound design: Designing for experience blend in a historic house museum. Journal on Computing and Cultural Heritage (JOCCH), 17(1):10:1– 10:??, March 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3633476.

Yue:2022:DDA

[YLFM22] Xuebin Yue, Hengyi Li, Yoshiyuki Fujikawa, and Lin Meng. Dynamic dataset augmentation for deep learning-based oracle bone inscriptions recognition. Journal on Computing and Cultural Heritage (JOCCH), 15(4):76:1– 76:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3532868.

Yengui:2020:STA

[YN20] Ameni Yengui and Mahmoud Neji. SIAT: Tunisian Archaeological Information Systems. Journal on Computing and Cultural Heritage (JOCCH), 13(3):19:1– 19:16, October 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3383781.

Yang:2017:ASP

[YPGR17] Ying Yang, Ruggero Pintus, Enrico Gobbetti, and Holly Rushmeier. Automatic single page-based algorithms for medieval manuscript analysis. Journal on Computing and Cultural Heritage (JOCCH), 10(2):9:1–9:??, April 2017. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Yan:2024:IRT

[YZH^{+}24] Yao Yan, Rui Zhang, Hao He, Tong Lei, Xusheng Zhang, and Chao Jiang. Image restoration technology of Tang Dynasty tomb murals using adversarial edge learning. Journal on Computing and Cultural Heritage (JOCCH), 17(3):52:1– 52:??, September 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3674984.

Zachos:2024:UTU

[ZA24] Aristeidis Zachos and Christos-Nikolaos Anagnostopoulos. Using TLS, UAV, and MR methodologies for 3D modelling and historical recreation of religious heritage monuments. Journal on Computing and Cultural Heritage (JOCCH), 17(4):56:1– 56:??, December 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3679021.

Zouaoui:2021:PCH

[ZDC21] Mohamed Amine Zouaoui, Boualem Djebri, and Antonio Capsoni. From point cloud to HBIM to FEA, the case of a vernacular architecture: Aggregate of the Kasbah of Algiers. Journal on Computing and Cultural Heritage (JOCCH), 14(1):4:1– 4:21, February 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3418039.

Zheng:2024:HAT

[ZJCN24] Xia Zheng, Yicheng Jiang, Hua Cheng, and Aiqing Nie. How to arrange texts and pictures for online visitors — comparing basic ceramic display forms with eye tracking. Journal on Computing and Cultural Heritage (JOCCH), 17(2): 22:1–22:??, June 2024. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3639938.

Zhao:2020:SCC

[ZKB^{+}20] Shichao Zhao, David Kirk, Simon Bowen, David Chatting, and Peter Wright. Supporting the cross-cultural appreciation of traditional Chinese puppetry through a digital gesture library. Journal on Computing and Cultural Heritage (JOCCH), 12(4):28:1–28:??, January 2020. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3341882.

Zhou:2022:IDD

[ZLS^{+}22] Zhiheng Zhou, Xinran Liu, Junyuan Shang, Junchu Huang, Zhihao Li, and Haiping Jia. Inpainting digital Dunhuang murals with structure-guided deep network. Journal on Computing and Cultural Heritage (JOCCH), 15(4):77:1– 77:??, December 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3532867.

Zohar:2013:ATC

[ZLSD13] Hadas Zohar, Chaya Liebeskind, Jonathan Schler, and Ido Dagan. Automatic thesaurus construction for cross generation corpus. Journal on Computing and Cultural Heritage (JOCCH), 6(1):4:1– 4:??, March 2013. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Zhao:2020:SDV

[ZM20] Zhenjie Zhao and Xiaojuan Ma. ShadowPlay2.5D: a 360-degree video authoring tool for immersive appreciation of classical Chinese poetry. Journal on Computing and Cultural Heritage (JOCCH), 13(1):5:1– 5:20, February 2020. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3352590.

Zabulis:2022:DTC

[ZMD^{+}22] Xenophon Zabulis, Carlo Meghini, Arnaud Dubois, Paraskevi Doulgeraki, Nikolaos Partarakis, Ilia Adami, Effie Karuzaki, Anne-Laure Carre, Nikolaos Patsiouras, Danae Kaplanidi, Daniele Metilli, Valentina Bartalesi, Chris Ringas, Eleana Tasiopoulou, and Zinovia Stefanidi. Digitisation of traditional craft processes. Journal on Computing and Cultural Heritage (JOCCH), 15(3):53:1– 53:??, September 2022. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3494675.

Zilio:2021:FSG

[ZOZ21] Daniel Zilio, Nicola Orio, and Luca Zamparo. Fake-Muse: a serious game on authentication for cultural heritage. Journal on Computing and Cultural Heritage (JOCCH), 14(2):17:1– 17:22, June 2021. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3441627.

Zeppelzauer:2016:ISR

[ZPS^{+}16] Matthias Zeppelzauer, Georg Poier, Markus Seidl, Christian Reinbacher, Samuel Schulter, Christian Breiteneder, and Horst Bischof. Interactive 3D segmentation of rock-art by enhanced depth maps and gradient preserving regularization. Journal on Computing and Cultural Heritage (JOCCH), 9(4):19:1– 19:??, December 2016. CO-DEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Zollhofer:2016:LCR

[ZSV^{+}16] M. Zollhöfer, C. Siegl, M. Vetter, B. Dreyer, M. Stamminger, Serdar Aybek, and F. Bauer. Lowcost real-time 3D reconstruction of large-scale excavation sites. Journal on Computing and Cultural Heritage (JOCCH), 9(1): 2:1–2:??, February 2016. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).